### v2

# When Respondents Don’t Care Anymore: Identifying the Onset of Careless Responding

Abstract
Questionnaires in the behavioral and organizational sciences tend to be lengthy: survey measures comprising hundreds of items are the norm rather than the exception. However, literature suggests that the longer a questionnaire takes, the higher the probability that participants lose interest and start responding carelessly. Consequently, in long surveys a large number of participants may engage in careless responding, posing a major threat to internal validity. We propose a novel method for identifying the onset of careless responding (or an absence thereof) for each participant. It is based on combined measurements of multiple dimensions in which carelessness may manifest, such as inconsistency and invariability. Since a structural break in either dimension is potentially indicative of carelessness, the proposed method searches for evidence for changepoints along the combined measurements. It is highly flexible, based on machine learning, and provides statistical guarantees on its performance. An empirical application on data from a seminal study on the incidence of careless responding reveals that the reported incidence has likely been substantially underestimated due to the presence of respondents that were careless for only parts of the questionnaire. In simulation experiments, we find that the proposed method achieves high reliability in correctly identifying carelessness onset, discriminates well between careless and attentive respondents, and captures a variety of careless response types, even when a large number of careless respondents are present. Furthermore, we provide freely available open source software to enhance accessibility and facilitate adoption by empirical researchers.
Keywords: careless responding, changepoint detection, machine learning, survey methodology
Introduction
Research in the behavioral and organizational sciences often involves the administration of lengthy (online) self-report questionnaires. For instance, common personality measures tend to consist of hundreds of items, such as the Revised NEO Personality Inventory (Costa & McCrae,, 1992) with 240 items or the Minnesota Multiphasic Personality Inventory-2 Restructured Form (Ben-Porath & Tellegen,, 2008) with 338 items. Even if one does not use such extensive measures, the number of items can easily reach three digits by including several shorter measures. However, it is well known that questionnaire length can have a concerning adverse effect on measurement accuracy: Respondents may experience fatigue or boredom as they progress through a lengthy questionnaire, which can provoke careless responding (e.g., Bowling et al.,, 2021; Gibson & Bowling,, 2020; Galesic & Bosnjak,, 2009), in particular in online administrations (e.g., Johnson,, 2005; Meade & Craig,, 2012). We call such participants partially careless respondents as they initially provide accurate responses, but resort to careless responding after a certain item and remain careless for the remainder of the questionnaire.
According to the definition of Ward & Meade, (2023), “careless responding occurs when participants are not basing their response on the item content, and it can occur when a respondent does not read an item, does not understand an item, or is unmotivated to think about what the item is asking”.11endnote: 1Perhaps more commonly used, Huang et al., (2012) define careless responding as “a response set in which the respondent answers a survey measure with low or little motivation to comply with survey instructions, correctly interpret item content, and provide accurate responses”. We prefer the definition of Ward & Meade, (2023), as it highlights that careless responding may not be intentional (cf. Huang et al.,, 2012; Ward & Pond,, 2015; Ward et al.,, 2017). Also note that Huang et al., (2012) refer to careless responding as insufficient effort responding in their definition. Other synonyms for careless responding are participant inattention (Maniaci & Rogge,, 2014), inconsistent responding (Greene,, 1978), protocol invalidity (Johnson,, 2005), and random responding (e.g., Beach,, 1989; Berry et al.,, 1992; Emons,, 2008). Notably, Schroeders et al., (2022) point out that the latter—random responding—may be a misnomer since careless responding can also arise in non-random patterns (e.g., a recurring sequence of 1-2-3-4-5). Careless responding has been identified as a major threat to the validity of research findings through a variety of psychometric and statistical issues such as lower scale reliability, errors in statistical hypothesis testing, and deteriorated model fit (e.g., Arias et al.,, 2020; Huang et al.,, 2012, 2015; Kam & Meyer,, 2015). It is widely prevalent (e.g., Bowling et al.,, 2016; DeSimone et al.,, 2015; Curran,, 2016; Meade & Craig,, 2012), suspected to be present in all questionnaire data (Ward & Meade,, 2023), and already incidence rates as low as 5% can be problematic (Arias et al.,, 2020; Credé,, 2010). Often viewed as merely an insidious source of measurement error, careless responding can be a valuable source of information about study participants and the study itself. For instance, carelessness has been linked to participant personality (e.g., Bowling et al.,, 2016; DeSimone et al.,, 2020) and design factors such as study length (e.g., Bowling et al.,, 2021) and item wording. In particular negatively-worded items are prone to misresponses (e.g., Swain et al.,, 2008; Weijters & Baumgartner,, 2012) triggered by carelessness. Consequently, it is widely recommended to carefully screen survey data for the presence of careless responding (e.g., DeSimone et al.,, 2015; Huang et al.,, 2015).
Numerous methods for identifying careless respondents after questionnaire administration (so-called screeners) have been proposed, for instance consistency indicators such as psychometric synonyms (Meade & Craig,, 2012), longstring indices (Johnson,, 2005), or multivariate outlier analyses (e.g., Curran,, 2016). We refer to Arthur et al., (2021), DeSimone et al., (2015), and Ward & Meade, (2023) for detailed reviews.22endnote: 2A popular pre-administration method for the detection of careless responding is the inclusion of detection items (e.g., Table 1 in Arthur et al.,, 2021). Such detection items are based on the presumption that an attentive respondent will respond in a specific way, while careless respondent may fail to do so. For instance, it is expected that an attentive respondent would strongly disagree to so-called bogus items such as “I am paid biweekly by leprechauns” (Meade & Craig,, 2012). A careless respondent may accidentally “agree” to this item as a consequence of inattention. Alternative types of detection items are instructed items and self-report items (Arthur et al.,, 2021). We discuss detection items and preventive measures against carelessness in detail in a later section. More recently, Schroeders et al., (2022), Alfons & Welz, (2024), and Ozaki, (2024) have studied machine learning techniques for identifying careless respondents. Still, commonly used screeners tend to suffer from three major limitations, which we discuss in the following.
First, extant methods commonly follow the implicit assumption that careless respondents respond carelessly to all questionnaire items or at least a vast majority of items. In particular, they indicate whether a respondent engaged in careless responding, but they typically do not provide any indication regarding which items are affected. Although there exist many comparative studies (e.g., Curran,, 2016; DeSimone & Harms,, 2018; Goldammer et al.,, 2020; Huang et al.,, 2012) on such screeners, their behavior under partial careless responding is not well-understood. This is at odds with the substantial evidence that many careless respondents tend to begin a survey as attentive and truthful respondents, but may resort to careless responding due to fatigue or boredom as a survey progresses (Bowling et al.,, 2021; Brower,, 2018; Galesic & Bosnjak,, 2009; Gibson & Bowling,, 2020; Ward et al.,, 2017). In general, Ward & Meade, (2023) believe it is rare that careless respondents respond carelessly from the beginning to the end of a questionnaire.
Second, common detection methods are typically designed to detect a single specific type of carelessness. For instance, the longstring index of Johnson, (2005) counts the maximum number of consecutive identical responses, which is intended to capture straightlining behavior. However, careless responding may manifest in multiple distinctive ways (e.g., Curran,, 2016; DeSimone et al.,, 2015; DeSimone & Harms,, 2018; Edwards,, 2019; Ward & Meade,, 2023), implying that a screener that is designed to detect one type of carelessness may fail to detect other types. It is therefore recommended to combine multiple screeners to detect various types of careless respondents (e.g., Goldammer et al.,, 2020; Huang et al.,, 2012; Meade & Craig,, 2012; Ward & Meade,, 2023).
Third, cutoff values for flagging careless respondents are often ad-hoc and lack statistical foundation. Screeners typically calculate a score for each participant, with a participant being flagged as careless if their score is sufficiently extreme (i.e., beyond a certain cutoff value). Ideally, such a cutoff value results in high sensitivity (flagging true careless respondents as careless) and high specificity (not flagging true attentive respondents as careless). For certain screeners, various cutoff values have been suggested in the literature, but individual suggestions vary substantially. For instance, for the same screeners, Goldammer et al., (2020) argues for different cutoff values than Curran, (2016) or Johnson, (2005), which suggests that the results for specific cutoff values may depend on factors such as the population or the constructs being measured (cf. Meade & Craig,, 2012, p. 452). For a specific screener, it would be desirable to obtain a cutoff value with statistical guarantees on sensitivity and specificity that does not depend on population or study characteristics.
In wake of these three issues with extant methods, we propose a novel screener for the identification of careless responding that is designed to overcome these issues. We achieve this by focusing on a more nuanced question. Motivated by the prominence of partial careless responding due to survey fatigue or boredom, we focus on identifying at which point in the survey a given participant becomes careless (if at all). Specifically, our method estimates the item at which carelessness onsets in each questionnaire participant, or an absence thereof. Hence, we call our method careless onset detection in extensive rating-scale surveys (CODERS). To capture various types of carelessness, CODERS can combine multiple dimensions of evidence for careless responding to construct a score that, for each item, measures if a given respondent has started responding carelessly by that item. It achieves this by searching for changepoints along a multidimensional series of indices that are potentially indicative of carelessness, such as measures of response consistency or variability.
We argue that the notion of a changepoint is intuitive when studying partial carelessness: Once a participant starts responding carelessly, we expect a structural break in their responses. In particular, such a respondent may abandon content-based responding and resort to careless response types, while no such break occurs in the absence of carelessness. CODERS is highly flexible as it does not assume a statistical model, nor does it require predefining what types of careless response types may exist, and it is primarily intended for lengthy multi-scale surveys. Moreover, CODERS builds on nonparametric estimation theory for cutoff values and statistical guarantees regarding sensitivity and specificity.
Re-examining data from a a seminal study on careless responding in Big 5 measurements (Johnson,, 2005), we demonstrate novel insights that can be obtained by estimating onset items of partially careless responding. Most notably, we find strong evidence that Johnson, (2005) considerably underestimates the incidence of careless respondents due to hitherto undetected presence of partially careless respondents. Through extensive simulation experiments, we find that the proposed method CODERS accurately estimates carelessness onset (or a lack thereof) and thereby achieves excellent sensitivity and specificity in detecting careless responding.
To the best of our knowledge, CODERS is the first attempt to systematically detect the onset of careless responding. Assuming that participants respond accurately in the beginning of the survey, CODERS is designed to segment the responses of each respondent into a segment of accurate responses and—if a changepoint was identified—a segment of careless responses. With this knowledge, researchers can restrict their primary analyses to the segments of accurate responses (by treating the careless responses as missing values) without having to discard all responses of a partially careless respondent. In addition, researchers can separately study the segments of careless responses to build theory on and obtain better understanding of the nature of (partially) careless responding (cf. Gibbert et al.,, 2021, for theory-building based on outlying data points). Finally, we provide the freely available software package carelessonset (Welz & Alfons,, 2024), which implements CODERS in the statistical programming environment R (R Core Team,, 2024) can be downloaded from https://github.com/mwelz/carelessonset. As such, it is a useful and accessible tool for any researcher who is concerned with survey fatigue or careless responding in general.
Introducing CODERS
The literature generally identifies three distinct ways in which careless responding may manifest: low internal consistency of the given responses (inconsistency), low variability of given responses (invariability), and impossibly fast responses (e.g., Curran,, 2016; DeSimone et al.,, 2015; DeSimone & Harms,, 2018; Edwards,, 2019; Ward & Meade,, 2023). We expect partial careless responding to result in a change in respondent behavior in at least one of these dimensions. CODERS therefore offers a general framework for searching for a joint changepoint in any combination of measurements of these dimensions of carelessness. While the majority of this paper is focused on inconsistency and invariability, we revisit response times in a later section.
We propose to measure the first dimension, inconsistency, by means of reconstructions of observed responses (cf. Alfons & Welz,, 2024). These reconstructions are generated by an auto-associative neural network (henceforth autoencoder) (Kramer,, 1992), which learns response patterns that characterize attentive responding. We expect that random-like, content-independent responses cannot be learned well and are thus poorly reconstructed by the autoencoder, leading to a changepoint in reconstruction performance.
For measuring the second dimension, invariability, we propose a novel algorithm that is inspired by the famous longstring index of Johnson, (2005). Since long sequences of identical responses or constant response patterns are not expected in surveys that use positively and negatively coded items in a randomized order, we expect a changepoint in response variability once a respondent commences to respond carelessly through straightlining or pattern responding behavior.
Overall, CODERS attempts to capture evidence for the onset of carelessness by combining multiple indicators that are potentially indicative of such an onset, where the different indicators may capture different manifestations of carelessness. Conversely, if a respondent never becomes careless, we do not expect a changepoint in any dimension. Combining multiple indicators is a generally recommended practice to capture various types of careless responding (Goldammer et al.,, 2020; Huang et al.,, 2012; Meade & Craig,, 2012; Ward & Meade,, 2023).
We provide a detailed description of our assumptions in Appendix A and a technical definition of CODERS in Appendix B.
Quantifying Inconsistency with Autoencoders
Ward & Meade, (2023) define internal consistency of responses as patterns that are expected based on theoretical/logical grounds or trends in the data. For instance, items that are part of the same construct are expected to correlate strongly in most participants, provided that participants are attentive. In contrast, inconsistent careless responding “generate[s] responses that fail to meet an expected level of consistency” (Ward & Meade,, 2023). Respondents may engage in inconsistent careless responding if they attempt to conceal their carelessness, for instance by randomly choosing from all response options (Ward & Meade,, 2023). For our purposes, we consider the defining characteristic of inconsistent careless responding to be content-independent responses that are approximately randomly chosen from the response options, where the probability of being chosen may differ between response options, such as preferring extreme response options.
In order to identify inconsistent careless responding, we propose to use the machine learning method of autoencoders (Kramer,, 1992), which were originally developed to filter random noise in signal processing applications. An autoencoder is a neural network (e.g., Bishop,, 2006; Goodfellow et al.,, 2016) that attempts to reconstruct its input. In other words, the output variables are equal to the input variables. The idea behind reconstructing the input data is to learn the internal structures of the data, thereby separating the signal from random noise that lacks structure. Since we consider the defining characteristic of inconsistent careless responding to be near-randomly chosen responses—which may be viewed as random noise in the context of questionnaire data—we expect autoencoders to perform well in distinguishing between attentive responses and inconsistent careless responses.
To achieve the goal of learning the internal structures of a dataset, it is crucial to avoid that an autoencoder simply passes the input data through its network architecture. Consequently, autoencoders first express the input data in terms of a low-dimensional representation before reconstructing the input data again from this low-dimensional representation (cf. Kramer,, 1992, and Chapter 14.6 in Goodfellow et al.,, 2016). The use of such a low-dimensional representation makes autoencoders appealing for questionnaire data, as questionnaires typically contain multiple items to measure each construct of interest. In this context, the autoencoder aims to learn the patterns between the latent constructs in the low-dimensional space. For attentive responses, randomness that can be attributed to taking multiple measurements of the latent constructs is thereby filtered out. Since the amount of randomness from these multiple measurements should be relatively low for a well-designed questionnaire, attentive responses are expected to be reconstructed well from the learned patterns in the low-dimensional space of the latent constructs. Inconsistent careless responses, on the other hand, should contain little to no structure also in this low-dimensional space, and are therefore expected to be reconstructed poorly.
A schematic example of the network architecture of an autoencoder is given in Figure 1. Concretely, an autoencoder is a fully connected network whose nodes are organized in multiple layers. Each node takes a linear combination of the nodes from the previous layer and subsequently applies a so-called activation function in order to allow for nonlinear relationships. A crucial part of an autoencoder is the central layer, known as the bottleneck layer, which stores the low-dimensional representations of the input data. The network architecture of an autoencoder is symmetric around this bottleneck layer such that the can be viewed as having an encoder-decoder architecture: the encoder compresses the data into the low-dimensional representations, while the decoder reconstructs the input data again. It follows that an autoencoder can be seen as a dimension reduction technique, specifically as a nonlinear generalization of principal component analysis (PCA; see Kramer,, 1991).
Figure 1: Schematic example of a neural network with an autoencoder structure.
$\bm{x}_{1}$$\bm{x}_{2}$$\bm{x}_{3}$$\bm{x}_{4}$$\bm{\widehat{z}}_{1}$$\bm{\widehat{z}}_{2}$$\bm{\widehat{x}}_{1}$$\bm{\widehat{x}}_{2}$$\bm{\widehat{x}}_{3}$$\bm{\widehat{x}}_{4}$nput LayerMapping LayerBottleneck LayerDemapping LayerOutput Layer
Note. The network consists of five fully connected layers and is symmetric around the central layer. Each dimension is represented by one node, and nodes within the same layer are aligned vertically. In each node, a so-called activation function is applied to a linear combination of the nodes from the previous layer. The first layer (input layer) represents the input variables (here $\bm{x}_{1},\dots,\bm{x}_{4}$). The second layer (mapping layer) prepares the data to be compressed in the third and central layer (bottleneck layer), which holds low-dimensional representations of the input data (here $\bm{\widehat{z}}_{1}$ and $\bm{\widehat{z}}_{2}$). The fourth layer (demapping layer) prepares the low-dimensional respresentations for reconstructing the input data in the final layer (output layer), yielding reconstructions in the original dimension (here $\bm{\widehat{x}}_{1},\dots,\bm{\widehat{x}}_{4}$).
Hence, the number of nodes in the bottleneck layer is an important design choice in the architecture of an autoencoder. In the context of questionnaire data, we recommend setting this number equal to the number of constructs measured in the questionnaire. For example, consider data from the Revised NEO Personality Inventory (NEO-PI-R; Costa & McCrae,, 1992), which contains 240 items and measures six so-called facets of each of the Big Five personality traits. Thus, this instrument measures a total of 30 facets, such as anxiety and modesty, so that we would set the number of nodes in the bottleneck layer to 30.
A thorough mathematical definition of autoencoders can be found in Appendix B.1. Although there are other design choices besides the number of nodes in the bottleneck layer (see Appendix B.1.2), our software provides default values that we recommend researchers to use. The only input required by the user is the number of constructs measured in the survey (for the number of nodes in the bottleneck layer).
Suppose now that we have obtained autoencoder reconstructions of the responses of each survey participant. Let there be $n$ participants and $p$ items, and denote the observed response of the $i$-th participant to the $j$-th item by $x_{ij}$ and the corresponding reconstruction by $\widehat{x}_{ij}$, for $i=1,\dots,n$ and $j=1,\dots,p$. We define the reconstruction error $\textsf{RE}_{ij}$ associated with the response of participant $i$ to item $j$ to be the squared difference between the observed response and its reconstruction, scaled by the range of the answer categories. Formally, for participants $i=1,\dots,n$ and items $j=1,\dots,p$,
$\textsf{RE}_{ij}=\left(\frac{x_{ij}-\widehat{x}_{ij}}{L_{j}-1}\right)^{2},$ (1)
where $L_{j}$ denotes the number of answer categories of item $j$. Recall that we expect the reconstruction errors $\textsf{RE}_{ij}$ to be low in case of attentive responding but high in the presence of inconsistent careless responding. Hence, if participant $i$ starts providing inconsistent careless responses at item $k\in\{1,\dots,p\}$, we expect a changepoint at position $k$ in the participant’s series of $p$ reconstruction errors, as reconstruction errors are expected to be higher from item $k$ onward. An example can be found in the top panel of Figure 2(a).
Figure 2: Illustrative example output from CODERS using responses of a participant to a 300-item personality inventory.
(a) Two dimensions indicative of carelessness: autoencoder reconstruction errors (RE) in the top panel, and the LongstringPattern (LSP) sequence in the bottom panel.
(b) The per-response statistic of the changepoint detection test. The dashed horizontal line corresponds to the associated critical value at an extremely conservative significance level of 0.1%.
Note. The solid vertical lines indicate the identified carelessness onset, which is at the 148th item. Reconstruction errors are in general much larger after this item than before, indicating more random-like responses. The response data are taken from Johnson, (2005) and are discussed in detail in a later section.
Quantifying Invariability
We propose to measure response invariability through a novel algorithm inspired by the longstring index of Johnson, (2005) that we call LongstringPattern (LSP). This algorithm exploits that invariable careless responses are characterized by content-independent response patterns. A pattern sequence has the defining property that it consists of repeated occurrences of the same response pattern. Consider the sequence of responses $-2-1-2-\ldots-1-$ 2 - … - 1 - 2, which consists of repeated occurrences of the pattern $1-2$. Denoting the length of a pattern by $l$, the above example exhibits a pattern of length $l=2$. Analogously, the sequence $-3-1-2-3-\ldots-1-2-$ 3 - … - 1 - 2 - 3 consists of repeated occurrences of the pattern $1-2-3$ of length $l=3$, whereas the straightlining sequence $1-1-\ldots-$ 1 - … - 1 consists of repeated occurrences of 1, which is a pattern of length $l=1$.
Define $l\textsf{-pattern}_{ij}$ as the number of consecutive items in the longest sequence of a recurring pattern of length $l$ in which the response of participant $i$ to item $j$ takes part, or assign the value 1 if the response is not part of a recurring pattern of length $l$. Consider the following two illustrative examples. First, for $l=2$, the response sequence $1-2-1-2-1-4-3-5-4-5-4$ will be assigned the 2-pattern sequence $4-4-4-4-1-1-1-4-4-4-4$ because the first four responses and last four responses comprise two occurrences of the distinct patterns $1-2$ and $5-4$, respectively, which are both of length $l=2$. The central three responses ($1-4-3$) are not part of any recurring pattern of length $l=2$, hence they are each assigned the value 1. Second, for a recurring pattern of length $l=1$ (i.e., consecutive identical responses), the response sequence $3-2-3-3-1-4-1-1-1$ is assigned the 1-pattern sequence $1-1-2-2-1-1-3-3-3$, since the subsequences of consecutive identical responses $3-3$ and $1-1-$ 1 - 1 comprise two and three responses, respectively.33endnote: 3The 1-pattern sequence can be seen as a generalization of the the longstring index of Johnson, (2005), which we recover by calculating $\max_{1\leq j\leq p}1\textsf{-pattern}_{ij}$ for each respondent $i=1,\dots,n$. In short, higher values in the $l$-pattern sequence capture a tendency to respond in fixed patterns of length $l$. If a participant starts engaging in invariable careless responding, we expect a changepoint in the $l$-pattern sequence (for some $l$) from relatively low to relatively high values, assuming that the order of items is randomized.
However, an $l$-pattern sequence crucially depends on the choice of pattern length $l$, which is restrictive since invariable careless respondents may choose widely different patterns. Consequently, a single choice of pattern length $l$ is unlikely to capture all careless response patterns that may emerge. To tackle this issue, we propose to calculate the $l$-pattern sequence for multiple choices of $l$, and we define
$\textsf{LSP}_{ij}=\max_{l}\ l\textsf{-pattern}_{ij}$ (2)
for participant $i=1,\dots,n$ and item $j=1,\dots,p$. In this paper, we vary $l$ from 1 to the number of answer categories in the survey. The rationale behind this choice is that we consider it unlikely that careless respondents choose complicated individual patterns whose length exceeds the number of answer categories.
We provide a detailed description of the LSP algorithm in Appendix B.2. An example of an LSP sequence is shown in the bottom panel of Figure 2(a).
Identifying Carelessness Onset via Changepoint Detection
For each participant, we can obtain multiple individual series of length equal to the total number of items in the survey. Each individual series should thereby measure one of the dimensions indicative of carelessness. Although we focus on using reconstruction errors from (1) to measure inconsistency as well as the LSP sequence from (2) to measure invariability, it should be noted that in principle CODERS can also be used with any single measurement or any other combination of different measurements of carelessness.
As motivated in the previous two sections, we expect a changepoint in any of the measurements of carelessness from the onset of careless responding (e.g., Figure 2(a)), and no changepoint in the absence of carelessness. We therefore propose to apply the nonparametric cumulative sum self-normalization test of Shao & Zhang, (2010), which was developed in the context of time series analysis. It searches for the location of a possible changepoint in a multivariate series and has attractive theoretical guarantees (see Shao & Zhang,, 2010; Zhao et al.,, 2022, and our detailed description in Appendix B.3).
For each point in a multivariate series (we use measurements of carelessness associated with each questionnaire item), this changepoint detection test calculates a statistic that compares a quantity of interest (we take the mean) in the periods before and after the current point (item). If the maximum of this statistic exceeds a critical value, the point (item) where this maximum occurs is flagged as a changepoint. Otherwise no changepoint is flagged. The critical value thereby depends only on the significance level $\alpha$ and the dimension of the multivariate series. Importantly, it controls the size of the test regardless of the distribution of the individual series. This gives CODERS great flexibility, as it allows us to use any combination of measurements of carelessness without making possibly unrealistic assumptions on their distributions or obtaining cutoff values in an ad-hoc manner. In line with literature that suggests that there should be overwhelming evidence when labeling respondents as careless (e.g., Huang et al.,, 2012), we recommend using a conservative significance level no greater than 1% for identifying changepoints.
As an example, consider again the two-dimensional series in Figure 2(a), which measures carelessness of a certain survey participant in terms of autoencoder reconstruction errors (RE) and the LongstringPattern (LSP) sequence. Figure 2(b) shows the corresponding series of statistics, where the maximum value at the 148th item clearly exceeds the critical value. That is, CODERS flags this respondent as being careless from the 148th item onwards. Indeed, from roughly this item onwards, the reconstruction errors are in general much larger than before. It follows that this respondent likely is a partially inconsistent careless respondent.
Empirical Application
Ideally, new methodology is empirically validated by conducting an experiment that emulates the phenomenon of interest in lab conditions. For instance, Schroeders et al., (2022) conduct an experiment in which they instruct one part of the participants to respond truthfully and another part to respond carelessly in order to investigate whether various screeners accurately identify the subset of careless respondents. For any of the compared methods, at most 19% of the respondents flagged as careless were in fact instructed to be careless. The authors therefore express doubts whether the study participants complied with the experimental instructions. In addition to the risk of non-compliance, Ulitzsch et al., 2022a ; Ulitzsch et al., 2022b stress that another major limitation of this experimental approach is the doubtful assumption that participants instructed to respond carelessly behave in a comparable manner to careless respondents outside of such lab experiments. Indeed, it seems plausible that careless responding stems from an intrinsic state-of-mind rather than being a conscious action that can be instructed (cf. Alfons & Welz,, 2024, for additional discussion). Hence, rather than conducting such an experiment, we take data from a seminal study on careless responding and further examine the responses for partial carelessness.
Data
We apply CODERS to an empirical data set collected by Johnson, (2005). The data concern the five factor model of personality (Goldberg,, 1992), which assumes that variation in a measurement of personality can be explained by the so-called Big Five personality traits (factors): openness, conscientiousness, extraversion, agreeableness, and neuroticism. Each of the Big Five factors can be further decomposed into six subdomains (facets): for instance, depression and anxiety are facets of the neuroticism trait. Consquently, common Big Five instruments, such as the 240-item NEO-PI-R (Costa & McCrae,, 1992), typically measure a total of $5\times 6=30$ facets. The data set of Johnson, (2005) consists of internet-collected responses to a 300-item representation of the NEO-PI-R inventory (Goldberg,, 1999), in which each of the 30 facets is measured by 10 items on a five-point Likert scale.
[…]
In general, it seems likely that survey fatigue is present in a sample of responses to an online questionnaire with a lengthy battery of 300 items. For instance, Bowling et al., (2021) estimate that if one wishes that at least 90% of participants do not respond carelessly to more than 20% of all items, one should not include more than 176 items in an online questionnaire. For the study of Johnson, (2005), it is important to keep in mind that the author applied two methods for detecting inconsistent careless respondents to the data from which straightliners are already removed. The two methods are designed to detect participants who respond carelessly to all or at least the vast majority of survey items. It is unclear how well they perform in identifying partially careless respondents, who may respond carelessly to only a relatively small number of items. In addition, Johnson, (2005) did not apply any method designed to identify invariable careless responding other than straightlining. By applying CODERS, which is designed to identify partially careless responding for various types of carelessness, new insights into the incidence and nature of (partially) careless responding may emerge.
Results
[…]
Comparison with Benchmark Methods
For obtaining the autoencoder reconstruction errors in CODERS, we set the number of nodes in the bottleneck layer to 30, i.e., equal the number of facets in the Big Five model. Table 1 summarizes the results of CODERS. Depending on the significance level $\alpha$, CODERS flags between 164 (0.7%) participants for $\alpha=0.001$ and 850 (3.8%) participants for $\alpha=0.01$ as partially careless. On average, the location of the carelessness onset item for flagged participants is estimated around the 150th item, meaning that carelessness onsets on average after half of the 300 total items. These findings are broadly in line with the results of Bowling et al., (2021, Figure 1), whose model suggests a low probability that a representative participant starts responding carelessly at the 150th item. The earliest onset occurs at item 12, whereas the latest onset is at item 294. A more detailed visual overview of the detected onset items for the flagged participants can be found in Figure C.1 in Appendix C.2.
For comparison, we consider personal reliability and psychometric antonym scores. Along the given responses, Johnson, (2005) provides per-individual estimates for either score, which we use in our analysis. Importantly, Johnson, reverses psychometric antonym scores so that low scores are associated with carelessness in both methods. Following Johnson,, we flag respondents as careless whose personal reliability and psychometric antonym scores are lower than $0.3$ and $-0.03$, respectively. The results are shown in Table 1. Personal reliability flags 40 participants (0.2%) as careless, while psychometric antonym flags 186 respondents (0.8%). CODERS, personal reliability, and psychometric antonyms generally show little agreement regarding the detected careless respondents. Personal reliability and psychometric antonym scores correlate by a moderate value of 0.488, yielding only 13 participant that are flagged by both methods. Depending on the significance level $\alpha$, CODERS and personal reliability have between 1 and 4 flagged participants in common, while the number of jointly flagged respondents for CODERS and psychometric antonym ranges between 5 and 10. All three methods share a single flagged participant.
Table 1: Comparison of CODERS with benchmark methods for the data of Johnson, (2005).
Also flagged by CODERS
Method Number of flagged respondents $\alpha=0.001$ $\alpha=0.005$ $\alpha=0.01$
Reliability 40 (0.2%) 1 3 4
Antonym 186 (0.8%) 5 7 10
Both 13 1 1 1
[…]
It is not surprising that there is little to no consensus of CODERS with the two benchmark methods, as the former is designed for a different task: finding respondents that have been careless in parts of the questionnaire (including identifying the onset item) rather than throughout the entire questionnaire. This disagreement indicates that partial carelessness may be present in the data, which the benchmark methods may be unable to identify. We therefore take a closer look at a representative selection of respondents and analyze their per-item reconstruction errors and LSP sequences, which are shown in Figure 3. Table 2 contains additional summary statistics for these measurements before and after the estimated careless onset item (if such an item is found by CODERS). Two of the selected participants are flagged as careless by CODERS but not by either of the two benchmark methods, one participant is flagged by both CODERS and psychometric antonym, and one participant is flagged by both benchmark methods but not CODERS.
Figure 3: Scores and careless onset item detected by CODERS in selected participants from the data of Johnson, (2005).
(a) Flagged by CODERS only, presumably inconsistent. Onset is estimated at item 148.
(b) Flagged by CODERS only, presumably invariable. Onset is estimated at item 58.
(c) Flagged by antonym and CODERS, presumably inconsistent. CODERS estimates onset at item 119.
(d) Flagged by antonym and personal reliability, but not CODERS (i.e., there is no estimated onset item).
Note. In each subplot, autoencoder reconstruction errors (RE) and LongstringPattern (LSP) sequences are shown in separate panels. The solid vertical lines indicate the identified careless onset items. Summary statistics for the corresponding participants are provided in Table 2.
Table 2: Average scores before and after the careless onset item detected by CODERS in selected participants from the data of Johnson, (2005).
Before After
RE 0.019 (0.035) 0.069 (0.082)
LSP 4.197 (1.396) 3.497 (1.433)
(a) Flagged by CODERS only, presumably inconsistent. Onset is estimated at item 148.
Before After
RE 0.078 (0.103) 0.025 (0.035)
LSP 2.421 (0.706) 19.770 (10.941)
(b) Flagged by CODERS only, presumably invariable. Onset is estimated at item 58.
Before After
RE 0.019 (0.022) 0.039 (0.059)
LSP 3.619 (1.371) 2.687 (0.972)
(c) Flagged by antonym and CODERS, presumably inconsistent. CODERS estimates onset at item 119.
Overall
RE 0.064 (0.066)
LSP 2.660 (0.753)
(d) Flagged by antonym and personal reliability, but not CODERS (i.e., there is no estimated onset item).
Note. Standard deviations are given in parentheses. The subtables correspond to the four selected participants from Figure 3. Positioning of subtables is consistent with positioning of the subfigures in Figure 3, e.g., Table 2(a) refers to the same participant as Figure 3(a).
First, we focus on Figure 3(a) and Table 2(a). CODERS identifies carelessness onset at item 148, but neither benchmark method flags this respondent as careless. We observe that the average reconstruction error, as well as the variability in the reconstruction errors, is much larger before the estimated onset item than afterward. Since the pre-onset responses could be reconstructed well (except for a brief spike around item 110) while post-onset responses could not, there is evidence that the participant started the questionnaire attentively, but succumbed to survey fatigue about halfway through and subsequently engaged in inconsistent careless respondening.
The second example from Figure 3(b) and Table 2(b) shows another participant that is flagged by CODERS but no benchmark method. The estimated onset of careless responding is at item 58. While the LSP sequence clearly changes around this item, even more excessive values are found for later items, with a maximum LSP value of 31. In fact, only four out of the $n=20,994$ participants in the sample yield more extreme LSP values. Hence, there is strong evidence that this participant responded with (increasingly) lengthy sequences of fixed patterns as the questionnaire progressed. That is, they likely started engaging in invariable careless responding.
The third example, shown in Figure 3(c) and Table 2(c), is similar to the first example in the sense that it suggests an initially attentive respondent starting to give inconsistent careless responses from some point onward in the questionnaire, as indicated by a higher average and a larger variability in the reconstruction errors. Here, the careless onset is estimated at item 119. The crucial difference with the first example is that this participant was not only flagged as careless by CODERS, but also by psychometric antonym. A plausible explanation for this difference may be found in the earlier careless onset. If a participant responds carelessly to a large enough majority of the items, a method designed to detect carelessness throughout the questionnaire (such as the benchmark methods) has a better chance at flagging that participant as careless.
[…]
While these four examples are naturally selective, similar plots as in Figure 3 for all participants flagged by CODERS or the two benchmark methods can be found in our online repository at https://github.com/mwelz/carelessonset_plots. In general, two key insights can be derived from the above analysis. First, we find empirical evidence that CODERS can indeed detect various types of partial careless responding, specifically inconsistent and invariable carelessness. Common screeners for careless responding throughout the questionnaire, such as personal reliability or psychometric antonym, are typically only effective for one type of careless responding. Second, CODERS is not a replacement for existing screeners, but rather a complimentary tool that provides new insights if participants display careless responding behavior only in part of the questionnaire.
Comparison with Johnson, (2005)
We now compare our findings with those reported by Johnson, (2005) regarding the incidence of careless responding. This comparison is summarized in Table 3.
Table 3: Incidence of careless responding in the data of Johnson, (2005).
Johnson, (2005) Our findings
Type of careless responding Method Number Method Number
Straightliners Longstring + screeplot 808 (3.6%)
Longstring + cutoff by Costa & McCrae, (2008) 1455 (6.5%)
Inconsistent Reliability 41 (0.2%) Reliability 40 (0.2%)
Antonym 185 (0.8%) Antonym 186 (0.8%)
Either 213 (1.0%) Either 213 (1.0%)
Partial CODERS ($\alpha=0.001$) 158 (0.7%)
CODERS ($\alpha=0.005$) 580 (2.6%)
CODERS ($\alpha=0.010$) 836 (3.7%)
[…]
Johnson, explores two different cutoff values for the longstring index to identify straightliners: one obtained via a screeplot and another suggested by Costa & McCrae, (2008),66endnote: 6Johnson, (2005) refers to a pre-publication version of Costa & McCrae, (2008). with the latter yielding almost twice as many flagged respondents (3.6% vs. 6.5%). As the data set provided by the author already excludes straightliners (among others), we could not attempt to replicate these findings.
Regarding inconsistent careless responding, we find minor differences in the number of flagged respondents compared to Johnson,: 40 vs. 41 for personal reliability, and 186 vs. 185 for psychometric antonym. These minor differences are potentially due to rounding: the psychometric antonym and personal reliability scores provided by Johnson, are rounded by two decimal places, whereas the scores used to compute the reported results in Johnson, (2005) were likely unrounded. As in Johnson, (2005), we find that personal reliability and psychometric antonym agree on 13 of the flagged participants.
Following strategies that “[eliminate] more non-attentive responders, albeit at the cost of more false positives”, Johnson, (2005, p. 116) applies the cutoff value of Costa & McCrae, (2008) for identifying straightliners, and considers respondents that are found by either personal reliability or psychometric antonym as inconsistent careless respondents. This results in a total of 1668 detected careless respondents (7.4%). Employing a similar strategy regarding CODERS, one may choose a significance level of 1%. Adding the corresponding 836 participants yields a total of 2504 flagged careless respondents (11.2%).
Hence, we find evidence that the incidence of careless responding might be substantially underreported in Johnson, (2005) due to hitherto undetected partially careless respondents. Moreover, an incidence of 11.2% in this data set would be higher than the 5–10% that the literature deems problematic for further analysis (e.g., Arias et al.,, 2020; Credé,, 2010).
Simulation Experiments
After demonstrating how CODERS yields new insights in data from a seminal study, we further study the behavior of CODERS via simulation experiments inspired by existing survey measurements.
Data Generation
We generate rating-scale responses of $n=500$ respondents to $p=300$ items. Each item has five Likert-type answer categories (anchored by 1 = “strongly disagree” and 5 = “strongly agree”). The simulated questionnaire measures 30 constructs, each of which is measured by a scale of 10 items. We simulate reliable scales by imposing that items from the same scale have a correlation of 0.7, which results in a Cronbach’s $\alpha$ value of 0.959.77endnote: 7We use package simstudy (Goldfeld & Wujciak-Jens,, 2020) to generate the data, for which the target values of correlations only hold approximately on the population level. Hence, the same applies to Cronbach’s $\alpha$ values. Items from different scales are uncorrelated. All 10 items within a given scale follow the same response probability distribution. We thereby consider the two types of distributions defined in Table 4: centered about the midpoint and skewed towards agreeing. Each of the two distributions is used in 15 scales, yielding the total of $2\times 15\times 10=300$ items. We reverse five items per scale (i.e., half the items) and randomize the order of the items. Note that distributions skewed towards disagreement are implicitly covered by this design because of the item reversal.
Table 4: Response probability distributions of items within different scales.
Type of distribution $P[X=1]$ $P[X=2]$ $P[X=3]$ $P[X=4]$ $P[X=5]$
Centered 0.15 0.20 0.30 0.20 0.15
Skewed 0.10 0.15 0.20 0.25 0.30
Note. The random variable $X$ denotes the response to an item with five Likert-type answer categories.
Subsequently, we select a fixed percentage $\gamma\in\{0,0.02,0.04,0.1,0.2,\dots,0.6,0.8,1\}$ of respondents to be partially careless. That is, $\gamma$ denotes the prevalence of careless responding. For each of the selected partially careless respondents, we sample a carelessness onset item from which onward the remaining responses are careless. We consider two settings for the carelessness onset: (i) early onset, which we randomly sample from the item set $\{30,31,\dots,269,270\}$ so that careless responding starts between 10% and 90% of all items, and (ii) later onset, which we randomly sample from $\{150,151,\dots,269,270\}$, i.e., between half and 90% of all items.
[…]
We repeat the above described data generating process 100 times. Importantly, we keep the location of carelessness onset item of each participant constant across the 100 repetitions, which allows for neat visualization of the results. We compare three variations of CODERS based on (i) both reconstruction errors and the LSP sequence, (ii) reconstruction errors only, and (iii) the LSP sequence only. For the autoencoder, we set the number of nodes in the bottleneck layer equal to the number of measured constructs (i.e., 30).
Performance Evaluation
To quantify how accurately the onset of careless responding is detected, we compute the adjusted Rand index (ARI) (Hubert & Arabie,, 1985, see Appendix D.1 for details). The ARI is best known as a measure of clustering performance: a value of 0 is expected for a random clustering, while a perfectly recovered clustering results in a value of 1. In the careless onset detection problem for a given respondent, we can view the items before the careless onset as one cluster and the items from the onset onwards as another. Specifically, the closer the ARI is to its maximum value 1, the better CODERS performs at accurately locating the true carelessness onset item. On the other hand, an ARI value close to 0 indicates that the estimated onset is far from the true onset. If a respondent is attentive throughout the questionnaire but CODERS incorrectly identifies a changepoint, the ARI takes a value of 0. The same holds if CODERS fails to flag a changepoint for a partially careless respondent.
We report the average ARI across respondents. Note that averaging the ARI over attentive respondents yields the proportion of attentive respondents who are correctly identified as attentive. Conversely, subtracting this number from 1 gives the proportion of attentive respondents that are incorrectly flagged as partially careless.
Results
We start presenting the results of CODERS for the three dimensions by means of a visual inspection. For a carelessness prevalence level of $\gamma=0.6$ and the early onset setting, Figure 4 displays a heatmap holding the indices of the participants in its rows and the item indices in its columns. For participant $i\in\{1,\dots,n\}$ and item $j\in\{1,\dots,p\}$, the color scale of the cell at position $(i,j)$ measures the frequency across the 100 repetitions with which the response of participant $i$ to item $j$ has been flagged as changepoint: a more saturated red indicates that the corresponding item was flagged more frequently as a changepoint. The black rectangles visualize the true location of the careless onset items (recall that those are fixed across the repetitions). The $y$-axis of the plot is rearranged so that the participant indices are ordered according to their respective carelessness onset. Good detection performance is characterized by flagging cells that are close to the corresponding onset items. Indeed, we can see that there is a saturated red band that closely follows the black rectangles. Note that the top 40% of the rows in Figure 4 very rarely see flagged changepoints. Those are the respondents who never become careless. Based on this visual examination, CODERS seems to not only accurately estimate the careless onset, but also discriminate well between partially careless respondents and fully attentive respondents. Results are similar for other combinations of carelessness prevalence and onset, hence we omit the corresponding heatmaps.
Figure 4: Heatmap of carelessness onset detection performance of CODERS.
Note. Of the $n=500$ respondents, 60% are partially careless according to the early onset setting. The significance level of CODERS is set to 0.1%. A more saturated red in a cell corresponds to a higher frequency across the 100 repetitions of the cell being flagged as carelessness onset item ($x$-axis) of a given respondent ($y$-axis). The black rectangles denote the true carelessness onset items (which are constant across the repetitions) and the $y$-axis is rearranged according to the location of true carelessness onset. An absence of a black rectangle (occurs in 40% of all respondents) indicates that a given respondent never becomes careless.
Figure 5: ARI of CODERS for various carelessness prevalence levels.
Note. The ARI is averaged over fully attentive respondents and partially careless respondents (columns) for different carelessness onset settings (rows), and the significance level of CODERS is set to 0.1%.
Varying the prevalence of careless responding, Figure 5 visualizes the ARI averaged over attentive respondents (left column) and partially careless respondents (right column) for the early onset (top row) and later onset (bottom row) settings. Each of the three variations of CODERS rarely flag changepoints in fully attentive respondents, except when carelessness prevalence is extremely high. While the percentage of false positives is close to the significance level 0.1% until 40% prevalence, overrejection remains moderate at 1% for 60% prevalence, before it rises to about 5% for early onset and 10% for later onset at 80% prevalence. For partially careless respondents, the ARI of CODERS with both dimensions remains remarkably constant throughout carelessness prevalence levels at about 0.75 and 0.85 for early and later onset, respectively, mirroring an excellent performance of flagging changepoints in correct locations. In contrast, the ARI using only the LSP sequences amounts to about 0.6 throughout prevalence levels in both onset settings, whereas the ARI using only the reconstruction errors drops from about 0.5 for 2% prevalence to about 0.3 for 100% prevalence with early onset, and from about 0.75 to about 0.35 with later onset. It is important to note that such a drop in performance is expected when using only one indicator of carelessness due to various types of careless responding being simulated. Reconstruction errors are designed to capture inconsistent carelessness and may therefore miss invariant carelessness, while LSP sequences are designed to capture invariant carelessness and may therefore miss inconsistent carelessness.
Figure 6: ARI of CODERS for various carelessness prevalence levels.
Note. The ARI is averaged over the participants corresponding to each careless response type (columns) for different carelessness onset settings (rows), and the significance level of CODERS is set to 0.1%. Note that for straightlining and pattern responding, the results for CODERS using all two dimensions and LongstringPattern only are nearly identical, resulting in overplotted lines.
To explore this further, Figure 6 displays the ARI averaged over subgroups of participants corresponding to each type of careless responding (separate columns) for early onset (top row) and later onset (bottom row). Indeed, reconstruction errors yield much better performance for inconsistent carelessness (random and extreme responding) than for invariable carelessness (straightlining and pattern responding), in particular in the later onset setting. Conversely, using LSP sequences results in near-perfect detection of straightlining and pattern responding with an ARI of about 0.99, but performs particularly poorly for random responding with an ARI close to 0. This highlights the importance of combining the two dimensions in CODERS for accurately detecting various types of partially careless responding.
We now zoom in on the results in Figure 6 for CODERS using reconstruction errors only, which vary by type of (inconsistent) carelessness, as well as by carelessness prevalence and onset location. First, the performance is clearly better for extreme responding than for random responding, with the ARI being in general higher and remaining more stable across prevalence levels. Considering the relatively flat probability distributions of the attentive responses (see Table 4), distinguishing random responses is simply a much more challenging problem. Second, the heterogeneity in performance across prevalence levels and onset location is likely due to the autoencoder requiring sufficiently many attentive responses to successfully distinguish between attentive and (inconsistent) careless responding. If carelessness tends to onset later in the survey, the autoencoder has more information to learn attentive response patterns, making it easier to distinguish between attentive and careless responses. Similar reasoning applies for carelessness prevalence. When fewer responses are careless, the autoencoder has access to more attentive responses from which to learn. In general, these results suggest that CODERS can accurately identify inconsistent careless responding unless careless responses are excessively present in the data through a high prevalence of careless respondents whose carelessness onset occurs early in the survey. However, having an overwhelmingly large number of participants with largely uninformative responses suggests a fundamental issue with data quality, in case of which any subsequent analysis may yield questionable results.
To summarize, this simulation experiment suggests that CODERS succeeds in combining complimentary indicators for carelessness onset in order to capture different manifestations of careless responding.
Additional Simulations
We have conducted a variety of additional simulation experiments, which are described and discussed in Appendix D.2. Specifically, we investigate different choices of the significance level in CODERS, varying correlation structures (and therefore measurement reliability), as well as varying sample sizes. Overall, the findings presented above are representative for those situations. For low to moderate carelessness prevalence (as in the empirical application), choosing a higher significance level (e.g., 1%) leads to an increase in power while retaining size control. On the other hand, for high prevalence levels (about 40% or higher), increasing the significance level yields too many false positives. As expected, power decreases (increases) slightly with decreasing (increasing) measurement reliability or sample size.
We conclude from our simulation experiments that CODERS reliably detects carelessness onset while successfully discriminating between partially careless respondents and fully attentive respondents. In addition, our simulations highlight the importance of using multiple indicators to capture different manifestations of carelessness (cf. Goldammer et al.,, 2020; Huang et al.,, 2012; Meade & Craig,, 2012; Ward & Meade,, 2023).
Extensions, Limitations and Alternatives
Below, we address the potential incorporation of response times into CODERS, and we discuss general limitations as well as potential alternatives.
Response Times Revisited
In addition to inconsistency and invariability, fast responding is considered the third major manifestation of carelessness (e.g., Ward & Meade,, 2023). In (online) surveys, response time is typically measured by the total time a participant spent on the questionnaire, time spent on each questionnaire page, or (less common) time spent on each questionnaire item.
If per-item response times are available, they can easily be incorporated into CODERS as an additional series to be used in the changepoint detection. If only per-page response times are available and pages are relatively short, an additional series can be obtained by assigning to each response the average page time, i.e., the time the participant spent on the page on which the item is located divided by the number of items on that page (which inspired by the page time index of Bowling et al.,, 2023). If a participant at some point in the survey, for instance, starts responding without reading the items, we expect a changepoint in response time towards considerably faster responses (cf. Bowling et al.,, 2023, 2016; Meade & Craig,, 2012; Huang et al.,, 2012).
Nevertheless, it is unclear if a changepoint in response times is indeed indicative of careless responding in an empirical setting. This may depend on the survey design: if items in one part of the survey are sufficiently longer than in other parts, a changepoint in response times might occur naturally for attentive respondents. On the other hand, if all items are of similar length, a changepoint to (much) shorter response times may indicate carelessness. For instance, there exist instruments whose individual items comprise of single words, such as the Big Five unipolar markers (Goldberg,, 1992) with items like “timid” or “envious”. In such measurements, one would not expect large differences in response times across items. For such survey designs with similar item length, incorporating response times into CODERS may prove fruitful, but a thorough validation is beyond the scope of this paper and therefore left for future research.
Limitations
Throughout this paper, we have assumed that all participants start as truthful and accurate respondents, while some of them may begin to respond carelessly from a certain item onward. However, in certain scenarios, this assumption may not hold true.
First, as we have already seen in our empirical application, some participants may respond carelessly throughout all survey items (cf. Figure 3(d) and Table 2(d)). Consequently, CODERS is unlikely to flag a changepoint due to an absence of a change in behavior. Nevertheless, the literature has established a plethora of methods for detecting respondents who have been careless throughout the questionnaire. In this sense, CODERS—which is intended for detecting partial carelessness—is complementary to those existing methods. If one suspects that respondents who have been careless throughout all items are present, we recommend to apply established detection methods following the guidelines of Arthur et al., (2021) and Ward & Meade, (2023). Respondents that are flagged as careless by such methods, but for whom no changepoint was flagged by CODERS, have likely been careless throughout the entire survey.
Second, there may be some participants who alternate between periods of attentive and careless responding (Berry et al.,, 1992; Clark et al.,, 2003; Meade & Craig,, 2012). In this case, CODERS may only flag one of the possibly multiple changepoints. CODERS should therefore be viewed as a promising first step towards detecting periods of careless responding, with an extension for the detection of multiple changepoints being an area of further research.
Finally, the changepoint detection procedure of Shao & Zhang, (2010) employed by CODERS is based on asymptotic arguments when the number of items is large. Hence, CODERS is designed for lengthy questionnaires, and it is likely that the statistical power drops considerably in short questionnaires. However, partial careless responding due to fatigue may be less prevalent in short questionnaires (cf. Bowling et al.,, 2021). Studying the onset of careless responding in shorter surveys is therefore beyond the scope of this paper.
Potential Alternatives to CODERS
First, the recommendation that one should not include more than three detection items (Meade & Craig,, 2012) is fairly restrictive as carelessness may have onset at a much earlier item than the detection item, resulting in a failure to exclude a potentially large number of careless responses. In principle, a form of carelessness onset detection may be possible by a priori including detection items in the survey, such as bogus, instructed, or self-report items: one may consider a failed detection item (i.e., the participant provides a response that would not be expected from an attentive respondent) as an estimate of the onset of careless responding. However, there are a number of practical and theoretical disadvantages with this approach (see Curran,, 2016, pp. 13–15, for a more general discussion on the caveats of detection items). First, the estimates of the carelessness onset may be crude at best. Recommendations of Meade & Craig, (2012) suggest including at most three bogus, instructed, or self-report items, spaced every 50–100 items. Consequently, the carelessness onset may occur much earlier than at at the detection item, leaving a potentially large number of careless responses undetected. Second, attentive respondents may ironically choose an unexpected response to a bogus item because they find doing so humorous (Meade & Craig,, 2012) or due to misinterpretation of the item (Curran & Hauser,, 2019).88endnote: 8For instance, Curran & Hauser, (2019) find that some respondents would agree to the popular bogus item “All my friends say I would make a great poodle” because they believe that they are loyal friends and they associate loyalty with dogs. Third, a careless respondent may choose the expected response by mere chance or lie in self-report items (DeSimone & Harms,, 2018; Edwards,, 2019), resulting in failure to identify this careless respondent. Fourth, it is unclear if, e.g., the response “somewhat disagree” to a detection item for which “strongly disagree” is expected should be considered careless (Ward & Meade,, 2023). Finally, it may not be possible to include detection items in the survey for legal or ethical reasons, or because the survey analyst has no control over the survey design.
In addition, the software packages careless (Yentes & Wilhelm,, 2021) and ufs (Peters & Gruijters,, 2021) support the calculation of an intra-individual variance (IRV) across subsets of responses.99endnote: 9Specifically, this is implemented in the functions irv() from package careless and carelessReport() from package ufs. If a set of responses towards the end of the questionnaire exhibits a sufficiently larger (Marjanovic et al.,, 2015) or lower (Dunn et al.,, 2018) variance than an earlier set of responses, there might evidence for an onset of partially careless responding. First, these packages only support splitting the responses into subsets of equal length, which does not allow for accurate detection of the carelessness onset. Second, a more general implementation for confirming the presence of partial carelessness in arbitrary subsets would require a priori knowledge on the onset item that defines the subsets, which is unrealistic in practice. Third, it is unclear when a variance can be considered sufficiently low or high to infer carelessness: determining cutoff values below or above which a participant is to be flagged as careless may depend on survey design and survey content.1010endnote: 10For example, if 5-point scales are used a lower cutoff is necessary than for a 9-point scale, since IRVs will be very different for those scales.
Instead of screening for careless responding after data collection, one may attempt to prevent such behavior during survey administration. Preventive measures against careless responding may entail promising rewards for responding accurately (Gibson & Bowling,, 2020), as well as threatening with punishment for careless responding or the presence of a proctor (Bowling et al.,, 2021). There is evidence that each of these three preventive measures can prevent or at least delay carelessness onset (Bowling et al.,, 2021; Gibson & Bowling,, 2020; Huang et al.,, 2012). However, either of these three preventive measures can be undesirable or unethical in practice. Administering proctored questionnaires is often impractical and expensive, and is typically infeasible in online studies. Arthur et al., (2021) and Bowling et al., (2021) advise against using threats if one wishes to maintain a positive long-term relationship with questionnaire participants, such as when there is a professional relationship between researcher and participants. Moreover, threats seem inappropriate in longitudinal studies or when the participants are members of a questionnaire panel (as is common in survey research), or when they are suspected to suffer from anxiety, depression, or trauma (as is common, e.g., in research on wellbeing in the workplace, or more general in psychological research). On the other hand, promising rewards may cross ethical boundaries in organizational surveys and in the context of employment-related decisions (Arthur et al.,, 2021).
Notwithstanding, we stress that preventive methods against carelessness are complementary to detection methods. If preventative methods are appropriate, combining them with screening techniques such as CODERS may further enhance quality and reliability of the collected data for subsequent analysis.
Discussion
We introduced a novel method, CODERS, to identify the onset of careless responding (or absence thereof) for each participant in a given survey. It is motivated by literature stressing that in lengthy surveys—which are common in the behavioral and organizational sciences—a large proportion of survey participants may respond partially carelessly due to fatigue or boredom. CODERS is highly flexible and can combine multiple indicators for the potential onset of carelessness, such as autoencoder reconstruction errors as an indicator of inconsistency, and the proposed LongstringPattern (LSP) sequence as an indicator of invariability. Re-analyzing data from a seminal study on careless responding, we find evidence of partial careless responding that was hitherto undetected. Additionally, simulation experiments indicate excellent performance of CODERS in identifying the onset of careless responding, and highlight the importance of incorporating multiple indicators that capture different manifestations of carelessness.
While CODERS is a promising first step towards detecting periods of careless responding rather than identifying which respondents are careless, various pertinent questions remain for future research. For instance, the LSP sequence as a measurement of response invariability may require additional validation, as does the incorporation of response times into CODERS, which we leave for empirical follow-up work. In addition, extending or adapting CODERS with changepoint detection methods that allow for multiple changepoints, or that retain higher power in shorter surveys, may be fruitful directions for further research.
Whereas we apply machine learning and nonparametric techniques in our approach, other literature is focused on model-based approaches. For instance, Ulitzsch et al., 2022b propose a mixture model based on item response theory that is designed to detect careless respondents, which is adapted in Ulitzsch et al., 2022a to incorporate response times. Other approaches that explicitly model carelessness as part of a mixture model are proposed in Arias et al., (2020), Steinmann et al., (2022), and Van Laar & Braeken, (2022). One advantage of model-based approaches is that their estimates can be used to construct person-weights that can be used to downweight the adverse influence of potentially careless respondents in subsequent analyses, which is discussed in detail in Ulitzsch et al., 2022a ; Ulitzsch et al., 2022b and Hong & Cheng, (2019). However, model-based approaches typically impose strong and potentially unverifiable assumptions on the properties of attentive and careless responding which might be violated in practice.
On a final note, Arthur et al., (2021) envisage “artificial intelligence and machine learning being used to detect careless responding and response distortion in real time in the foreseeable future.” Using machine learning techniques in the form of autoencoders and nonparametric changepoint detection, our proposed method CODERS takes a step towards that future.
Computational Details
Throughout this paper, we have followed recent guidelines by Alfons & Welz, (2024) on the application of machine learning methods for the detection of careless responding. All computations were performed with R version 4.3.3 (R Core Team,, 2024) and Python version 3.10 (Van Rossum & Drake,, 2009). Our implementation of CODERS in the R package carelessonset is available from https://github.com/mwelz/carelessonset. Replication files will be made publicly available upon acceptance of this manuscript and a link will be included here.
Acknowledgments
We thank Donald Bergh, Dennis Fok, Patrick Groenen, Nick Koning, Robin Lumsdaine, Jakob Raymakers, and Kevin ten Haaf for valuable comments and feedback. This work was supported by a grant from the Dutch Research Council (NWO), research program Vidi (Project No. VI.Vidi.195.141).
Appendix A Setup and Assumptions
Throughout our paper, we consider the following setup. Let $\bm{X}$ be an $n\times p$ data matrix holding the rating-scale responses of $n$ survey participants to $p$ items. We make the standard assumption that all observations (i.e., all participants whose responses the survey collects) are independently and identically distributed. Moreover, it is a priori unknown if and when careless responding occurs in $\bm{X}$. We require the following assumptions.
Assumption 1.
The responses in $\bm{X}$ admit an $n\times s$ lower-dimensional representation, $\bm{S}$, where $s\ll p$. The dimension $s$ is known and corresponds to the number of constructs that the survey measures.
Assumption 2.
The survey that generated $\bm{X}$ is reliable in the sense that if all participants responded accurately, $\bm{X}$ would accurately measure all constructs.
Assumption 3.
The items in the survey are in a randomized order.
Assumption 4.
All participants begin the survey as attentive respondents by providing accurate and truthful responses. As the survey progresses, some (possibly none or all) participants start responding carelessly and continue to do so for the remainder of the survey.
Assumption 5.
The onset of partial carelessness is characterized by a changepoint in one or more indicators. We use autoencoder reconstruction errors and LongstringPattern (LSP) sequences as indicators.
Assumption 1 is very mild in survey data because surveys typically measure multiple constructs and the number of constructs is typically known. For instance, the $p=300$ item representation of the NEO-PI-R inventory (Goldberg,, 1999) measures $s=30$ constructs.
Assumption 2 is again mild in survey data. There exist well-established and reliable survey measures for a large variety of variables. For instance, the International Personality Item Pool (IPIP; Goldberg,, 1999) is a pool of more than 250 personality scales. It is in general recommended to use well-established measures to ensure a highly reliable measurement (Babbie,, 2020, p. 150). The reliability of a measurement can be estimated in multiple ways, such as Cronbach’s alpha, the omega coefficient and variations thereof, as well as the $H$ coefficient (see, e.g., McNeish,, 2018, for a description of these and other reliability measures). Hence, Assumption 2 is satisfied when the survey data were collected with measures of high reliability.
Assumption 3 is common practice in surveys in the social and behavioral sciences. The randomized order of items prevents that recurring response patters emerge as a result of the survey design rather than due to invariable careless responding.
Assumption 4 is discussed in the section on limitations, and Assumption 5 is motivated in detail in the section introducing the proposed methodology.
Appendix B Technical Details on CODERS
B.1 Details on Autoencoders
B.1.1 Network Architecture and Estimation
Denote by $x_{ij}$ the response of the $i$-th participant to the $j$-th survey item. Collect the responses of the $i$-th participant to all $p$ items in a $p$-dimensional vector $\mathbf{x}_{i}=(x_{i1},x_{i2},\dots,x_{ip})^{\top}$so that the $n\times p$ data matrix $\bm{X}$ is given by $\bm{X}=(\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{n})^{\top}$.
Introduced by Kramer, (1992), an autoencoder is a network consisting of an uneven number of $M$ layers, in which the first layer (input layer) holds the input data $\mathbf{x}_{i}$ and the last layer (output layer) holds the reconstructed input data $\widehat{\mathbf{x}}_{i}$. Denoting the number of nodes in the $\ell$-th layer by $N^{(\ell)}$, we have that $N^{(1)}=N^{(M)}=p$. Moreover, the central layer (bottleneck layer) consists of $s<p$ nodes to prevent that the input is simply passed through the network, and the network architecture is symmetric around the bottleneck layer.
Each node in a given layer stores so-called activations, which are obtained by transforming a linear combination of the activations from the previous layer. This transformation is defined by an activation function $g_{\ell}(\cdot)$. Formally, for participant $i=1,\dots,n$, the activation $a_{ij}^{(\ell)}$ of the $j$-th node, $j=1,\dots,N^{(\ell)},$ of the $\ell$-th layer, $\ell=2,\dots,M$, is given by
$a_{ij}^{(\ell)}$ $=g_{\ell}\left(z_{ij}^{(\ell)}\right),$
$z_{ij}^{(\ell)}$ $=\sum_{k=1}^{N^{(\ell-1)}}\omega_{jk}^{(\ell)}a_{ik}^{(\ell-1)}+b_{j}^{(\ell)}.$
The $\omega_{jk}^{(\ell)}$ and $b_{j}^{(\ell)}$ are unknown weights and intercept terms, respectively. Observe that in the first layer, we have $a_{ij}^{(1)}=x_{ij}$. The activations of the last layer, $a_{ij}^{(M)}$, hold the network’s output, namely the autoencoder’s reconstructions $\widehat{x}_{ij}$ of responses $x_{ij}$.
We jointly refer to the weights and intercept terms as the network’s parameters, which need to be estimated, and we collect them in a vector $\bm{\theta}$. To emphasize the dependence on the parameter vector $\bm{\theta}$, we define for a fixed $\bm{\theta}$ a prediction function $\bm{f}_{\bm{\theta}}(\cdot)$ of the network, which corresponds to the activations in the last layer:
$\bm{f}_{\bm{\theta}}(\mathbf{x}_{i})=\Big{(}f_{1,\bm{\theta}}(\mathbf{x}_{i}),f_{2,\bm{\theta}}(\mathbf{x}_{i}),\dots,f_{p,\bm{\theta}}(\mathbf{x}_{i})\Big{)}^{\top}=\left(a_{i1}^{(M)},a_{i2}^{(M)},\dots,a_{ip}^{(M)}\right)^{\top},$
for participants $i=1,\dots,n$. The reconstruction of the input vector $\mathbf{x}_{i}$ is then given by $\widehat{\mathbf{x}}_{i}~{}=~{}\bm{f}_{\bm{\theta}}(\mathbf{x}_{i})$. Due to the fact that each node is a function of nodes from a previous layer, the nodes in adjacent layers are often visualized by connecting edges. Figure B.1 provides such a schematic overview of a general autoencoder neural network.
Figure B.1: Schematic overview of an autoencoder neural network with $M$ layers.
[…]
Note. The first and last layer are referred to as input and output layer, respectively. For participants $i=1,\dots,n$, the input in the first layer are the input responses $\mathbf{x}_{i}=(x_{i1},\dots,x_{ip})^{\top}$, and the output in the last layer are the activations $f_{j,\bm{\theta}}(\mathbf{x}_{i})=a_{ij}^{(M)},\ j=1,\dots,p,$ that reconstruct the input responses for a parameter vector $\bm{\theta}$ to be estimated. The $j$-th node in the $\ell$-th layer holds the activation $a_{ij}^{(\ell)}$, for $\ell=2,\dots,M,j=1,\dots,N^{(\ell)}$. Besides the parameter vector $\bm{\theta}$, the prediction function $\bm{f}_{\bm{\theta}}(\cdot)$ depends implicitly on the prespecified choice of the activation functions $g_{\ell}(\cdot)$.
To fit an autoencoder, we aim at minimizing its reconstruction error. For a prespecified loss function $\mathcal{L}(\cdot)$, we fit the neural network by finding the $\bm{\theta}$ that yields the best average reconstruction error,
$\widehat{\bm{\theta}}=\arg\min_{\bm{\theta}}\Bigg{\{}\frac{1}{n}\sum_{i=1}^{n}\sum_{j=1}^{p}\mathcal{L}\Big{(}x_{ij}-f_{j,\bm{\theta}}(\mathbf{x}_{i})\Big{)}\Bigg{\}}.$ (B.1)
The prediction function of the fitted network is subsequently given by $\bm{f}_{\widehat{\bm{\theta}}}(\cdot)$. For the loss function in (B.1), we choose the smooth and robust Pseudo-Huber loss, which is defined for a fixed $\delta>0$ as
$\mathcal{L}(z)=\delta^{2}\left(\sqrt{1+(z/\delta)^{2}}-1\right).$ (B.2)
This choice results in quadratic loss for small values of $z$ and linear loss for large values of $z$. Consequently, the Pseudo-Huber loss function avoids that large individual reconstruction errors strongly affect the fit, which would make it hard for the network to distinguish attentive responses from inconsistent careless responses, as the latter typically lead to large prediction errors. To solve the optimization problem in (B.1), we use stochastic gradient descent. We refer to Chapter 8.5 in Goodfellow et al., (2016) for details.
B.1.2 Modeling Choices
Following Kramer, (1992), we specify $M=5$ layers in the autoencoder. We set the number of nodes in layers 2 and 4 (mapping an demapping layer, respectively) to $\lfloor 1.5\times p\rfloor$, as a relatively large number of nodes is expected to give the autoencoder flexibility to learn many different types of response behavior (attentive and careless). Based on Assumption 1, the number of nodes in the bottleneck layer equals the number of constructs measured in the questionnaire.
Concerning the activation functions, we again follow a recommendation in Kramer, (1992) and propose to use nonlinear activation functions in the mapping and demapping layers, as well as a linear activation function in the bottleneck layer. Specifically, we propose to use the hyperbolic tangent activation in the (de-)mapping layers and the identity mapping in the bottleneck layer. Definitions are given in Table B.1, which summarizes our proposed autoencoder architecture.
Table B.1: Choices regarding the network architecture of the autoencoder.
Layer Number of nodes Activation function
Mapping layer $\lfloor 1.5\times p\rfloor$ $\textsf{tanh}(x)=2/(1+\exp(-2x))-1$
Bottleneck layer $s$ $\textsf{identity}(x)=x$
Demapping layer $\lfloor 1.5\times p\rfloor$ $\textsf{tanh}(x)=2/(1+\exp(-2x))-1$
Note. The constants $p$ and $s$ denote the number of items in the questionnaire and the number of constructs that the questionnaire measures, respectively. For a real number $x$, $\lfloor x\rfloor$ denotes the largest integer smaller than or equal to $x$.
For fitting the autoencoder, we use the stochastic gradient descent algorithm with a batch size of 10, a learning rate of 0.0001, and 100 epochs. In the presudo-Huber loss (B.2), we set constant $\delta=1$.
B.2 LongstringPattern Algorithm
Formal algorithms for computing the $l$-pattern and LSP sequences are provided in Algorithms 1 and 2, respectively.
Input: Series of observed rating-scale responses $\{x_{j}\}_{j=1}^{p}$, pattern length $l\in\mathbb{N}$ to consider (number of items that an individual pattern encompasses). For instance, the individual pattern $1-2-3$ encompasses $l=3$ items in the response sequence $-3-1-2-$ 3 - 1 - 2 - 3. Note that the algorithm measures straightlining when setting $l=1$.
Output: Sequence of measures $\{l\textsf{-pattern}_{j}\}_{j=1}^{p}$
Initialization: $streak\leftarrow 1,startindex\leftarrow 1$
1
/* fill in the first $l-1$ indices */
2 for $k\leftarrow 1$ to $p-l-1$ by $1$ do
3       if $x[k]==x[k+l]$ then
            increment $streak$ by 1 ;
             // streak is ongoing
4
5      else
             Assign value $streak$ to all elements in $\{l\textsf{-pattern}_{j}\}_{j=startindex}^{k}$ ;
             // streak broke
6             $startindex\leftarrow k+1$;
7             streak $\leftarrow 1$;
8
9       end if
10
11 end for
12
/* fill in the last $l$ indices (there are 3 edge cases) */
13 if $\texttt{streak}\geq l$ then
       /* case 1: current streak is ongoing and we have seen it completely at least once, so assume it will last until the end and complete the current streak (which has not yet been broken) */
14       Assign value $streak+l-1$ to all elements in $\{l\textsf{-pattern}_{j}\}_{j=startindex}^{p}$;
15
16 else if $1<\texttt{streak}<l$ then
       /* case 2: there is an ongoing streak, but we have no chance of seeing it completely (since streak < l), so we do not know how the streak looks like. In this case, fill in the current value of "streak" until the (p-l)-th item and fill in 1 for the last l items (because we do not know what the complete streak looks like) */
17       Assign value $streak-1$ to all elements in $\{l\textsf{-pattern}_{j}\}_{j=startindex}^{p-l}$;
18       Assign value $1$ to all elements in $\{l\textsf{-pattern}_{j}\}_{j=p-l+1}^{p}$;
19
20 else
       /* case 3: there is no ongoing streak, so fill in 1 for last $l$ indices */
21       Assign value 1 to all elements in $\{l\textsf{-pattern}_{j}\}_{j=p-l+1}^{p}$;
22
23 end if
Algorithm 1 $l$-pattern($\{x_{j}\}_{j=1}^{p},l$)
Input: Series of observed rating-scale responses $\{x_{j}\}_{j=1}^{p}$, maximum pattern length $l_{max}\in\mathbb{N}$ to consider
Output: Set of LongstringPattern measures $\{\textsf{LSP}_{j}\}_{j=1}^{p}$
1
/* calculate $l$-pattern for fixed pattern length $l$/
2 for $l\leftarrow 1$ to $l_{max}$ by $1$ do
3       $\{\textsf{LSP}^{l}_{j}\}_{j=1}^{p}\leftarrow l\textsf{-pattern}(\{x_{j}\}_{j=1}^{p},l)$;
4
5 end for
6
/* for each element $\{1,\dots,p\}$, report the largest $l$-pattern value */
7 Initialize $\{\textsf{LSP}_{j}\}_{j=1}^{p}$;
8 for $j\leftarrow 1$ to $p$ by $1$ do
9       $\textsf{LSP}_{j}\leftarrow\max\big{\{}\textsf{LSP}_{j}^{1},\textsf{LSP}_{j}^{2},\dots,\textsf{LSP}_{j}^{l}\big{\}}$;
10
11 end for
Algorithm 2 LongstringPattern($\{x_{j}\}_{j=1}^{p},l_{max}$)
B.3 Details on Changepoint Detection
Let $\{\bm{Y}_{j}\}_{j=1}^{p}$ be a series of length $p$ consisting of $d$-dimensional random variables. The goal of the cumulative sum self-normalization test of Shao & Zhang, (2010) is to estimate the location of a possible changepoint in the (multivariate) series, i.e., to divide the series into two subsets $\{\bm{Y}_{j}\}_{j=1}^{k-1}$ and $\{\bm{Y}_{j}\}_{j=k}^{p}$ in which a $d$-dimensional parameter of interest $\bm{\theta}$ differs. Here, we focus on the mean as the parameter $\bm{\theta}$. Define by $\widehat{\bm{\theta}}_{a,b}=(b-a+1)^{-1}\sum_{j=a}^{b}\bm{Y}_{j}$ the mean of the series calculated on the subset implied by periods $a\leq b$. For location $k\in\{2,\dots,p\}$, define the statistic
$T_{p}(k)=\bm{D}_{p}(k)^{\top}\bm{V}_{p}(k)^{-1}\bm{D}_{p}(k),$
where
$\bm{D}_{p}(k)$ $=\frac{(k-1)(p-k+1)}{p^{3/2}}\left(\widehat{\bm{\theta}}_{1,k-1}-\widehat{\bm{\theta}}_{k,p}\right),$
$\bm{V}_{p}(k)$ $=\bm{L}_{p}(k)+\bm{R}_{p}(k),$
with
$\bm{L}_{p}(k)=\left\{\begin{array}[]{ll}\displaystyle\sum_{i=1}^{k-2}\frac{i^{2}(k-1-i)^{2}}{p^{2}(k-1)^{2}}\left(\widehat{\bm{\theta}}_{1,i}-\widehat{\bm{\theta}}_{i+1,k-1}\right)\left(\widehat{\bm{\theta}}_{1,i}-\widehat{\bm{\theta}}_{i+1,k-1}\right)^{\top},&\qquad\text{if }k>2,\\ 0,&\qquad\text{if }k=2,\\ \end{array}\right.$
and
$\bm{R}_{p}(k)=\left\{\begin{array}[]{ll}\displaystyle\sum_{i=k+1}^{p}\frac{(p-i+1)^{2}(i-k)^{2}}{p^{2}(k-1)^{2}}\left(\widehat{\bm{\theta}}_{i,p}-\widehat{\bm{\theta}}_{k,i-1}\right)\left(\widehat{\bm{\theta}}_{i,p}-\widehat{\bm{\theta}}_{k,i-1}\right)^{\top},&\qquad\text{if }k<p,\\ 0,&\qquad\text{if }k=p,\\ \end{array}\right.$
The final test statistic is then defined as
$SN_{p}=\max_{k=2,\dots,p}T_{p}(k).$ (B.3)
Consider a given threshold $K_{d}>0$. If the test statistic exceeds the threshold, $SN_{p}>K_{d}$, an estimate of the changepoint location is obtained via
$\widehat{k}=\operatorname*{arg\,max}_{k=2,\dots,p}T_{p}(k).$
Otherwise, no changepoint is detected. Shao & Zhang, (2010) and Zhao et al., (2022) derive theoretical guarantees for this procedure. Specifically, Shao & Zhang, (2010) obtain the asymptotic distribution of the test statistic $SN_{p}$ for $p\rightarrow\infty$ and provide the corresponding critical values $K_{d}$ for a given significance level $\alpha$ and dimension $d$. We summarize relevant critical values in Table B.2.
Table B.2: Critical values $K_{d}$ for relevant significance levels $\alpha$ and dimensions $d$.
Dimension $d$
$\alpha$ 1 2
0.010 68.6 117.7
0.005 84.6 135.3
0.001 121.9 192.5
Note. The critical values are taken from Table 1 in Shao & Zhang, (2010).
In the context of this paper, we use the test statistic from (B.3) to test for changepoints in the two-dimensional series
$\bm{Y}_{j}=\begin{pmatrix}\textsf{RE}_{j}\\ \textsf{LSP}_{j}\end{pmatrix},\qquad j=1,\dots,p,$
where $\textsf{RE}_{j}$ is the autoencoder reconstruction error from (1) of a response to item $j$, and $\textsf{LSP}_{j}$ is the value of the LSP sequence from (2) assigned to this response. If a changepoint $\widehat{k}$ is detected, under Assumptions 4 and 5, the subsets $\{\bm{Y}_{j}\}_{j=1}^{\widehat{k}-1}$ and $\{\bm{Y}_{j}\}_{j=\widehat{k}}^{p}$ may be interpreted as attentive and careless responses, respectively. That is, the detected changepoint $\widehat{k}$ may be viewed as the onset of careless responding. We obtain such a series for each of the $n$ survey participants and subsequently test each of the $n$ series for changepoints.
On a final technical note, LSP sequences may exhibit low variation. Computing their variation over a relatively narrow interval—as required by the test statistic $SN_{p}$—may result in a singular (and therefore not invertible) variation matrix $\bm{V}_{p}(k)$ so that $SN_{p}$ cannot be computed. To avoid this issue, we by default inject Gaussian noise of tiny magnitude to each $\textsf{LSP}_{j},j=1,\dots,p$. Specifically, we inject draws from a normal distribution with mean zero and variance $0.01^{2}$. Doing so results in negligible but nonzero variation, which renders $\bm{V}_{p}(k)$ invertible.
Appendix C Details on the Empirical Application
C.1 Data Preprocessing
The data of Johnson, (2005), provided at https://osf.io/sxeq5/, contain the following information for each of the the $n=20,993$ respondents. First, per-item responses to $p=300$ IPIP-NEO-300 items, five domain scores (directly calculated from the responses), scores for psychometric antonyms and personal reliability, as well as respondent sex and age.
The responses in the provided data have already underwent some preprocessing that is described in detail the accompanying README file. Specifically, responses to negatively-worded items were reverse-coded, that is, a given response of 5 was recoded to 1, 4 was recoded to 3, and so on. Hence, in order to recover the original responses, we reverted this recoding. Furthermore, a total of 33,691 individual responses are missing (0.53% of all responses). Note that Johnson, had already removed participants with too many missing responses from the provided data set, hence there are relatively few missing responses for the remaining participants. Due to the low number of missing responses, we treat them as a separate answer category by coding them as 0.
C.2 Additional Results
Figure C.1 visualizes the detected onset items for the respondents that are flagged as careless by CODERS at significance levels 0.1%, 0.5%, and 1%.
Figure C.1: Estimated onset of careless responding for participants flagged by CODERS in the data of Johnson, (2005).
Note. Results for different significance levels $\alpha\in\{0.01,0.05,0.1\}$ are presented in separate panels. The $x$-axis shows the indices of the participants flagged as partially carelessness by CODERS. The participants are sorted in ascending order of the estimated carelessness onset item, which is in turn depicted on the $y$-axis. The dashed horizontal line represents the average location of carelessness onset across the flagged participants.
Figure C.2 provides a histogram as well as a density estimate for the per-respondent average reconstruction errors. We can see that the density has a long right tail, which might be due to careless respondents. We highlight the value 0.064 by a vertical dotted line, which is the average reconstruction error of the participant from Figure 3(d) and Table 2(d). This value is in the 99.7% percentile of all $n=20,993$ average reconstruction errors. Given the extremely high average reconstruction error and the fact that this respondent was flagged by both benchmark methods, it is likely that the participant responded carelessly throughout the survey, explaining why CODERS did not flag a changepoint.
Figure C.2: Histogram of average reconstruction errors for the data of Johnson, (2005).
Note. The $x$-axis contains the per-respondent average reconstruction errors, $\textsf{RE}_{i}=\frac{1}{p}\sum_{j=1}^{p}\textsf{RE}_{ij}$, $i=1,\dots,n$, where averages are taken over the $p=300$ items. On the $y$-axis, a histogram of the average reconstruction errors for $n=20,993$ respondents is displayed together with an empirical Gaussian kernel density. The vertical dotted line denotes the value of $\textsf{RE}_{i}$ for the participant in Figure 3(d) that was flagged by both benchmark methods, but not CODERS. This value, 0.064, lies in the 99.7% percentile of all average reconstruction errors.
Appendix D Details and Additional Simulations
D.1 Details on the Adjusted Rand Index
In the context of our simulations, the adjusted Rand index (ARI) for a given respondent is computed as
$ARI=\begin{cases}1&\text{ if the respondent is attentive and no changepoint isflagged},\\ \frac{A-BC/D}{(B+C)/2-BC/D}&\textnormal{ otherwise,}\end{cases}$
where
$A$ $=\binom{p_{11}}{2}+\binom{p_{00}}{2}+\binom{p_{10}}{2}+\binom{p_{01}}{2},$
$B$ $=\binom{p_{00}+p_{01}}{2}+\binom{p_{10}+p_{11}}{2},$
$C$ $=\binom{p_{00}+p_{10}}{2}+\binom{p_{01}+p_{11}}{2},$
$D$ $=\binom{p}{2},$
with $p_{11}$ denoting the number of careless responses that are correctly identified as careless, $p_{00}$ the number of attentive responses that are correctly identified as attentive, $p_{10}$ the number of careless responses that are incorrectly identified as attentive, and $p_{01}$ the number of attentive responses that are incorrectly identified as careless. Intuitively, $p_{11}$ and $p_{00}$ measure correct classification, whereas $p_{01}$ and $p_{10}$ measure misclassification. Observe that the ARI assumes value 1 if there is no misclassification, which is the maximum value it can take. Hence, for accurate estimates of the carelessness onset item, the ARI will be close to value 1. For very inaccurate estimates of carelessness onset, the ARI will be close to 0, or potentially even below that. It is unclear how to interpret negative ARI values, but such values only occur rarely. We refer to Steinley, (2004) for a detailed discussion of the properties of the ARI.
If a changepoint is flagged in an attentive respondent, the ARI takes value 0. Similarly, if no changepoint is flagged in an attentive respondent, the ARI takes value 1. Hence, when averaging the ARI over attentive respondents, the ensuing average can be interpreted as the proportion of attentive respondents who are correctly identified as attentive. On the other hand, averages of ARIs calculated on partially careless respondents do not have a direct interpretation other than higher values indicating better performance.
D.2 Additional Simulation Results
We investigate the effect of various choices in our simulations. As the results are similar to those in the main paper other than some shifts in size or power, we only present plots of the average ARI for attentive and careless respondents (similar to Figure 5) in order to keep the appendix concise.
D.2.1 Choosing a Different Significance Level
Figure D.1: ARI of CODERS at a significance level of 0.5%.
Note. The ARI is averaged over fully attentive respondents and partially careless respondents (columns) for different carelessness onset settings (rows).
Figure D.2: ARI of CODERS at a significance level of 1%.
Note. The ARI is averaged over fully attentive respondents and partially careless respondents (columns) for different carelessness onset settings (rows).
We use the baseline simulation design from the main text, but we set the significance level in CODERS to 0.5% and to 1%, respectively. The corresponding results are shown in Figures D.1 and D.2. Compared to the significance level of 0.1% in the main text (cf. Figure 5), there is a gradual increase in power in the changepoint detection, yielding a small increase in the ARI across prevalence levels for careless respondents. Further investigation suggests that the increase in ARI stems mainly from better performance in detecting random and extreme responding (results not shown to keep the appendix at a reasonable length). However, for high prevalence levels from 40% onwards, this comes at a cost of too many false positives in attentive respondents, translating to a substantial decrease in ARI for the highest prevalence levels.
D.2.2 Varying Correlation Structures
In this variation of the simulation design, we keep the same settings as in the baseline design from the main text, but we vary the correlation structure in the following ways:
  1. 1.
    Randomly sampled moderate correlations. In every repetition, correlations of item pairs from the same scale are randomly drawn from the interval $[0.4,0.6]$. Items from different scales remain uncorrelated.
  2. 2.
    Randomly sampled strong correlations. In every repetition, correlations of item pairs from the same scale are randomly drawn from the interval $[0.6,0.8]$. Items from different scales remain uncorrelated.
  3. 3.
    Randomly sampled grouped correlation structure. Multiple scales form groups so that there are 5 groups consisting of 6 scales each. Since there are 10 items in each scale, there are $6\times 10=60$ items in each group. In each repetition, correlations of item pairs from the same scale are randomly drawn from the interval $[0.6,0.8]$, correlations of item pairs in different scales of the same group are randomly drawn from the interval $[0,0.2]$, and items from different groups are uncorrelated.
While the baseline simulation design uses a relatively simple correlation structure (items from the same scale have a fixed correlation of 0.7 before recoding, items from different scales are uncorrelated), the three variations above introduce more randomness into the design. Due to this randomness, it may occur that a sampled correlation matrix is not positive definite. In such a case, we find the closest positive definite correlation matrix via the function nearPD() in the R package Matrix (Bates et al.,, 2023).
Figure D.3: ARI of CODERS for randomly sampled moderate correlations.
Note. The ARI is averaged over fully attentive respondents and partially careless respondents (columns) for different carelessness onset settings (rows).
Figure D.4: ARI of CODERS for randomly sampled strong correlations.
Note. The ARI is averaged over fully attentive respondents and partially careless respondents (columns) for different carelessness onset settings (rows).
Figure D.5: ARI of CODERS for the randomly sampled grouped correlation structure.
Note. The ARI is averaged over fully attentive respondents and partially careless respondents (columns) for different carelessness onset settings (rows).
Figures D.3D.5 display the results corresponding to the different correlation structures at significance level 0.1%. The results of the design with randomly sampled strong correlations (Figure D.4) and the grouped correlation structure (Figure D.5) are essentially equivalent to the results of the baseline design with fixed correlation (Figure 5 in the main text). The design with randomly sampled moderate correlations (Figure D.3) also yields similar results, save for a small power drop in the changepoint detection (resulting in a lower ARI for careless respondents) and fewer false positives in extreme carelessness prevalence levels (yielding a higher ARI in attentive respondents).
D.2.3 Varying Sample Size
This variation of the simulation design again uses the same settings as in the baseline design from the main text, but now we change the sample size to $n=250$ and $n=1000$, respectively.
Figures D.6 and D.7 visualize the corresponding results at significance level 0.1%. They are similar to the results for $n=500$ in the main text (cf. Figure 5), other than a small decrease ARI in careless respondents for $n=250$ as well as a small increase for $n=1000$.
Figure D.6: ARI of CODERS for sample size $n=250$.
Note. The ARI is averaged over fully attentive respondents and partially careless respondents (columns) for different carelessness onset settings (rows).
Figure D.7: ARI of CODERS for sample size $n=1000$.
Note. The ARI is averaged over fully attentive respondents and partially careless respondents (columns) for different carelessness onset settings (rows).

### v3

@@
 Abstract
-Questionnaires in the behavioral and organizational sciences tend to be lengthy: survey measures comprising hundreds of items are the norm rather than the exception.
-However, literature suggests that the longer a questionnaire takes, the higher the probability that participants lose interest and start responding carelessly.
+Questionnaires in the behavioral and organizational sciences tend to be lengthy.
+However, literature suggests that survey length is a contributing factor to careless responding, with longer questionnaires yielding higher probability that participants start responding carelessly.
 Consequently, in long surveys a large number of participants may engage in careless responding, posing a major threat to internal validity.
-We propose a novel method for identifying the onset of careless responding (or an absence thereof) for each participant.
-It is based on combined measurements of multiple dimensions in which carelessness may manifest, such as inconsistency and invariability.
-Since a structural break in either dimension is potentially indicative of carelessness, the proposed method searches for evidence for changepoints along the combined measurements.
-It is highly flexible, based on machine learning, and provides statistical guarantees on its performance.
-An empirical application on data from a seminal study on the incidence of careless responding reveals that the reported incidence has likely been substantially underestimated due to the presence of respondents that were careless for only parts of the questionnaire.
-In simulation experiments, we find that the proposed method achieves high reliability in correctly identifying carelessness onset, discriminates well between careless and attentive respondents, and captures a variety of careless response types, even when a large number of careless respondents are present.
-Furthermore, we provide freely available open source software to enhance accessibility and facilitate adoption by empirical researchers.
+We propose a novel method for identifying the onset of careless responding (or an absence thereof) that searches for a changepoint in combined measurements of multiple dimensions in which carelessness may manifest, such as inconsistency and invariability.
+It is highly flexible, based on machine learning, and provides statistical guarantees for controlling the false positive rate.
+In simulation experiments, the proposed method achieves high accuracy in identifying carelessness onset and discriminates well between attentive and various types of careless responding, even when a large number of careless respondents are present.
+An empirical application highlights how identifying partial carelessness uncovers novel insights on careless responding behavior.
+Furthermore, we provide the freely available open source software package carelessonset to facilitate adoption by empirical researchers.
 Keywords: careless responding, changepoint detection, machine learning, survey methodology
@@
 Careless responding has been identified as a major threat to the validity of research findings through a variety of psychometric and statistical issues such as lower scale reliability, errors in statistical hypothesis testing, and deteriorated model fit (e.g., Arias et al.,, 2020; Huang et al.,, 2012, 2015; Kam & Meyer,, 2015).
-It is widely prevalent (e.g., Bowling et al.,, 2016; DeSimone et al.,, 2015; Curran,, 2016; Meade & Craig,, 2012), suspected to be present in all questionnaire data (Ward & Meade,, 2023), and already incidence rates as low as 5% can be problematic (Arias et al.,, 2020; Credé,, 2010).
+It is widely prevalent (e.g., Bowling et al.,, 2016; Curran,, 2016; DeSimone et al.,, 2015; Meade & Craig,, 2012), suspected to be present in all questionnaire data (Ward & Meade,, 2023), and already incidence rates as low as 5% can be problematic (Arias et al.,, 2020; Credé,, 2010; Welz et al., 2024a, ; Welz et al., 2024b, ).
 Often viewed as merely an insidious source of measurement error, careless responding can be a valuable source of information about study participants and the study itself.
@@
 Consequently, it is widely recommended to carefully screen survey data for the presence of careless responding (e.g., DeSimone et al.,, 2015; Huang et al.,, 2015).
-Numerous methods for identifying careless respondents after questionnaire administration (so-called screeners) have been proposed, for instance consistency indicators such as psychometric synonyms (Meade & Craig,, 2012), longstring indices (Johnson,, 2005), or multivariate outlier analyses (e.g., Curran,, 2016).
+Numerous methods for identifying careless respondents after questionnaire administration (so-called screeners) have been proposed, for instance inconsistency indicators such as psychometric synonyms or antonyms (Meade & Craig,, 2012), invariability indicators such as the longstring index (Johnson,, 2005), or multivariate outlier analyses (e.g., Curran,, 2016).
 We refer to Arthur et al., (2021), DeSimone et al., (2015), and Ward & Meade, (2023) for detailed reviews.22endnote: 2A popular pre-administration method for the detection of careless responding is the inclusion of detection items (e.g., Table 1 in Arthur et al.,, 2021).
@@
 For instance, it is expected that an attentive respondent would strongly disagree to so-called bogus items such as “I am paid biweekly by leprechauns” (Meade & Craig,, 2012).
-A careless respondent may accidentally “agree” to this item as a consequence of inattention.
 Alternative types of detection items are instructed items and self-report items (Arthur et al.,, 2021).
-We discuss detection items and preventive measures against carelessness in detail in a later section.
-More recently, Schroeders et al., (2022), Alfons & Welz, (2024), and Ozaki, (2024) have studied machine learning techniques for identifying careless respondents.
-Still, commonly used screeners tend to suffer from three major limitations, which we discuss in the following.
-First, extant methods commonly follow the implicit assumption that careless respondents respond carelessly to all questionnaire items or at least a vast majority of items.
-In particular, they indicate whether a respondent engaged in careless responding, but they typically do not provide any indication regarding which items are affected.
-Although there exist many comparative studies (e.g., Curran,, 2016; DeSimone & Harms,, 2018; Goldammer et al.,, 2020; Huang et al.,, 2012) on such screeners, their behavior under partial careless responding is not well-understood.
-This is at odds with the substantial evidence that many careless respondents tend to begin a survey as attentive and truthful respondents, but may resort to careless responding due to fatigue or boredom as a survey progresses (Bowling et al.,, 2021; Brower,, 2018; Galesic & Bosnjak,, 2009; Gibson & Bowling,, 2020; Ward et al.,, 2017).
-In general, Ward & Meade, (2023) believe it is rare that careless respondents respond carelessly from the beginning to the end of a questionnaire.
-Second, common detection methods are typically designed to detect a single specific type of carelessness.
-For instance, the longstring index of Johnson, (2005) counts the maximum number of consecutive identical responses, which is intended to capture straightlining behavior.
-However, careless responding may manifest in multiple distinctive ways (e.g., Curran,, 2016; DeSimone et al.,, 2015; DeSimone & Harms,, 2018; Edwards,, 2019; Ward & Meade,, 2023), implying that a screener that is designed to detect one type of carelessness may fail to detect other types.
-It is therefore recommended to combine multiple screeners to detect various types of careless respondents (e.g., Goldammer et al.,, 2020; Huang et al.,, 2012; Meade & Craig,, 2012; Ward & Meade,, 2023).
-Third, cutoff values for flagging careless respondents are often ad-hoc and lack statistical foundation.
-Screeners typically calculate a score for each participant, with a participant being flagged as careless if their score is sufficiently extreme (i.e., beyond a certain cutoff value).
-Ideally, such a cutoff value results in high sensitivity (flagging true careless respondents as careless) and high specificity (not flagging true attentive respondents as careless).
-For certain screeners, various cutoff values have been suggested in the literature, but individual suggestions vary substantially.
-For instance, for the same screeners, Goldammer et al., (2020) argues for different cutoff values than Curran, (2016) or Johnson, (2005), which suggests that the results for specific cutoff values may depend on factors such as the population or the constructs being measured (cf.
-Meade & Craig,, 2012, p.
-452).
-For a specific screener, it would be desirable to obtain a cutoff value with statistical guarantees on sensitivity and specificity that does not depend on population or study characteristics.
-In wake of these three issues with extant methods, we propose a novel screener for the identification of careless responding that is designed to overcome these issues.
-We achieve this by focusing on a more nuanced question.
-Motivated by the prominence of partial careless responding due to survey fatigue or boredom, we focus on identifying at which point in the survey a given participant becomes careless (if at all).
-Specifically, our method estimates the item at which carelessness onsets in each questionnaire participant, or an absence thereof.
-Hence, we call our method careless onset detection in extensive rating-scale surveys (CODERS).
+More recently, Alfons & Welz, (2024), Ozaki, (2024), and Schroeders et al., (2022) have studied machine learning techniques for identifying careless respondents.
+Existing methods and their implementations in freely available software packages such as careless (Yentes & Wilhelm,, 2023) make screening for careless respondents a simple task for empirical researchers, thus elevating reliability and validity of research findings and playing an instrumental role in gaining insights into the phenomenon of careless responding.
+In addition, many popular screeners are intuitive and easy to understand.
+It should be noted that common indicators typically compute a single score for each participant, with respondents deemed careless and removed from subsequent analysis if their score exceeds a certain cutoff value.
+Although there exist many comparative studies on such screeners (e.g., Curran,, 2016; DeSimone & Harms,, 2018; Goldammer et al.,, 2020; Huang et al.,, 2012), their behavior under partially careless responding is unfortunately not well-understood.
+Furthermore, excluding partially careless respondents—particularly in lengthy surveys—may leave a lot of valuable information unused, namely the respective parts of the survey in which they responded attentively.
+We therefore aim to address a more nuanced question: Can we identify at which point in the survey participants start responding carelessly, if at all?
+Indeed, there is evidence that many—though not all—careless respondents tend to begin a survey as attentive and truthful respondents, but may resort to careless responding as the survey progresses, for instance due to fatigue or boredom (Bowling et al.,, 2021; Brower,, 2018; Galesic & Bosnjak,, 2009; Gibson & Bowling,, 2020; Ward et al.,, 2017); cf.
+also the literature on back random responding (e.g., Clark et al.,, 2003) or partial random responding (e.g.
+Berry et al.,, 1992; Gallen & Berry,, 1997).
+Successful identification of the onset of careless responding can therefore yield valuable, complementary insights to those from existing screeners.
+For subsequent analysis, the issue of partially careless responding may be transformed into a missing data problem.
+Rather than excluding partially careless respondents altogether, researchers can restrict their primary analyses to the parts of attentive responses, treating the careless responses as missing values.
+Common guidelines for handling missing data (e.g., Newman,, 2014) can thereby be followed.
+In addition, accurately identifying the onset of careless responding may uncover implications for survey design (e.g., if many participants start responding carelessly at a similar point in the survey, it may be beneficial to rephrase or replace those items, or to place an intervention) and may lead to new theory and a deeper understanding of careless responding behavior (e.g., potentially linking certain personality traits to a tendency for an earlier or later onset of careless responding; cf.
+Gibbert et al.,, 2021, for theory-building based on outlying data).
+For the purpose of identifying at which point in the survey a given participant becomes careless, if at all, we introduce a novel method called careless onset detection in extensive rating-scale surveys (CODERS).
+Specifically, CODERS estimates the item at which carelessness onsets in each questionnaire participant, or an absence thereof.
+An important issue is that careless responding may manifest in multiple distinctive ways (e.g., Curran,, 2016; DeSimone et al.,, 2015; DeSimone & Harms,, 2018; Edwards,, 2019; Ward & Meade,, 2023).
 To capture various types of carelessness, CODERS can combine multiple dimensions of evidence for careless responding to construct a score that, for each item, measures if a given respondent has started responding carelessly by that item.
-It achieves this by searching for changepoints along a multidimensional series of indices that are potentially indicative of carelessness, such as measures of response consistency or variability.
-We argue that the notion of a changepoint is intuitive when studying partial carelessness: Once a participant starts responding carelessly, we expect a structural break in their responses.
-In particular, such a respondent may abandon content-based responding and resort to careless response types, while no such break occurs in the absence of carelessness.
+It achieves this by searching for changepoints along a multidimensional series of indices that are potentially indicative of carelessness, such as measures of response inconsistency or invariability.
+We argue that the notion of a changepoint is intuitive when studying partial carelessness: if a participant starts responding carelessly, we expect a structural break in their responses.
+In particular, such a respondent may abandon content-based responding and resort to careless response types from the changepoint onward.
 CODERS is highly flexible as it does not assume a statistical model, nor does it require predefining what types of careless response types may exist, and it is primarily intended for lengthy multi-scale surveys.
-Moreover, CODERS builds on nonparametric estimation theory for cutoff values and statistical guarantees regarding sensitivity and specificity.
-Re-examining data from a a seminal study on careless responding in Big 5 measurements (Johnson,, 2005), we demonstrate novel insights that can be obtained by estimating onset items of partially careless responding.
-Most notably, we find strong evidence that Johnson, (2005) considerably underestimates the incidence of careless respondents due to hitherto undetected presence of partially careless respondents.
-Through extensive simulation experiments, we find that the proposed method CODERS accurately estimates carelessness onset (or a lack thereof) and thereby achieves excellent sensitivity and specificity in detecting careless responding.
+Moreover, CODERS builds on nonparametric estimation theory for cutoff values with statistical guarantees regarding the false positive rate (i.e., the probability of attentive respondents being incorrectly flagged as partially careless).
 To the best of our knowledge, CODERS is the first attempt to systematically detect the onset of careless responding.
-Assuming that participants respond accurately in the beginning of the survey, CODERS is designed to segment the responses of each respondent into a segment of accurate responses and—if a changepoint was identified—a segment of careless responses.
-With this knowledge, researchers can restrict their primary analyses to the segments of accurate responses (by treating the careless responses as missing values) without having to discard all responses of a partially careless respondent.
-In addition, researchers can separately study the segments of careless responses to build theory on and obtain better understanding of the nature of (partially) careless responding (cf.
-Gibbert et al.,, 2021, for theory-building based on outlying data points).
-Finally, we provide the freely available software package carelessonset (Welz & Alfons,, 2024), which implements CODERS in the statistical programming environment R (R Core Team,, 2024) can be downloaded from https://github.com/mwelz/carelessonset.
+Through extensive simulation experiments, we find that CODERS accurately estimates carelessness onset (or a lack thereof) and achieves excellent sensitivity and specificity in detecting partially careless responding.
+Crucially, CODERS should not be viewed as a replacement of existing screeners for careless respondents, but as a complementary tool for a related but more nuanced task, as highlighted by an empirical application.
+Re-examining data from a a seminal study on careless responding in Big Five personality measurements (Johnson,, 2005), we demonstrate novel insights that can be obtained by estimating onset items of partially careless responding.
+Reflecting on those insights, we subsequently discuss limitations and potential extensions of CODERS.
+Our concluding discussion highlights how CODERS complements various existing approaches for the detection of careless respondents.
+Finally, we provide the freely available software package carelessonset (Welz & Alfons,, 2024), which implements CODERS in the statistical programming environment R (R Core Team,, 2024) and can be downloaded from https://github.com/mwelz/carelessonset.
 As such, it is a useful and accessible tool for any researcher who is concerned with survey fatigue or careless responding in general.
@@
 The literature generally identifies three distinct ways in which careless responding may manifest: low internal consistency of the given responses (inconsistency), low variability of given responses (invariability), and impossibly fast responses (e.g., Curran,, 2016; DeSimone et al.,, 2015; DeSimone & Harms,, 2018; Edwards,, 2019; Ward & Meade,, 2023).
-We expect partial careless responding to result in a change in respondent behavior in at least one of these dimensions.
+We expect partially careless responding to result in a change in respondent behavior in at least one of these dimensions.
 CODERS therefore offers a general framework for searching for a joint changepoint in any combination of measurements of these dimensions of carelessness.
 While the majority of this paper is focused on inconsistency and invariability, we revisit response times in a later section.
-We propose to measure the first dimension, inconsistency, by means of reconstructions of observed responses (cf.
+As used here, CODERS thus consists of three main components:
+  1. 1.
+    For each participant, a nonparametric changepoint detection test (Shao & Zhang,, 2010) is applied to a combined series of participant-item level measurements of careless responding.
+  2. 2.
+    As a measurement of inconsistency, reconstruction errors (RE) obtained via an auto-associative neural network (autoencoder) (Kramer,, 1992) are used.
+An autoencoder can be viewed as a nonlinear generalization of principal component analysis (PCA), and the (participant-item level) reconstruction errors can be interpreted as the difference between observed responses and what is expected under attentive responding.
+  3. 3.
+    As a measurement of invariability, we propose a novel algorithm called LongStringPattern (LSP).
+LSP can be viewed as a generalization of the famous longstring index of Johnson, (2005) to the participant-item level, also allowing for recurring patterns involving multiple answer categories.
+Regarding the first dimension, inconsistency, the autoencoder learns response patterns that characterize attentive responding (cf.
 Alfons & Welz,, 2024).
-These reconstructions are generated by an auto-associative neural network (henceforth autoencoder) (Kramer,, 1992), which learns response patterns that characterize attentive responding.
 We expect that random-like, content-independent responses cannot be learned well and are thus poorly reconstructed by the autoencoder, leading to a changepoint in reconstruction performance.
-For measuring the second dimension, invariability, we propose a novel algorithm that is inspired by the famous longstring index of Johnson, (2005).
-Since long sequences of identical responses or constant response patterns are not expected in surveys that use positively and negatively coded items in a randomized order, we expect a changepoint in response variability once a respondent commences to respond carelessly through straightlining or pattern responding behavior.
+Regarding the second dimension, invariability, long sequences of identical responses or fixed response patterns are not expected in well-designed surveys (e.g., using positively and negatively coded items, randomized in some form).
+Hence, we expect a changepoint in response variability once a respondent commences to respond carelessly through straightlining or pattern responding behavior.
+To obtain a granular estimate of carelessness onset, both reconstruction errors and LSP sequences are measured on the participant-item level so that a changepoint can be detected on the item-level for each participant.
 Overall, CODERS attempts to capture evidence for the onset of carelessness by combining multiple indicators that are potentially indicative of such an onset, where the different indicators may capture different manifestations of carelessness.
@@
 Combining multiple indicators is a generally recommended practice to capture various types of careless responding (Goldammer et al.,, 2020; Huang et al.,, 2012; Meade & Craig,, 2012; Ward & Meade,, 2023).
-We provide a detailed description of our assumptions in Appendix A and a technical definition of CODERS in Appendix B.
+We provide a detailed description of our assumptions in Appendix A and additional details on CODERS in Appendix B.
 Quantifying Inconsistency with Autoencoders
@@
 Since we consider the defining characteristic of inconsistent careless responding to be near-randomly chosen responses—which may be viewed as random noise in the context of questionnaire data—we expect autoencoders to perform well in distinguishing between attentive responses and inconsistent careless responses.
-To achieve the goal of learning the internal structures of a dataset, it is crucial to avoid that an autoencoder simply passes the input data through its network architecture.
-Consequently, autoencoders first express the input data in terms of a low-dimensional representation before reconstructing the input data again from this low-dimensional representation (cf.
+To achieve the goal of learning the internal structures of a dataset, autoencoders first express the input data in terms of a low-dimensional representation before reconstructing the input data again from this low-dimensional representation (cf.
 Kramer,, 1992, and Chapter 14.6 in Goodfellow et al.,, 2016).
-The use of such a low-dimensional representation makes autoencoders appealing for questionnaire data, as questionnaires typically contain multiple items to measure each construct of interest.
+The use of such a low-dimensional representation makes autoencoders appealing for questionnaire data, as questionnaires typically contain multiple items to measure each construct of interest, such as attitudes, sentiments, or personality traits.
 In this context, the autoencoder aims to learn the patterns between the latent constructs in the low-dimensional space.
@@
 Inconsistent careless responses, on the other hand, should contain little to no structure also in this low-dimensional space, and are therefore expected to be reconstructed poorly.
-A schematic example of the network architecture of an autoencoder is given in Figure 1.
-Concretely, an autoencoder is a fully connected network whose nodes are organized in multiple layers.
-Each node takes a linear combination of the nodes from the previous layer and subsequently applies a so-called activation function in order to allow for nonlinear relationships.
-A crucial part of an autoencoder is the central layer, known as the bottleneck layer, which stores the low-dimensional representations of the input data.
-The network architecture of an autoencoder is symmetric around this bottleneck layer such that the can be viewed as having an encoder-decoder architecture: the encoder compresses the data into the low-dimensional representations, while the decoder reconstructs the input data again.
-It follows that an autoencoder can be seen as a dimension reduction technique, specifically as a nonlinear generalization of principal component analysis (PCA; see Kramer,, 1991).
 Figure 1: Schematic example of a neural network with an autoencoder structure.
@@
 The second layer (mapping layer) prepares the data to be compressed in the third and central layer (bottleneck layer), which holds low-dimensional representations of the input data (here $\bm{\widehat{z}}_{1}$ and $\bm{\widehat{z}}_{2}$).
-The fourth layer (demapping layer) prepares the low-dimensional respresentations for reconstructing the input data in the final layer (output layer), yielding reconstructions in the original dimension (here $\bm{\widehat{x}}_{1},\dots,\bm{\widehat{x}}_{4}$).
+The fourth layer (demapping layer) prepares the low-dimensional representations for reconstructing the input data in the final layer (output layer), yielding reconstructions in the original dimension (here $\bm{\widehat{x}}_{1},\dots,\bm{\widehat{x}}_{4}$).
+A schematic example of the network architecture of an autoencoder is given in Figure 1.
+Concretely, an autoencoder is a fully connected network whose nodes are organized in multiple layers.
+Each node takes a linear combination of the nodes from the previous layer and subsequently applies a so-called activation function in order to allow for nonlinear relationships.
+A crucial part of an autoencoder is the central layer, known as the bottleneck layer, which stores the low-dimensional representations of the input data.
+The network architecture of an autoencoder is symmetric around this bottleneck layer such that it can be viewed as having an encoder-decoder architecture: the encoder compresses the data into the low-dimensional representations, while the decoder reconstructs the input data again.
+It follows that an autoencoder can be seen as a dimension reduction technique, specifically as a nonlinear generalization of principal component analysis (PCA; see Kramer,, 1991).
 Hence, the number of nodes in the bottleneck layer is an important design choice in the architecture of an autoencoder.
@@
 Thus, this instrument measures a total of 30 facets, such as anxiety and modesty, so that we would set the number of nodes in the bottleneck layer to 30.
-A thorough mathematical definition of autoencoders can be found in Appendix B.1.
-Although there are other design choices besides the number of nodes in the bottleneck layer (see Appendix B.1.2), our software provides default values that we recommend researchers to use.
+Further intuition and a thorough mathematical definition of autoencoders can be found in Appendix B.2.
+Although there are other design choices besides the number of nodes in the bottleneck layer (see Appendix B.2.3), our software implementation in package carelessonset (Welz & Alfons,, 2024) provides default values that we recommend researchers to use.
 The only input required by the user is the number of constructs measured in the survey (for the number of nodes in the bottleneck layer).
@@
 Hence, if participant $i$ starts providing inconsistent careless responses at item $k\in\{1,\dots,p\}$, we expect a changepoint at position $k$ in the participant’s series of $p$ reconstruction errors, as reconstruction errors are expected to be higher from item $k$ onward.
-An example can be found in the top panel of Figure 2(a).
-Figure 2: Illustrative example output from CODERS using responses of a participant to a 300-item personality inventory.
-(a) Two dimensions indicative of carelessness: autoencoder reconstruction errors (RE) in the top panel, and the LongstringPattern (LSP) sequence in the bottom panel.
-(b) The per-response statistic of the changepoint detection test. The dashed horizontal line corresponds to the associated critical value at an extremely conservative significance level of 0.1%.
-Note.
-The solid vertical lines indicate the identified carelessness onset, which is at the 148th item.
-Reconstruction errors are in general much larger after this item than before, indicating more random-like responses.
-The response data are taken from Johnson, (2005) and are discussed in detail in a later section.
+An example can be found in Figure 2, where we show simulated responses of a hypothetical participant in a 300-item questionnaire (5 answer categories per item) who starts choosing answer categories completely at random from the 172nd item, resulting in systemically larger autoencoder reconstruction errors from that item onward.
+Figure 2: Simulated illustrative example of autoencoder reconstruction errors of an inconsistent partially careless respondent to a 300-item questionnaire comprising five-point Likert scales.
+Note. The simulated participant started choosing answer categories completely at random from the 172nd item onward.
 Quantifying Invariability
@@
 In short, higher values in the $l$-pattern sequence capture a tendency to respond in fixed patterns of length $l$.
-If a participant starts engaging in invariable careless responding, we expect a changepoint in the $l$-pattern sequence (for some $l$) from relatively low to relatively high values, assuming that the order of items is randomized.
+If a participant starts engaging in invariable careless responding, we expect a changepoint in the $l$-pattern sequence (for some $l$) from relatively low to relatively high values, assuming that the survey design precludes such patterns in attentive responses.
 However, an $l$-pattern sequence crucially depends on the choice of pattern length $l$, which is restrictive since invariable careless respondents may choose widely different patterns.
@@
 for participant $i=1,\dots,n$ and item $j=1,\dots,p$.
+We call the series of within-participant values $\textsf{LSP}_{ij}$ across all items a LongStringPattern (LSP) sequence.
 In this paper, we vary $l$ from 1 to the number of answer categories in the survey.
-The rationale behind this choice is that we consider it unlikely that careless respondents choose complicated individual patterns whose length exceeds the number of answer categories.
-We provide a detailed description of the LSP algorithm in Appendix B.2. An example of an LSP sequence is shown in the bottom panel of Figure 2(a).
+The rationale behind this choice is that a careless respondent may repeat, e.g., the pattern $1-2-3-4-5$ in case of five answer categories.
+At the same time, we consider it unlikely that careless respondents choose complicated individual patterns whose length exceeds the number of answer categories.
+We provide a detailed description of the LSP algorithm in Appendix B.3.
+An example of an LSP sequence is shown in Figure 3, which shows simulated responses of a hypothetical participant in a 300-item questionnaire (5 answer categories per item) who starts responding carelessly using the recurring pattern $3-4$ from the 260th item, resulting in systematically larger LSP values from that item onward.
+Figure 3: Simulated illustrative example of a LongStringPattern sequence of an invariable partially careless respondent to a 300-item questionnaire comprising five-point Likert scales.
+Note. The simulated participant started responding carelessly with the recurring pattern $3-4$ from the 260th item onward.
 Identifying Carelessness Onset via Changepoint Detection
@@
 Each individual series should thereby measure one of the dimensions indicative of carelessness.
-Although we focus on using reconstruction errors from (1) to measure inconsistency as well as the LSP sequence from (2) to measure invariability, it should be noted that in principle CODERS can also be used with any single measurement or any other combination of different measurements of carelessness.
-As motivated in the previous two sections, we expect a changepoint in any of the measurements of carelessness from the onset of careless responding (e.g., Figure 2(a)), and no changepoint in the absence of carelessness.
+Although we focus on using reconstruction errors from (1) to measure inconsistency as well as the LSP sequence from (2) to measure invariability, in principle CODERS can also be used with any single measurement or any other combination of different measurements of carelessness on the participant-item level.
+As motivated in the previous two sections, we expect a changepoint in any of the measurements of carelessness from the onset of careless responding (cf.
+Figures 2 and 3), and no changepoint in the absence of carelessness.
 We therefore propose to apply the nonparametric cumulative sum self-normalization test of Shao & Zhang, (2010), which was developed in the context of time series analysis.
@@
 Otherwise no changepoint is flagged.
-The critical value thereby depends only on the significance level $\alpha$ and the dimension of the multivariate series.
+The critical value thereby depends only on a pre-specified significance level $\alpha$ and the dimension of the multivariate series.
 Importantly, it controls the size of the test regardless of the distribution of the individual series.
@@
 In line with literature that suggests that there should be overwhelming evidence when labeling respondents as careless (e.g., Huang et al.,, 2012), we recommend using a conservative significance level no greater than 1% for identifying changepoints.
-As an example, consider again the two-dimensional series in Figure 2(a), which measures carelessness of a certain survey participant in terms of autoencoder reconstruction errors (RE) and the LongstringPattern (LSP) sequence.
-Figure 2(b) shows the corresponding series of statistics, where the maximum value at the 148th item clearly exceeds the critical value.
-That is, CODERS flags this respondent as being careless from the 148th item onwards.
-Indeed, from roughly this item onwards, the reconstruction errors are in general much larger than before.
-It follows that this respondent likely is a partially inconsistent careless respondent.
+As an illustrative example, consider Figure 4(a).
+It illustrates two dimensions computed from a given participant’s responses, namely autoencoder reconstruction errors and the LSP sequence.
+The former series has been shown before in Figure 2.
+The bottom panel of Figure 4(a) shows the test statistics computed from this two-dimensional series, where the maximum value at the 172nd item clearly exceeds the critical value.
+That is, CODERS identifies inconsistent carelessness from the 172nd item onward, as the reconstruction errors are in general much larger than before.
+Indeed, this simulated participant started choosing responses at random at the identified item.
+Figure 4(b) is similar, but computed from a different respondent, whose LSP sequence has been shown before in Figure 3.
+CODERS flags a changepoint at the 261st item.
+From this item onward, the LSP values are clearly elevated, indicating invariable carelessness.
+In fact, this simulated participant started repeating the pattern $3-4$ just one item before the estimated onset.
+Figure 4: Illustrative example output from CODERS using responses of two simulated partially careless respondents to a 300-item questionnaire.
+(a) Inconsistent partial carelessness.
+(b) Invariable partial carelessness.
+Note.
+The left subfigure shows the output from CODERS for a simulated partially inconsistent respondent who starts responding randomly from the 172nd item onward, while the right subfigure does so for a simulated partially invariable respondent who starts repeating the pattern $3-4$ from the 260th item onward.
+Each subfigure visualizes the autoencoder reconstruction errors (RE), LongStringPattern (LSP) sequence, and test statistics associated with the corresponding participant’s responses.
+The blue solid vertical lines indicate the carelessness onset items identified by CODERS at significance level 0.1%, which correspond to the true onset item 172 for the partially inconsistent respondent, and to item 261 (one item after true onset) for the partially invariable respondent.
+The dashed horizontal line is the critical value of the test statistic at the 0.1% significance level (192.5; see Table B.2 in Appendix B.4).
+The reconstruction errors in the left subfigure are those shown in Figure 2, and the LSP sequence in the right subfigure is that shown in Figure 3.
+Simulation Experiments
+After introducing CODERS, we now investigate its behavior via simulation experiments modeled after data collected by Johnson, (2005).
+We use the data from this seminal study on careless responding in our empirical application in a later section, where the data are described in more detail.
+Data Generation
+The survey instrument used by Johnson, (2005) is a 300-item representation of the NEO-PI-R (Goldberg,, 1999) measurement of the Big Five personality traits (openness, conscientiousness, extraversion, agreeableness, and neuroticism; Goldberg,, 1992) by means of five-point Likert scales.
+Each of the five traits comprises six subdomains (facets) so that a total of $5\times 6=30$ facets are measured.
+We base our simulation design on the between-item correlation structure and per-item response distributions in the data collected by Johnson, (2005), calculated after removing any presumably careless respondents detected in our empirical application.
+Figure 5 visualizes the resulting correlation matrix and marginal response distributions.
+Of the 300 items, 152 are positively scored while 148 are negatively scored.
+In the correlation matrix, one can identify the five traits and, within each block of within-trait correlations, also the six corresponding facets.
+Using these correlations and marginal response distributions, we generate rating-scale responses of $n=500$ participants to the $p=300$ items,44endnote: 4We use package simstudy (Goldfeld & Wujciak-Jens,, 2020) to generate the data, for which the target values of correlations only hold approximately on the population level.
+with the items in the same randomized order as in Johnson, (2005).
+Figure 5: Correlation matrix and marginal response distributions used in simulations.
+(a) Correlation matrix of the $p=300$ items. A more saturated blue (red) in a cell corresponds to a stronger positive (negative) correlation between the corresponding items.
+(b) Marginal response distributions of the $p=300$ items ($x$-axis), each of which has five response options ($y$-axis).
+A darker shade of gray indicates a higher probability of the corresponding response option.
+Note.
+The correlation matrix and the marginal response distributions are calculated from the data of Johnson, (2005) after removing presumably careless respondents.
+Of the $p=300$ items, 152 are positively scored and 148 are negatively scored.
+The items in both plots are presented in an order such that items measuring the same facet are adjacent.
+Subsequently, a fixed percentage $\gamma\in\{0,0.02,0.04,0.1,0.2,0.3,0.4,0.5,0.6,0.8,1\}$ of respondents is selected to be partially careless.
+That is, $\gamma$ denotes the prevalence of careless responding.
+For each of the selected partially careless respondents, we sample a carelessness onset item from which onward the remaining responses are careless.
+We consider three regimes for the carelessness onset: (i) the baseline setting where the onset item is randomly sampled from $\{30,31,\dots,269,270\}$ so that carelessness starts between 10% and 90% of all items, (ii) early onset randomly sampled from items $\{30,31,\dots,149,150\}$ (between 10% and 50% of all items), and (iii) late onset randomly sampled from items $\{150,151,\dots,269,270\}$ (between 50% and 90% of all items).
+[…]
+We repeat the above described data generating process 100 times.
+We compare three variations of CODERS based on (i) both reconstruction errors (RE) and the LSP sequence, (ii) RE only, and (iii) the LSP sequence only.
+For the autoencoder, we set the number of nodes in the bottleneck layer equal to the number of measured constructs (i.e., 30 facets).
+In all three variations of CODERS, we look for changepoints with three different significance levels $\alpha$, namely 0.1%, 0.5%, and 1% (from most to least conservative).
+Performance Evaluation
+To quantify how accurately partially careless respondents are detected, we compute the false positive rate (FPR) (proportion of true attentive respondents incorrectly flagged as careless) and the false negative rate (FNR) (proportion of true careless respondents incorrectly not flagged as careless).
+For CODERS, we consider a respondent to be flagged as careless if a changepoint is detected.
+In addition, to quantify how accurately the onset of careless responding is detected, we compute the mean absolute error (MAE) (difference between the true onset item and the detected onset item in absolute value, averaged over the set of true careless respondents for whom a changepoint is detected).
+We report averages of these performance measures across the 100 simulation runs.
+For all three measures, a lower value (close to 0) indicates better performance.
+Results
+We focus on discussing the results for CODERS using both measurements of carelessness (RE and LSP), except when we explicitly state otherwise.
+Varying the prevalence of careless responding, Figure 6 visualizes the FPR of CODERS in fully attentive respondents for the different carelessness onset regimes (in separate columns).
+Overall, CODERS rarely flags changepoints in fully attentive respondents.
+Even with our most liberal choice of significance level $\alpha$ (1%), overrejection increases only slowly to about 3% for 60% prevalence, before it reaches between 3.5% and 5.7% at 80% prevalence (depending on the onset regime).
+Other than for extremely high prevalence levels, the FPR stays reasonably close to the respective nominal level, indicating that CODERS controls Type I errors fairly well.
+Figure 6: False positive rate of CODERS for attentive respondents.
+Note. The $x$-axis displays various prevalence levels of partially careless respondents. Results for different onset regimes are shown in separate columns, averaged across 100 repetitions.
+For partially careless respondents, Figure 7 (top row) contains a similar plot of the FNR.
+Most importantly, the FNR of CODERS with both dimensions remains remarkably constant throughout prevalence levels, at about 18–26% in the baseline onset regime depending on the significance level.
+While the FNR increases to about 36% with the most conservative significance level (0.1%) in the early onset regime, it decreases to about 8% with the most liberal significance level (1%) in the late onset regime.
+Overall, the relatively low FNR implies that CODERS successfully identifies a large majority of the true careless respondents.
+Combined with its low FPR, we find that CODERS discriminates well between attentive and partially careless respondents.
+In addition, Figure 7 (bottom row) displays the MAE of the estimated onset in partially careless respondents.
+The prevalence level and onset regime have little effect on the accuracy of the estimates, with an MAE hovering around 2 to 3 depending on the significance level.
+That is, on average the estimated onset is only two to three items away from the true onset, indicating that CODERS achieves excellent accuracy in estimating carelessness onset.
+Figure 7: Performance measures of CODERS for partially careless respondents.
+Note.
+The $x$-axis displays various prevalence levels.
+The top row shows results for the false negative rate, while the bottom row visualizes the accuracy of onset estimation via the mean absolute error.
+Results for different onset regimes are shown in separate columns, averaged across 100 repetitions.
+On the other hand, Figure 7 further reveals that the FNR of CODERS using only RE or LSP is considerably higher in all settings.
+Furthermore, the MAE of the estimated onset is much higher when using only RE.
+It is important to note that such a drop in performance is expected when using only one measurement of carelessness due to the various types of careless responding being simulated.
+RE is designed to capture inconsistent carelessness and may therefore miss invariable carelessness, whereas it is the other way around for LSP.
+To investigate this further, Figures 8 and 9 show the FNR and the MAE of onset estimation averaged over subgroups of participants corresponding to each type of careless responding (in separate rows).
+For the two inconsistent types of carelessness (top two rows), using both dimensions yields overall good results similar to to those using only RE, whereas performance is clearly worse when using only LSP.
+For the two invariable types of carelessness (bottom two rows), the results when using both dimensions are near-perfect (with FNR and MAE close to 0) and indistinguishable from using only LSP, while performance deteriorates when using only RE.
+Hence, these findings highlight the benefit of combining measurements of different dimensions of carelessness in CODERS.
+Figure 8: False negative rate of CODERS for different types of partially careless respondents.
+Note. The $x$-axis displays various prevalence levels. Results for different types of careless responding (rows) in different onset regimes (columns) are averaged across 100 repetitions.
+Figure 9: Mean absolute error of carelessness onset estimation with CODERS for different types of partially careless respondents.
+Note. The $x$-axis displays various prevalence levels. Results for different types of careless responding (rows) in different onset regimes (columns) are averaged across 100 repetitions.
+We now zoom in on the results of CODERS for the two types of inconsistent carelessness in Figures 8 and 9, as there is notable heterogeneity.
+Specifically, CODERS performs clearly better for extreme responding than for random responding, with the FNR and the MAE of onset estimation in general being lower.
+Considering that the items exhibit a variety of different probability distributions for the attentive responses, with a low probability mass in at least one of the extreme answer categories (see Figure 5), distinguishing random responding is simply a much more challenging problem than distinguishing extreme responding.
+For random responding (see the first row in Figure 8), CODERS only achieves a decent FNR in the late onset regime with significance level 1% or 0.5% (about 32% and 40%, respectively).
+In the baseline onset regime, its FNR is higher than 50%.
+In the early onset regime, even our most liberal significance level of 1% only yields an FNR of about 75%.
+Nevertheless, this finding should not be mistaken as an intrinsic inability of CODERS to detect random responding.
+As the results greatly improve the later the onset of carelessness, it seems that the autoencoder needs enough attentive responses to learn the challenging task of distinguishing random responses.
+In other words, this is an issue of statistical power, which explains why it can be mitigated to some extent by choosing a less conservative significance level.
+For additional investigation, we apply traditional methods for identifying careless respondents also used by Johnson, (2005).
+They do not improve upon CODERS (see Appendix C.1), therefore further highlighting the difficulty of detecting random respondents in this simulation design.
+Coming back to CODERS, it shows similar but far less pronounced behavior for extreme responding (see the second row in Figure 8).
+The FNR remains decent in the early onset regime, while it is good in the baseline onset regime and excellent (close to 0) in the late onset regime.
+To summarize our findings, CODERS performs well for extreme responding and near-perfectly for the two types of invariable responding.
+On the other hand, it struggles with random responding in this simulation design, where decent power is only achieved in case of late carelessness onset and with a not-too-conservative significance level.
+However, results for traditional methods (see Appendix C.1) suggest that distinguishing random responding from attentive responding may be an intrinsically difficult problem for the chosen correlation structure and marginal distributions.
+Overall, we find that (i) CODERS reliably discriminates between attentive and partially careless respondents, as indicated by near-zero false positive rates and low overall false negative rates, (ii) CODERS accurately estimates the location of carelessness onset, and (iii) the performance of CODERS is stable across a wide range of prevalence levels, demonstrating robustness.
+Additional Simulations
+We conducted additional simulation experiments to further investigate the robustness of our findings and to obtain a better understanding of possible limitations of CODERS.
+While details and results can be found in Appendix C, we provide a brief outline of each experiment and the most relevant findings below.
+First, we vary the number of items to investigate the performance of CODERS in shorter surveys.
+Recall that our simulation design is based on measurements of the Big Five personality traits, with each trait being measured by 60 items (for a total of 300 items).
+We randomly select traits in a stepwise manner and drop all items of the selected traits, resulting in surveys of length $p\in\{240,180,120,60\}$.
+The overall FNR in partially careless respondents increases only slowly as survey length decreases.
+For $p=60$, only very few random respondents are detected, but the FNR remains decent for extreme responding and near-perfect for the invariable types of carelessness.
+Furthermore, the accuracy of the estimated onset of carelessness stays excellent throughout the investigated survey lengths.
+In summary, other than for random respondents, overall power remains adequate when using a not-too-conservative significance level (e.g., 0.5% or 1%).
+[…]
+Third, we consider an additional type of careless responding to investigate the performance of CODERS when the lines between attentive and careless responding are blurry.
+This type of carelessness, to which we refer as middling, is characterized by randomly choosing from the three middle answer categories (out of five).
+Given the correlation structure and marginal distributions of attentive responding (see Figure 5), there is plenty of natural variation among the central response categories, thus making middling extremely hard to distinguish.
+Indeed, although CODERS maintains a low FPR in fully attentive respondents, it detects only very few middling respondents (as do the traditional methods).
+This further highlights that CODERS requires (inconsistent) partially careless responding to be sufficiently distinct from attentive responding to successfully distinguish between them.
+Fourth, we emulate a situation of temporary carelessness with two changepoints: the first constitutes a change from attentive to careless responding, and the second embodies a change back to attentive responding.
+We find that CODERS fails to detect such temporarily careless respondents, but the same applies to the traditional methods (except for the longstring index, which successfully captures temporary straightliners).
+This highlights that while CODERS is effective in detecting a single changepoint, it is simply not designed for detecting multiple changepoints.
+Finally, having so far considered simulations based on personality assessments, we include a simulation inspired by a study from organizational research.
+This simulated survey is shorter, with $p=198$ items measuring a variety of organizationally relevant constructs, and we apply a different strategy for randomization better suited in organizational contexts.
+Furthermore, there exist complex relationships between (some of) the measured constructs.
+The FPR of CODERS is somewhat more elevated for high prevalence levels.
+Interestingly, the FNR is slightly lower for random responding but slightly higher for extreme responding, resulting in a comparable FNR overall.
+All in all, the results for CODERS remain relatively similar to those in the simulations based on personality assessments, indicating that its performance may not depend much on the institutional setting of the survey.
 Empirical Application
-Ideally, new methodology is empirically validated by conducting an experiment that emulates the phenomenon of interest in lab conditions.
-For instance, Schroeders et al., (2022) conduct an experiment in which they instruct one part of the participants to respond truthfully and another part to respond carelessly in order to investigate whether various screeners accurately identify the subset of careless respondents.
-For any of the compared methods, at most 19% of the respondents flagged as careless were in fact instructed to be careless.
-The authors therefore express doubts whether the study participants complied with the experimental instructions.
-In addition to the risk of non-compliance, Ulitzsch et al., 2022a ; Ulitzsch et al., 2022b stress that another major limitation of this experimental approach is the doubtful assumption that participants instructed to respond carelessly behave in a comparable manner to careless respondents outside of such lab experiments.
-Indeed, it seems plausible that careless responding stems from an intrinsic state-of-mind rather than being a conscious action that can be instructed (cf.
-Alfons & Welz,, 2024, for additional discussion).
-Hence, rather than conducting such an experiment, we take data from a seminal study on careless responding and further examine the responses for partial carelessness.
+After studying the performance of CODERS in simulations inspired by the data from Johnson, (2005), we now re-examine the responses from this seminal study for partial carelessness.
 Data
-We apply CODERS to an empirical data set collected by Johnson, (2005).
-The data concern the five factor model of personality (Goldberg,, 1992), which assumes that variation in a measurement of personality can be explained by the so-called Big Five personality traits (factors): openness, conscientiousness, extraversion, agreeableness, and neuroticism.
+The data collected by Johnson, (2005) concern the five factor model of personality (Goldberg,, 1992), which assumes that variation in a measurement of personality can be explained by the so-called Big Five personality traits (factors): openness, conscientiousness, extraversion, agreeableness, and neuroticism.
 Each of the Big Five factors can be further decomposed into six subdomains (facets): for instance, depression and anxiety are facets of the neuroticism trait.
-Consquently, common Big Five instruments, such as the 240-item NEO-PI-R (Costa & McCrae,, 1992), typically measure a total of $5\times 6=30$ facets.
+Consequently, common Big Five instruments, such as the 240-item NEO-PI-R (Costa & McCrae,, 1992), typically measure a total of $5\times 6=30$ facets.
 The data set of Johnson, (2005) consists of internet-collected responses to a 300-item representation of the NEO-PI-R inventory (Goldberg,, 1999), in which each of the 30 facets is measured by 10 items on a five-point Likert scale.
@@
 For instance, Bowling et al., (2021) estimate that if one wishes that at least 90% of participants do not respond carelessly to more than 20% of all items, one should not include more than 176 items in an online questionnaire.
-For the study of Johnson, (2005), it is important to keep in mind that the author applied two methods for detecting inconsistent careless respondents to the data from which straightliners are already removed.
-The two methods are designed to detect participants who respond carelessly to all or at least the vast majority of survey items.
-It is unclear how well they perform in identifying partially careless respondents, who may respond carelessly to only a relatively small number of items.
-In addition, Johnson, (2005) did not apply any method designed to identify invariable careless responding other than straightlining.
-By applying CODERS, which is designed to identify partially careless responding for various types of carelessness, new insights into the incidence and nature of (partially) careless responding may emerge.
+By applying CODERS—which is specifically designed to identify partially careless responding through a change in response behavior—in addition to the traditional screeners applied by Johnson, (2005), new and complimentary insights into the incidence and nature of (partially) careless responding may emerge.
 Results
 […]
-Comparison with Benchmark Methods
+General Overview
 For obtaining the autoencoder reconstruction errors in CODERS, we set the number of nodes in the bottleneck layer to 30, i.e., equal the number of facets in the Big Five model.
 Table 1 summarizes the results of CODERS.
-Depending on the significance level $\alpha$, CODERS flags between 164 (0.7%) participants for $\alpha=0.001$ and 850 (3.8%) participants for $\alpha=0.01$ as partially careless.
-On average, the location of the carelessness onset item for flagged participants is estimated around the 150th item, meaning that carelessness onsets on average after half of the 300 total items.
-These findings are broadly in line with the results of Bowling et al., (2021, Figure 1), whose model suggests a low probability that a representative participant starts responding carelessly at the 150th item.
-The earliest onset occurs at item 12, whereas the latest onset is at item 294.
-A more detailed visual overview of the detected onset items for the flagged participants can be found in Figure C.1 in Appendix C.2.
-For comparison, we consider personal reliability and psychometric antonym scores.
-Along the given responses, Johnson, (2005) provides per-individual estimates for either score, which we use in our analysis.
-Importantly, Johnson, reverses psychometric antonym scores so that low scores are associated with carelessness in both methods.
-Following Johnson,, we flag respondents as careless whose personal reliability and psychometric antonym scores are lower than $0.3$ and $-0.03$, respectively.
+Depending on the significance level $\alpha$, CODERS flags between 181 (0.8%) participants for $\alpha=0.001$ and 856 (3.8%) participants for $\alpha=0.01$ as partially careless.
+On average, the location of the carelessness onset item for flagged participants is estimated around the 145th item, meaning that carelessness onsets on average after nearly half of the 300 items.
+These findings are broadly in line with the results of Bowling et al., (2021, Figure 1), whose model suggests a low probability that a representative participant starts responding carelessly at the 145th item.
+The earliest onset occurs at item 15, whereas the latest onset is at item 294.
+Figure 10 provides a more detailed visual overview of the detected onset items for the flagged participants at different significance levels.
+Table 1: Results of CODERS for the data of Johnson, (2005).
+Detected onset
+Significance level Number of flagged respondents Minimum Average Maximum
+$\alpha=0.001$ 181 (0.8%) 19 143.7 292
+$\alpha=0.005$ 571 (2.5%) 19 145.8 294
+$\alpha=0.01\phantom{0}$ 856 (3.8%) 15 147.9 294
+[…]
+Figure 10: Estimated onset of careless responding for participants flagged by CODERS in the data of Johnson, (2005).
+Note.
+Results for different significance levels $\alpha\in\{0.01,0.05,0.1\}$ are presented in separate panels.
+The $x$-axis shows the indices of the participants flagged as partially carelessness by CODERS.
+Here, the participants are sorted in ascending order of the estimated carelessness onset item, which is in turn depicted on the $y$-axis.
+The dashed horizontal line represents the average location of carelessness onset across the flagged participants.
+We now turn to analyzing the agreement between CODERS and the two traditional methods applied by Johnson, (2005), personal reliability and psychometric antonym.
+Following Johnson,, we reverse psychometric antonym scores so that low scores are associated with carelessness in both methods, and we flag respondents as careless whose personal reliability and psychometric antonym scores are lower than $0.3$ and $-0.03$, respectively.
 The results are shown in Table 2.
-Personal reliability flags 40 participants (0.2%) as careless, while psychometric antonym flags 186 respondents (0.8%).
+Personal reliability flags 140 participants (0.6%) as careless, while psychometric antonym flags 284 respondents (1.3%).66endnote: 6We use the R package careless (Yentes & Wilhelm,, 2023) to compute personal reliability and psychometric antonym scores.
+The use of different software implementations (with potentially different default values) may explain differences with the numbers reported by Johnson, (2005).
 CODERS, personal reliability, and psychometric antonyms generally show little agreement regarding the detected careless respondents.
-Personal reliability and psychometric antonym scores correlate by a moderate value of 0.488, yielding only 13 participant that are flagged by both methods.
-Depending on the significance level $\alpha$, CODERS and personal reliability have between 1 and 4 flagged participants in common, while the number of jointly flagged respondents for CODERS and psychometric antonym ranges between 5 and 10.
+Personal reliability and psychometric antonym scores correlate by a low value of 0.314, yielding only 11 participants that are flagged by both methods.
+Depending on the significance level $\alpha$, CODERS and personal reliability have between 2 and 10 flagged participants in common, while the number of jointly flagged respondents for CODERS and psychometric antonym ranges between 9 and 24.
 All three methods share a single flagged participant.
-Table 1: Comparison of CODERS with benchmark methods for the data of Johnson, (2005).
+Table 2: Complementary findings of CODERS and traditional methods for the data of Johnson, (2005).
 Also flagged by CODERS
 Method Number of flagged respondents $\alpha=0.001$ $\alpha=0.005$ $\alpha=0.01$
-Reliability 40 (0.2%) 1 3 4
-Antonym 186 (0.8%) 5 7 10
-Both 13 1 1 1
+Reliability 140 (0.6%) 2 18 10
+Antonym 284 (1.3%) 9 18 24
+Both 11 1 11 21
 […]
-It is not surprising that there is little to no consensus of CODERS with the two benchmark methods, as the former is designed for a different task: finding respondents that have been careless in parts of the questionnaire (including identifying the onset item) rather than throughout the entire questionnaire.
-This disagreement indicates that partial carelessness may be present in the data, which the benchmark methods may be unable to identify.
-We therefore take a closer look at a representative selection of respondents and analyze their per-item reconstruction errors and LSP sequences, which are shown in Figure 3.
-Table 2 contains additional summary statistics for these measurements before and after the estimated careless onset item (if such an item is found by CODERS).
-Two of the selected participants are flagged as careless by CODERS but not by either of the two benchmark methods, one participant is flagged by both CODERS and psychometric antonym, and one participant is flagged by both benchmark methods but not CODERS.
-Figure 3: Scores and careless onset item detected by CODERS in selected participants from the data of Johnson, (2005).
-(a) Flagged by CODERS only, presumably inconsistent. Onset is estimated at item 148.
-(b) Flagged by CODERS only, presumably invariable. Onset is estimated at item 58.
-(c) Flagged by antonym and CODERS, presumably inconsistent. CODERS estimates onset at item 119.
-(d) Flagged by antonym and personal reliability, but not CODERS (i.e., there is no estimated onset item).
+It is not surprising that there is little consensus of CODERS with the two traditional methods, as they are designed for different tasks.
+The traditional methods aim at identifying which respondents are careless, whereas CODERS aims to detect the onset of careless responding in partially careless respondents.
+Hence, this disagreement indicates that (i) there may be respondents who respond carelessly throughout the survey, which CODERS cannot identify due to the absence of a changepoint (but which are found by the traditional methods), and (ii) partial carelessness may be present in the data, for which traditional methods may have low detection power (but which are found by CODERS).
+Moreover, CODERS uncovers novel insights on careless responding behavior, which we highlight with a selection of respondents below.
+Further Insights on Individual Cases
+To gain insights into the behavior of selected respondents, we analyze their per-item reconstruction errors and LSP sequences, which are shown in Figure 11.
+Table 3 contains additional summary statistics for these measurements before and after the estimated changepoint item (if such an item is found by CODERS).
+While these examples are naturally selective, similar plots for all participants flagged by CODERS or the two traditional methods can be found in our online repository at https://github.com/mwelz/carelessonset_plots.
+Figure 11: Scores and changepoint item detected by CODERS in selected participants from the data of Johnson, (2005).
+(a) Presumably inconsistent. Onset estimated by CODERS at item 148.
+(b) Presumably inconsistent. Onset estimated by CODERS at item 246.
+(c) Presumably inconsistent. Changepoint detected by CODERS at item 63.
+(d) Presumably inconsistent first, then invariable. Changepoint detected by CODERS at item 58.
+(e) Presumably invariable. Changepoint detected by CODERS at item 142.
+(f) Presumably inconsistent throughout the survey (no changepoint item detected by CODERS).
 Note.
 In each subplot, autoencoder reconstruction errors (RE) and LongStringPattern (LSP) sequences are shown in separate panels.
-The solid vertical lines indicate the identified careless onset items.
+The solid vertical lines indicate the identified changepoint items.
 Summary statistics for the corresponding participants are provided in Table 3.
-Table 2: Average scores before and after the careless onset item detected by CODERS in selected participants from the data of Johnson, (2005).
+Table 3: Average scores before and after the changepoint item detected by CODERS in selected participants from the data of Johnson, (2005).
 Before After
 RE 0.019 (0.035) 0.069 (0.082)
-LSP 4.197 (1.396) 3.497 (1.433)
-(a) Flagged by CODERS only, presumably inconsistent. Onset is estimated at item 148.
+LSP 4.122 (1.428) 3.373 (1.464)
+(a) Presumably inconsistent. Onset estimated by CODERS at item 148.
+Before After
+RE 0.017 (0.026) 0.070 (0.096)
+LSP 2.886 (1.189) 3.236 (1.815)
+(b) Presumably inconsistent. Onset estimated by CODERS at item 246.
+Before After
+RE 0.095 (0.119) 0.024 (0.036)
+LSP 2.565 (0.917) 2.668 (0.925)
+(c) Presumably inconsistent. Changepoint detected by CODERS at item 63.
 Before After
 RE 0.078 (0.103) 0.025 (0.035)
-LSP 2.421 (0.706) 19.770 (10.941)
-(b) Flagged by CODERS only, presumably invariable. Onset is estimated at item 58.
+LSP 2.158 (0.621) 19.708 (11.027)
+(d) Presumably inconsistent first, then invariable. Changepoint detected by CODERS at item 58.
 Before After
-RE 0.019 (0.022) 0.039 (0.059)
-LSP 3.619 (1.371) 2.687 (0.972)
-(c) Flagged by antonym and CODERS, presumably inconsistent. CODERS estimates onset at item 119.
+RE 0.018 (0.029) 0.024 (0.046)
+LSP 146.000 (0.000) 23.597 (25.878)
+(e) Presumably invariable. Changepoint detected by CODERS at item 142.
 Overall
-RE 0.064 (0.066)
-LSP 2.660 (0.753)
-(d) Flagged by antonym and personal reliability, but not CODERS (i.e., there is no estimated onset item).
+RE 0.079 (0.105)
+LSP 2.813 (1.115)
+(f) Presumably inconsistent throughout the survey (no changepoint item detected by CODERS).
 Note.
 Standard deviations are given in parentheses.
-The subtables correspond to the four selected participants from Figure 3.
+The subtables correspond to the selected participants from Figure 11.
 Positioning of subtables is consistent with positioning of the subfigures in Figure 11, e.g., Table 3(a) refers to the same participant as Figure 11(a).
-First, we focus on Figure 3(a) and Table 2(a).
-CODERS identifies carelessness onset at item 148, but neither benchmark method flags this respondent as careless.
+First, we focus on the two participants in Figures 11(a)11(b) and Tables 3(a)3(b).
+CODERS identifies carelessness onset at item 148 and item 246, respectively.
 We observe that the average reconstruction error, as well as the variability in the reconstruction errors, is much larger before the estimated onset item than afterward.
-Since the pre-onset responses could be reconstructed well (except for a brief spike around item 110) while post-onset responses could not, there is evidence that the participant started the questionnaire attentively, but succumbed to survey fatigue about halfway through and subsequently engaged in inconsistent careless respondening.
-The second example from Figure 3(b) and Table 2(b) shows another participant that is flagged by CODERS but no benchmark method.
-The estimated onset of careless responding is at item 58.
-While the LSP sequence clearly changes around this item, even more excessive values are found for later items, with a maximum LSP value of 31.
+Since the pre-onset responses could (generally) be reconstructed well while post-onset responses could not, there is evidence that these participants succumbed to survey fatigue and eventually engaged in inconsistent careless responding.
+These two cases are representative for the majority of respondents flagged by CODERS, likely starting attentively but starting to respond carelessly at some point.
+In addition, more respondents flagged by CODERS seem to exhibit inconsistency than invariability, as Johnson, already removed straightliners from the publicly available data set.
+However, CODERS also identifies participants with remarkably different behavior.
+The participant in Figure 11(c) and Table 3(c) shows similar behavior as the first two examples, albeit in reverse.
+CODERS identifies a changepoint at item 63, with reconstruction errors being large in terms of magnitude and variability before the changepoint while being small afterward.
+We conclude that this participant may have started the survey as an inconsistent careless respondent, but then shifted toward attentive responding.
+For the participant in Figure 11(d) and Table 3(d), CODERS identifies a changepoint at item 58.
+Both the reconstruction errors and the LSP sequence clearly change around this item, with even more excessive LSP values occurring in later items (maximum value of 31).
 In fact, only four out of the $n=20,994$ participants in the sample yield more extreme LSP values.
-Hence, there is strong evidence that this participant responded with (increasingly) lengthy sequences of fixed patterns as the questionnaire progressed.
-That is, they likely started engaging in invariable careless responding.
-The third example, shown in Figure 3(c) and Table 2(c), is similar to the first example in the sense that it suggests an initially attentive respondent starting to give inconsistent careless responses from some point onward in the questionnaire, as indicated by a higher average and a larger variability in the reconstruction errors.
-Here, the careless onset is estimated at item 119.
-The crucial difference with the first example is that this participant was not only flagged as careless by CODERS, but also by psychometric antonym.
-A plausible explanation for this difference may be found in the earlier careless onset.
-If a participant responds carelessly to a large enough majority of the items, a method designed to detect carelessness throughout the questionnaire (such as the benchmark methods) has a better chance at flagging that participant as careless.
+Furthermore, the reconstruction errors before the changepoint are of comparable magnitude and variability to those of the previous examples of inconsistent responding.
+Hence, there is strong evidence that this participant initially responded inconsistently, but switched to invariable responding characterized by (increasingly) lengthy sequences of fixed patterns as the questionnaire progressed.
+Next, we discuss the example in Figure 11(e) and Table 3(e), for whom CODERS detects a changepoint at item 142.
+It turns out that this respondent repeated the pattern 5-5-3-3-3 in the first half of the survey.
+Even after the changepoint, the LSP values are among the highest in the sample, with the participant switching between the similar patterns 1-1-3-3-3, 5-1-3-3-3, and 1-5-3-3-3 for the remainder of the questionnaire.
+Although there is a clear change in behavior, as found by CODERS, this participant seems to respond invariably throughout the survey.
 […]
-While these four examples are naturally selective, similar plots as in Figure 3 for all participants flagged by CODERS or the two benchmark methods can be found in our online repository at https://github.com/mwelz/carelessonset_plots.
-In general, two key insights can be derived from the above analysis.
-First, we find empirical evidence that CODERS can indeed detect various types of partial careless responding, specifically inconsistent and invariable carelessness.
-Common screeners for careless responding throughout the questionnaire, such as personal reliability or psychometric antonym, are typically only effective for one type of careless responding.
-Second, CODERS is not a replacement for existing screeners, but rather a complimentary tool that provides new insights if participants display careless responding behavior only in part of the questionnaire.
-Comparison with Johnson, (2005)
-We now compare our findings with those reported by Johnson, (2005) regarding the incidence of careless responding. This comparison is summarized in Table 3.
-Table 3: Incidence of careless responding in the data of Johnson, (2005).
-Johnson, (2005) Our findings
-Type of careless responding Method Number Method Number
-Straightliners Longstring + screeplot 808 (3.6%)
-Longstring + cutoff by Costa & McCrae, (2008) 1455 (6.5%)
-Inconsistent Reliability 41 (0.2%) Reliability 40 (0.2%)
-Antonym 185 (0.8%) Antonym 186 (0.8%)
-Either 213 (1.0%) Either 213 (1.0%)
-Partial CODERS ($\alpha=0.001$) 158 (0.7%)
-CODERS ($\alpha=0.005$) 580 (2.6%)
-CODERS ($\alpha=0.010$) 836 (3.7%)
-[…]
-Johnson, explores two different cutoff values for the longstring index to identify straightliners: one obtained via a screeplot and another suggested by Costa & McCrae, (2008),66endnote: 6Johnson, (2005) refers to a pre-publication version of Costa & McCrae, (2008).
-with the latter yielding almost twice as many flagged respondents (3.6% vs.
-6.5%).
-As the data set provided by the author already excludes straightliners (among others), we could not attempt to replicate these findings.
-Regarding inconsistent careless responding, we find minor differences in the number of flagged respondents compared to Johnson,: 40 vs.
-41 for personal reliability, and 186 vs.
-185 for psychometric antonym.
-These minor differences are potentially due to rounding: the psychometric antonym and personal reliability scores provided by Johnson, are rounded by two decimal places, whereas the scores used to compute the reported results in Johnson, (2005) were likely unrounded.
-As in Johnson, (2005), we find that personal reliability and psychometric antonym agree on 13 of the flagged participants.
-Following strategies that “[eliminate] more non-attentive responders, albeit at the cost of more false positives”, Johnson, (2005, p.
-116) applies the cutoff value of Costa & McCrae, (2008) for identifying straightliners, and considers respondents that are found by either personal reliability or psychometric antonym as inconsistent careless respondents.
-This results in a total of 1668 detected careless respondents (7.4%).
-Employing a similar strategy regarding CODERS, one may choose a significance level of 1%.
-Adding the corresponding 836 participants yields a total of 2504 flagged careless respondents (11.2%).
-Hence, we find evidence that the incidence of careless responding might be substantially underreported in Johnson, (2005) due to hitherto undetected partially careless respondents.
-Moreover, an incidence of 11.2% in this data set would be higher than the 5–10% that the literature deems problematic for further analysis (e.g., Arias et al.,, 2020; Credé,, 2010).
-Simulation Experiments
-After demonstrating how CODERS yields new insights in data from a seminal study, we further study the behavior of CODERS via simulation experiments inspired by existing survey measurements.
-Data Generation
-We generate rating-scale responses of $n=500$ respondents to $p=300$ items.
-Each item has five Likert-type answer categories (anchored by 1 = “strongly disagree” and 5 = “strongly agree”).
-The simulated questionnaire measures 30 constructs, each of which is measured by a scale of 10 items.
-We simulate reliable scales by imposing that items from the same scale have a correlation of 0.7, which results in a Cronbach’s $\alpha$ value of 0.959.77endnote: 7We use package simstudy (Goldfeld & Wujciak-Jens,, 2020) to generate the data, for which the target values of correlations only hold approximately on the population level.
-Hence, the same applies to Cronbach’s $\alpha$ values.
-Items from different scales are uncorrelated.
-All 10 items within a given scale follow the same response probability distribution.
-We thereby consider the two types of distributions defined in Table 4: centered about the midpoint and skewed towards agreeing.
-Each of the two distributions is used in 15 scales, yielding the total of $2\times 15\times 10=300$ items.
-We reverse five items per scale (i.e., half the items) and randomize the order of the items.
-Note that distributions skewed towards disagreement are implicitly covered by this design because of the item reversal.
-Table 4: Response probability distributions of items within different scales.
-Type of distribution $P[X=1]$ $P[X=2]$ $P[X=3]$ $P[X=4]$ $P[X=5]$
-Centered 0.15 0.20 0.30 0.20 0.15
-Skewed 0.10 0.15 0.20 0.25 0.30
-Note. The random variable $X$ denotes the response to an item with five Likert-type answer categories.
-Subsequently, we select a fixed percentage $\gamma\in\{0,0.02,0.04,0.1,0.2,\dots,0.6,0.8,1\}$ of respondents to be partially careless.
-That is, $\gamma$ denotes the prevalence of careless responding.
-For each of the selected partially careless respondents, we sample a carelessness onset item from which onward the remaining responses are careless.
-We consider two settings for the carelessness onset: (i) early onset, which we randomly sample from the item set $\{30,31,\dots,269,270\}$ so that careless responding starts between 10% and 90% of all items, and (ii) later onset, which we randomly sample from $\{150,151,\dots,269,270\}$, i.e., between half and 90% of all items.
-[…]
-We repeat the above described data generating process 100 times.
-Importantly, we keep the location of carelessness onset item of each participant constant across the 100 repetitions, which allows for neat visualization of the results.
-We compare three variations of CODERS based on (i) both reconstruction errors and the LSP sequence, (ii) reconstruction errors only, and (iii) the LSP sequence only.
-For the autoencoder, we set the number of nodes in the bottleneck layer equal to the number of measured constructs (i.e., 30).
-Performance Evaluation
-To quantify how accurately the onset of careless responding is detected, we compute the adjusted Rand index (ARI) (Hubert & Arabie,, 1985, see Appendix D.1 for details).
-The ARI is best known as a measure of clustering performance: a value of 0 is expected for a random clustering, while a perfectly recovered clustering results in a value of 1.
-In the careless onset detection problem for a given respondent, we can view the items before the careless onset as one cluster and the items from the onset onwards as another.
-Specifically, the closer the ARI is to its maximum value 1, the better CODERS performs at accurately locating the true carelessness onset item.
-On the other hand, an ARI value close to 0 indicates that the estimated onset is far from the true onset.
-If a respondent is attentive throughout the questionnaire but CODERS incorrectly identifies a changepoint, the ARI takes a value of 0.
-The same holds if CODERS fails to flag a changepoint for a partially careless respondent.
-We report the average ARI across respondents.
-Note that averaging the ARI over attentive respondents yields the proportion of attentive respondents who are correctly identified as attentive.
-Conversely, subtracting this number from 1 gives the proportion of attentive respondents that are incorrectly flagged as partially careless.
-Results
-We start presenting the results of CODERS for the three dimensions by means of a visual inspection.
-For a carelessness prevalence level of $\gamma=0.6$ and the early onset setting, Figure 4 displays a heatmap holding the indices of the participants in its rows and the item indices in its columns.
-For participant $i\in\{1,\dots,n\}$ and item $j\in\{1,\dots,p\}$, the color scale of the cell at position $(i,j)$ measures the frequency across the 100 repetitions with which the response of participant $i$ to item $j$ has been flagged as changepoint: a more saturated red indicates that the corresponding item was flagged more frequently as a changepoint.
-The black rectangles visualize the true location of the careless onset items (recall that those are fixed across the repetitions).
-The $y$-axis of the plot is rearranged so that the participant indices are ordered according to their respective carelessness onset.
-Good detection performance is characterized by flagging cells that are close to the corresponding onset items.
-Indeed, we can see that there is a saturated red band that closely follows the black rectangles.
-Note that the top 40% of the rows in Figure 4 very rarely see flagged changepoints.
-Those are the respondents who never become careless.
-Based on this visual examination, CODERS seems to not only accurately estimate the careless onset, but also discriminate well between partially careless respondents and fully attentive respondents.
-Results are similar for other combinations of carelessness prevalence and onset, hence we omit the corresponding heatmaps.
-Figure 4: Heatmap of carelessness onset detection performance of CODERS.
-Note.
-Of the $n=500$ respondents, 60% are partially careless according to the early onset setting.
-The significance level of CODERS is set to 0.1%.
-A more saturated red in a cell corresponds to a higher frequency across the 100 repetitions of the cell being flagged as carelessness onset item ($x$-axis) of a given respondent ($y$-axis).
-The black rectangles denote the true carelessness onset items (which are constant across the repetitions) and the $y$-axis is rearranged according to the location of true carelessness onset.
-An absence of a black rectangle (occurs in 40% of all respondents) indicates that a given respondent never becomes careless.
-Figure 5: ARI of CODERS for various carelessness prevalence levels.
-Note.
-The ARI is averaged over fully attentive respondents and partially careless respondents (columns) for different carelessness onset settings (rows), and the significance level of CODERS is set to 0.1%.
-Varying the prevalence of careless responding, Figure 5 visualizes the ARI averaged over attentive respondents (left column) and partially careless respondents (right column) for the early onset (top row) and later onset (bottom row) settings.
-Each of the three variations of CODERS rarely flag changepoints in fully attentive respondents, except when carelessness prevalence is extremely high.
-While the percentage of false positives is close to the significance level 0.1% until 40% prevalence, overrejection remains moderate at 1% for 60% prevalence, before it rises to about 5% for early onset and 10% for later onset at 80% prevalence.
-For partially careless respondents, the ARI of CODERS with both dimensions remains remarkably constant throughout carelessness prevalence levels at about 0.75 and 0.85 for early and later onset, respectively, mirroring an excellent performance of flagging changepoints in correct locations.
-In contrast, the ARI using only the LSP sequences amounts to about 0.6 throughout prevalence levels in both onset settings, whereas the ARI using only the reconstruction errors drops from about 0.5 for 2% prevalence to about 0.3 for 100% prevalence with early onset, and from about 0.75 to about 0.35 with later onset.
-It is important to note that such a drop in performance is expected when using only one indicator of carelessness due to various types of careless responding being simulated.
-Reconstruction errors are designed to capture inconsistent carelessness and may therefore miss invariant carelessness, while LSP sequences are designed to capture invariant carelessness and may therefore miss inconsistent carelessness.
-Figure 6: ARI of CODERS for various carelessness prevalence levels.
-Note.
-The ARI is averaged over the participants corresponding to each careless response type (columns) for different carelessness onset settings (rows), and the significance level of CODERS is set to 0.1%.
-Note that for straightlining and pattern responding, the results for CODERS using all two dimensions and LongstringPattern only are nearly identical, resulting in overplotted lines.
-To explore this further, Figure 6 displays the ARI averaged over subgroups of participants corresponding to each type of careless responding (separate columns) for early onset (top row) and later onset (bottom row).
-Indeed, reconstruction errors yield much better performance for inconsistent carelessness (random and extreme responding) than for invariable carelessness (straightlining and pattern responding), in particular in the later onset setting.
-Conversely, using LSP sequences results in near-perfect detection of straightlining and pattern responding with an ARI of about 0.99, but performs particularly poorly for random responding with an ARI close to 0.
-This highlights the importance of combining the two dimensions in CODERS for accurately detecting various types of partially careless responding.
-We now zoom in on the results in Figure 6 for CODERS using reconstruction errors only, which vary by type of (inconsistent) carelessness, as well as by carelessness prevalence and onset location.
-First, the performance is clearly better for extreme responding than for random responding, with the ARI being in general higher and remaining more stable across prevalence levels.
-Considering the relatively flat probability distributions of the attentive responses (see Table 4), distinguishing random responses is simply a much more challenging problem.
-Second, the heterogeneity in performance across prevalence levels and onset location is likely due to the autoencoder requiring sufficiently many attentive responses to successfully distinguish between attentive and (inconsistent) careless responding.
-If carelessness tends to onset later in the survey, the autoencoder has more information to learn attentive response patterns, making it easier to distinguish between attentive and careless responses.
-Similar reasoning applies for carelessness prevalence.
-When fewer responses are careless, the autoencoder has access to more attentive responses from which to learn.
-In general, these results suggest that CODERS can accurately identify inconsistent careless responding unless careless responses are excessively present in the data through a high prevalence of careless respondents whose carelessness onset occurs early in the survey.
-However, having an overwhelmingly large number of participants with largely uninformative responses suggests a fundamental issue with data quality, in case of which any subsequent analysis may yield questionable results.
-To summarize, this simulation experiment suggests that CODERS succeeds in combining complimentary indicators for carelessness onset in order to capture different manifestations of careless responding.
-Additional Simulations
-We have conducted a variety of additional simulation experiments, which are described and discussed in Appendix D.2.
-Specifically, we investigate different choices of the significance level in CODERS, varying correlation structures (and therefore measurement reliability), as well as varying sample sizes.
-Overall, the findings presented above are representative for those situations.
-For low to moderate carelessness prevalence (as in the empirical application), choosing a higher significance level (e.g., 1%) leads to an increase in power while retaining size control.
-On the other hand, for high prevalence levels (about 40% or higher), increasing the significance level yields too many false positives.
-As expected, power decreases (increases) slightly with decreasing (increasing) measurement reliability or sample size.
-We conclude from our simulation experiments that CODERS reliably detects carelessness onset while successfully discriminating between partially careless respondents and fully attentive respondents.
-In addition, our simulations highlight the importance of using multiple indicators to capture different manifestations of carelessness (cf.
-Goldammer et al.,, 2020; Huang et al.,, 2012; Meade & Craig,, 2012; Ward & Meade,, 2023).
-Extensions, Limitations and Alternatives
-Below, we address the potential incorporation of response times into CODERS, and we discuss general limitations as well as potential alternatives.
+In summary, we reiterate that the first two examples are representative for the majority of participants flagged by CODERS, in line with the assumption that respondents start attentive and possibly become careless.
+Although other cases indicate that this assumption is too restrictive in practice, CODERS—which processes item-level indicators of inconsistency and invariability for changepoints—successfully uncovered relevant changes in response behavior.
+In light of these findings, we proceed to discuss limitations and potential extensions.
+Limitations and Extensions
+We start by reviewing general limitations of CODERS together with (possible) remedies, before discussing the potential incorporation of response times into CODERS.
+Limitations
+First, some participants may respond carelessly throughout all survey items (cf.
+Figure 11(f) and Table 3(f)).
+Consequently, CODERS is unlikely to flag a changepoint due to an absence of a change in behavior.
+Nevertheless, the literature has established a plethora of methods for detecting careless respondents.
+In this sense, CODERS—which is intended for detecting partial carelessness through a change in behavior—is complementary to those existing methods.
+If one suspects that respondents who have been careless throughout all items are present, we recommend to apply established detection methods following the guidelines of Arthur et al., (2021) and Ward & Meade, (2023).
+Respondents that are flagged as careless by such methods, but for whom no changepoint was flagged by CODERS, have likely been careless throughout the survey.
+Second, a detected changepoint does not necessarily indicate a change from attentive responding to careless responding.
+A participant may respond carelessly in the beginning of the survey and switch to responding attentively later on (cf.
+Figure 11(c) and Table 3(c)), or they may switch between careless response behaviors through the course of the survey (cf.
+Figures 11(d)11(e) and Tables 3(d)3(e)).
+For such cases, our empirical application demonstrates that CODERS—combined with a post-hoc analysis of the identified segments of responses—remains a valuable tool for gaining detailed insights into careless response behaviors.
+Third, there may be participants who alternate between periods of attentive and careless responding (Berry et al.,, 1992; Clark et al.,, 2003; Meade & Craig,, 2012).
+Our simulations indicate that CODERS remains conservative in this case and rarely flags one of the changepoints (see Appendix C.5).
+CODERS should therefore be viewed as a promising first step towards detecting periods of careless responding, with an extension for the detection of multiple changepoints being an area of further research.
+Finally, the changepoint detection procedure of Shao & Zhang, (2010) employed by CODERS is based on asymptotic arguments when the number of items is large.
+Hence, the number of items in the questionnaire is the main determinant for the statistical power of CODERS.
+For the detection of inconsistent careless responding, the number of participants plays a secondary role regarding power: the more attentive responses the autoencoder can learn from, the better the reconstruction errors can distinguish between attentive and careless responding (cf.
+Appendix C.3).
+The LSP sequence of a given participant, on the other hand, depends only on their own responses, meaning that the number of observations should not play a role concerning power in detecting invariable careless responding.
+CODERS is thus primarily designed for relatively lengthy questionnaires, and our simulations suggest that there is a gradual drop in statistical power for detecting inconsistency as survey length decreases (see Appendix C.2).
+However, partial carelessness due to survey fatigue may be less prevalent in short questionnaires (cf.
+Bowling et al.,, 2021).
+A detailed study of partially careless responding in shorter surveys is therefore beyond the scope of this paper.
 Response Times Revisited
@@
 If per-item response times are available, they can easily be incorporated into CODERS as an additional series to be used in the changepoint detection.
-If only per-page response times are available and pages are relatively short, an additional series can be obtained by assigning to each response the average page time, i.e., the time the participant spent on the page on which the item is located divided by the number of items on that page (which inspired by the page time index of Bowling et al.,, 2023).
+If only per-page response times are available and pages are relatively short, an additional series can be obtained by assigning to each response the average page time, i.e., the time the participant spent on the page on which the item is located divided by the number of items on that page (which is inspired by the page time index of Bowling et al.,, 2023).
 If a participant at some point in the survey, for instance, starts responding without reading the items, we expect a changepoint in response time towards considerably faster responses (cf.
@@
 For such survey designs with similar item length, incorporating response times into CODERS may prove fruitful, but a thorough validation is beyond the scope of this paper and therefore left for future research.
-Limitations
-Throughout this paper, we have assumed that all participants start as truthful and accurate respondents, while some of them may begin to respond carelessly from a certain item onward.
-However, in certain scenarios, this assumption may not hold true.
-First, as we have already seen in our empirical application, some participants may respond carelessly throughout all survey items (cf.
-Figure 3(d) and Table 2(d)).
-Consequently, CODERS is unlikely to flag a changepoint due to an absence of a change in behavior.
-Nevertheless, the literature has established a plethora of methods for detecting respondents who have been careless throughout the questionnaire.
-In this sense, CODERS—which is intended for detecting partial carelessness—is complementary to those existing methods.
-If one suspects that respondents who have been careless throughout all items are present, we recommend to apply established detection methods following the guidelines of Arthur et al., (2021) and Ward & Meade, (2023).
-Respondents that are flagged as careless by such methods, but for whom no changepoint was flagged by CODERS, have likely been careless throughout the entire survey.
-Second, there may be some participants who alternate between periods of attentive and careless responding (Berry et al.,, 1992; Clark et al.,, 2003; Meade & Craig,, 2012).
-In this case, CODERS may only flag one of the possibly multiple changepoints.
-CODERS should therefore be viewed as a promising first step towards detecting periods of careless responding, with an extension for the detection of multiple changepoints being an area of further research.
-Finally, the changepoint detection procedure of Shao & Zhang, (2010) employed by CODERS is based on asymptotic arguments when the number of items is large.
-Hence, CODERS is designed for lengthy questionnaires, and it is likely that the statistical power drops considerably in short questionnaires.
-However, partial careless responding due to fatigue may be less prevalent in short questionnaires (cf.
-Bowling et al.,, 2021).
-Studying the onset of careless responding in shorter surveys is therefore beyond the scope of this paper.
-Potential Alternatives to CODERS
-First, the recommendation that one should not include more than three detection items (Meade & Craig,, 2012) is fairly restrictive as carelessness may have onset at a much earlier item than the detection item, resulting in a failure to exclude a potentially large number of careless responses.
-In principle, a form of carelessness onset detection may be possible by a priori including detection items in the survey, such as bogus, instructed, or self-report items: one may consider a failed detection item (i.e., the participant provides a response that would not be expected from an attentive respondent) as an estimate of the onset of careless responding.
-However, there are a number of practical and theoretical disadvantages with this approach (see Curran,, 2016, pp.
-13–15, for a more general discussion on the caveats of detection items).
-First, the estimates of the carelessness onset may be crude at best.
-Recommendations of Meade & Craig, (2012) suggest including at most three bogus, instructed, or self-report items, spaced every 50–100 items.
-Consequently, the carelessness onset may occur much earlier than at at the detection item, leaving a potentially large number of careless responses undetected.
-Second, attentive respondents may ironically choose an unexpected response to a bogus item because they find doing so humorous (Meade & Craig,, 2012) or due to misinterpretation of the item (Curran & Hauser,, 2019).88endnote: 8For instance, Curran & Hauser, (2019) find that some respondents would agree to the popular bogus item “All my friends say I would make a great poodle” because they believe that they are loyal friends and they associate loyalty with dogs.
-Third, a careless respondent may choose the expected response by mere chance or lie in self-report items (DeSimone & Harms,, 2018; Edwards,, 2019), resulting in failure to identify this careless respondent.
-Fourth, it is unclear if, e.g., the response “somewhat disagree” to a detection item for which “strongly disagree” is expected should be considered careless (Ward & Meade,, 2023).
-Finally, it may not be possible to include detection items in the survey for legal or ethical reasons, or because the survey analyst has no control over the survey design.
-In addition, the software packages careless (Yentes & Wilhelm,, 2021) and ufs (Peters & Gruijters,, 2021) support the calculation of an intra-individual variance (IRV) across subsets of responses.99endnote: 9Specifically, this is implemented in the functions irv() from package careless and carelessReport() from package ufs.
-If a set of responses towards the end of the questionnaire exhibits a sufficiently larger (Marjanovic et al.,, 2015) or lower (Dunn et al.,, 2018) variance than an earlier set of responses, there might evidence for an onset of partially careless responding.
-First, these packages only support splitting the responses into subsets of equal length, which does not allow for accurate detection of the carelessness onset.
-Second, a more general implementation for confirming the presence of partial carelessness in arbitrary subsets would require a priori knowledge on the onset item that defines the subsets, which is unrealistic in practice.
-Third, it is unclear when a variance can be considered sufficiently low or high to infer carelessness: determining cutoff values below or above which a participant is to be flagged as careless may depend on survey design and survey content.1010endnote: 10For example, if 5-point scales are used a lower cutoff is necessary than for a 9-point scale, since IRVs will be very different for those scales.
-Instead of screening for careless responding after data collection, one may attempt to prevent such behavior during survey administration.
-Preventive measures against careless responding may entail promising rewards for responding accurately (Gibson & Bowling,, 2020), as well as threatening with punishment for careless responding or the presence of a proctor (Bowling et al.,, 2021).
-There is evidence that each of these three preventive measures can prevent or at least delay carelessness onset (Bowling et al.,, 2021; Gibson & Bowling,, 2020; Huang et al.,, 2012).
-However, either of these three preventive measures can be undesirable or unethical in practice.
-Administering proctored questionnaires is often impractical and expensive, and is typically infeasible in online studies.
-Arthur et al., (2021) and Bowling et al., (2021) advise against using threats if one wishes to maintain a positive long-term relationship with questionnaire participants, such as when there is a professional relationship between researcher and participants.
-Moreover, threats seem inappropriate in longitudinal studies or when the participants are members of a questionnaire panel (as is common in survey research), or when they are suspected to suffer from anxiety, depression, or trauma (as is common, e.g., in research on wellbeing in the workplace, or more general in psychological research).
-On the other hand, promising rewards may cross ethical boundaries in organizational surveys and in the context of employment-related decisions (Arthur et al.,, 2021).
-Notwithstanding, we stress that preventive methods against carelessness are complementary to detection methods.
-If preventative methods are appropriate, combining them with screening techniques such as CODERS may further enhance quality and reliability of the collected data for subsequent analysis.
 Discussion
-We introduced a novel method, CODERS, to identify the onset of careless responding (or absence thereof) for each participant in a given survey.
-It is motivated by literature stressing that in lengthy surveys—which are common in the behavioral and organizational sciences—a large proportion of survey participants may respond partially carelessly due to fatigue or boredom.
-CODERS is highly flexible and can combine multiple indicators for the potential onset of carelessness, such as autoencoder reconstruction errors as an indicator of inconsistency, and the proposed LongstringPattern (LSP) sequence as an indicator of invariability.
-Re-analyzing data from a seminal study on careless responding, we find evidence of partial careless responding that was hitherto undetected.
-Additionally, simulation experiments indicate excellent performance of CODERS in identifying the onset of careless responding, and highlight the importance of incorporating multiple indicators that capture different manifestations of carelessness.
+We introduced a novel method, CODERS, designed to identify the onset of careless responding (or absence thereof) for each participant in a survey.
+It is motivated by literature stressing that survey length contributes to careless responding, in the sense that lengthy surveys—which are common in the behavioral and organizational sciences—may result in a large proportion of survey participants being partially careless, for instance due to fatigue or boredom.
+CODERS is highly flexible and can combine multiple indicators for the potential onset of carelessness, such as autoencoder reconstruction errors (RE) as an indicator of inconsistency, and the proposed LongStringPattern (LSP) sequence as an indicator of invariability.
+Simulation experiments indicate excellent performance of CODERS in identifying the onset of careless responding, and highlight the importance of incorporating multiple indicators that capture different manifestations of carelessness.
+Re-analyzing data from a seminal study on careless responding, we demonstrate how CODERS reveals novel insights on partial carelessness.
+In particular when combined with a post-hoc analysis of the identified response segments, CODERS is able to uncover various types of changes in (careless) response behavior.
 While CODERS is a promising first step towards detecting periods of careless responding rather than identifying which respondents are careless, various pertinent questions remain for future research.
@@
 In addition, extending or adapting CODERS with changepoint detection methods that allow for multiple changepoints, or that retain higher power in shorter surveys, may be fruitful directions for further research.
-Whereas we apply machine learning and nonparametric techniques in our approach, other literature is focused on model-based approaches.
+Furthermore, it could be interesting to study whether autoencoder reconstruction errors, which can be thought of comparing observed responses with an expectation, are able to distinguish between inconsistent careless responding and other sources of deviations from expected response behavior, such as misunderstanding or misinterpretation of survey items.
+CODERS naturally integrates with existing techniques that address careless responding.
+First, as illustrated in our empirical application, traditional post-hoc screeners and CODERS provide complementary insights.
+The former aim at identifying which respondents are careless, whereas CODERS specifically looks for changepoints in response behavior.
+It should be noted that some implementations allow traditional screeners to be computed on subsets of responses.
+For instance, the implementation of the intra-individual variance (IRV) (Dunn et al.,, 2018; Marjanovic et al.,, 2015) in the R package careless (Yentes & Wilhelm,, 2023) can be calculated for a given number of subsets of equal length.
+Although this approach is related to the detection of periods of careless responding, it analyzes predefined subsets of responses, while CODERS aims to identify the location of a changepoint.
+Second, pre-administered detection items such as bogus, instructed, or self-report items (e.g.
+Arthur et al.,, 2021; Curran,, 2016; Ward & Meade,, 2023) may provide further validation for the findings of CODERS.
+An advantage of detection items is that they can in principle detect any type of careless responding.
+By themselves, detection items do not provide accurate estimates of the location of a change in response behavior, at least when following recommendations of Meade & Craig, (2012) to include at most three detection items, spaced every 50–100 items.
+However, detection items can strengthen the findings of CODERS: if passed detection items (i.e., responses expected from an attentive participant) fall with the found attentive period while failed detection items fall within the found careless period, there is convergent evidence of such response behavior.
+Third, various methods have been proposed in the literature to prevent careless responding in the first place, such as promising rewards for responding accurately (Gibson & Bowling,, 2020), or threatening with punishment for careless responding (Bowling et al.,, 2021).
+There is evidence that such measures can prevent or at least delay the onset of careless responding (Bowling et al.,, 2021; Gibson & Bowling,, 2020; Huang et al.,, 2012).
+However, preventive measures can be undesirable or even unethical in practice.
+For instance, Arthur et al., (2021) and Bowling et al., (2021) advise against using threats if one wishes to maintain a positive long-term relationship with survey participants, while Arthur et al., (2021) stress that promising rewards may cross ethical boundaries in organizational surveys.
+Notwithstanding, if preventative methods are appropriate, combining them with screening techniques such as CODERS may further enhance quality and reliability of the collected data for subsequent analysis.
+Whereas we apply machine learning and nonparametric techniques in CODERS, other literature is focused on model-based approaches.
 For instance, Ulitzsch et al., 2022b propose a mixture model based on item response theory that is designed to detect careless respondents, which is adapted in Ulitzsch et al., 2022a to incorporate response times.
@@
 One advantage of model-based approaches is that their estimates can be used to construct person-weights that can be used to downweight the adverse influence of potentially careless respondents in subsequent analyses, which is discussed in detail in Ulitzsch et al., 2022a ; Ulitzsch et al., 2022b and Hong & Cheng, (2019).
-However, model-based approaches typically impose strong and potentially unverifiable assumptions on the properties of attentive and careless responding which might be violated in practice.
+However, model-based approaches typically impose strong and potentially unverifiable assumptions on the properties of attentive and careless responding.
 On a final note, Arthur et al., (2021) envisage “artificial intelligence and machine learning being used to detect careless responding and response distortion in real time in the foreseeable future.” Using machine learning techniques in the form of autoencoders and nonparametric changepoint detection, our proposed method CODERS takes a step towards that future.
@@
 Throughout this paper, we have followed recent guidelines by Alfons & Welz, (2024) on the application of machine learning methods for the detection of careless responding.
-All computations were performed with R version 4.3.3 (R Core Team,, 2024) and Python version 3.10 (Van Rossum & Drake,, 2009).
-Our implementation of CODERS in the R package carelessonset is available from https://github.com/mwelz/carelessonset.
+All computations were performed with R version 4.4.2 (R Core Team,, 2024) and Python version 3.10 (Van Rossum & Drake,, 2009).
+Our implementation of CODERS in the R package carelessonset (Welz & Alfons,, 2024) is available from https://github.com/mwelz/carelessonset.
 Replication files will be made publicly available upon acceptance of this manuscript and a link will be included here.
@@
 We require the following assumptions.
+Importantly, these assumptions should not be understood as a conceptual model for careless responding, but rather as statistical prerequisites for CODERS to work best.
+In the immortal words of John von Neumann, “[the truth] is much too complicated to allow anything but approximations.”
 Assumption 1.
-The responses in $\bm{X}$ admit an $n\times s$ lower-dimensional representation, $\bm{S}$, where $s\ll p$.
+The responses in $\bm{X}$ admit an $n\times s$ lower-dimensional representation $\bm{S}$, where $s\ll p$.
 The dimension $s$ is known and corresponds to the number of constructs that the survey measures.
+This assumption is very mild in survey data since surveys typically measure multiple constructs, with the number of constructs typically known to the survey designer.
+For instance, the $p=300$ item representation of the NEO-PI-R inventory (Goldberg,, 1999) measures $s=30$ constructs.
 Assumption 2.
 The survey that generated $\bm{X}$ is reliable in the sense that if all participants responded accurately, $\bm{X}$ would accurately measure all constructs.
-Assumption 3.
-The items in the survey are in a randomized order.
-Assumption 4.
-All participants begin the survey as attentive respondents by providing accurate and truthful responses.
-As the survey progresses, some (possibly none or all) participants start responding carelessly and continue to do so for the remainder of the survey.
-Assumption 5.
-The onset of partial carelessness is characterized by a changepoint in one or more indicators. We use autoencoder reconstruction errors and LongstringPattern (LSP) sequences as indicators.
-Assumption 1 is very mild in survey data because surveys typically measure multiple constructs and the number of constructs is typically known.
-For instance, the $p=300$ item representation of the NEO-PI-R inventory (Goldberg,, 1999) measures $s=30$ constructs.
-Assumption 2 is again mild in survey data.
-There exist well-established and reliable survey measures for a large variety of variables.
+This assumption is again mild in survey data.
+There exist well-established and reliable survey measures for a large variety of constructs.
 For instance, the International Personality Item Pool (IPIP; Goldberg,, 1999) is a pool of more than 250 personality scales.
@@
 150).
-The reliability of a measurement can be estimated in multiple ways, such as Cronbach’s alpha, the omega coefficient and variations thereof, as well as the $H$ coefficient (see, e.g., McNeish,, 2018, for a description of these and other reliability measures).
-Hence, Assumption 2 is satisfied when the survey data were collected with measures of high reliability.
-Assumption 3 is common practice in surveys in the social and behavioral sciences.
-The randomized order of items prevents that recurring response patters emerge as a result of the survey design rather than due to invariable careless responding.
-Assumption 4 is discussed in the section on limitations, and Assumption 5 is motivated in detail in the section introducing the proposed methodology.
-Appendix B Technical Details on CODERS
-B.1 Details on Autoencoders
-B.1.1 Network Architecture and Estimation
+The reliability of a measurement can be estimated in multiple ways, such as Cronbach’s alpha, McDonald’s omega and variations thereof, as well as the $H$ coefficient (see, e.g., McNeish,, 2018, for a description of these and other reliability measures).
+Hence, this assumption is satisfied when the survey data are collected with measures of high reliability.
+Assumption 3.
+The survey design precludes responses in (fixed) recurring patterns, meaning that such patterns are indicative of invariable careless responding.
+One way to satisfy this assumption is through randomization of the survey items, which is common practice in surveys in the social and behavioral sciences.
+If it is desirable to keep items of the same construct together to reduce the burden on respondents, as is often the case in organizational surveys, one may randomize the order of items within each construct, and randomize the order of constructs (or groups of constructs, if related constructs should be presented together).
+The use of negatively keyed items may also be beneficial.
+Assumption 4.
+Participants begin the survey as attentive respondents by providing accurate and truthful responses.
+As the survey progresses, some (possibly none or all) participants start responding carelessly and continue to do so for the remainder of the survey.
+Under this assumption, a detected changepoint (see also the next assumption) naturally divides a participant’s responses into an attentive segment before the changepoint and a careless segment from the changepoint onward.
+In practice, a changepoint may characterize different changes in response behavior.
+For instance, a participant may engage in careless responding in the beginning of the survey but later switch to attentive responding, or they may switch from one type of careless responding to another (e.g., from inconsistency to invariability).
+Via a post-hoc analysis of the identified response segments, CODERS remains useful for uncovering insights on such changes in response behavior, as highlighted in the empirical application with selected cases (cf.
+Figures 11(c)11(e) and Tables 3(c)3(e) in the main text).
+It is also conceivable that a respondent may switch repeatedly between attentive and careless responding.
+For instance, they may start attentively, then be distracted for a certain period of time, but switch back to attentive responding once the distraction goes away.
+CODERS in its current form is not designed to pick up such behavior, but an extension using multiple changepoint detection may be a fruitful venue for future research.
+Assumption 5.
+The onset of partial carelessness is characterized by a changepoint in one or more indicators. We use autoencoder reconstruction errors (RE) and LongStringPattern (LSP) sequences as indicators.
+A detailed motivation for this assumption is provided in the section introducing CODERS.
+Appendix B Details on CODERS
+B.1 Motivating Novel Methodology
+Before we present a technical description of CODERS, we believe that it is instructive to first take a step back and motivate the application of novel techniques for the task of detecting partially careless responding rather than using existing post-hoc screeners such as psychometric antonym, personal reliability, and the longstring index (see e.g., Alfons & Welz,, 2024, for a recent overview of traditional screeners).
+A primary motivation for devising new methodology is that most existing screeners compute a single carelessness score for each participant.
+While this can work very well for distinguishing careless from attentive respondents, having a singular score for every respondent does not provide much information beyond the attentive-careless classification.
+As outlined in the main text, recent literature suggests that partially careless responding might be a common occurrence.
+As such, it might be more informative to know when a given respondent starts responding carelessly (if at all) rather than a dichotomous attentive-careless distinction.
+For instance, this could be useful for knowing when the responses of a given respondent become unreliable so that one must not drop all of their responses in a primary analysis.
+In order to obtain a granular estimate of carelessness onset on the participant level, we require a carelessness score for every item for every respondent rather than a singular per-participant carelessness score.
+While some traditional screeners can be somewhat tweaked in this direction (for instance, one could calculate separate intra-individual response variances for separate sets of items; see e.g., Dunn et al.,, 2018; Marjanovic et al.,, 2015), accurate estimation of the onset of carelessness necessitates new methodology.
+As such, CODERS is based on two different carelessness scores on the participant-item level to capture a wide variety of different types of careless responding.
+The first dimension, reconstruction errors (RE) obtained via an auto-associative neural network (autoencoder) (Kramer,, 1992), is intended to measure inconsistent careless responding, such as choosing response categories at random, on the participant-item level.
+Autoencoders are described in detail in Section B.2.
+The second dimension, a LongStringPattern (LSP) sequence, is a generalization of the famous longstring index (Johnson,, 2005) that measures on the participant-item level if a given response is part of a content-independent recurring response sequence, such as $-2-1-2-\dots-1-$ 2 - ⋯ - 1 - 2.
+If so, it allows one to trace back when the recurring sequence has started, thereby providing evidence for the onset of invariable carelessness.
+LSP is explained in detail in Section B.3.
+B.2 Details on Autoencoders
+B.2.1 Further Intuition
+Assigning a participant-item score that measures how inconsistently a participant has responded to the item in question arguably is a challenging task.
+Since different attentive participants may respond differently to a given item (between-participant heterogeneity), one needs to somehow understand how attentive responding manifests on the participant-level so that a possible switch to inconsistent responding can be detected.
+Otherwise, there would be the risk of confusing attentive responding for inconsistent responding, and vice versa.
+We argue that an auto-associative neural network (autoencoder) is useful for learning within-participant attentive response structures.
+An autoencoder as used in CODERS essentially predicts the response of each respondent to each item.
+The underlying idea is that if the predicted response is notably different from the given response, the participant has given a response that is unexpected, judging from their previous response behavior.
+While an occasional unexpected response is not out of the ordinary, having multiple such unexpected responses occur consecutively may suggest that the participant has shifted from predictable responding to unpredictable responding, which might be viewed as indication of an onset of inconsistent responding.
+Since we assume that every participant has provided attentive responses at the beginning of the questionnaire (see Assumption 4), it should be conceptually possible to learn how attentive responding manifests for each participant so that somewhat accurate response predictions can be made up to carelessness onset.
+The autoencoder learns attentive response behavior by exploiting a key feature of questionnaires, namely that multiple items measure the same construct through psychometric scales.
+Since one usually knows how many scales are measured by the survey at hand, the information contained in a given participant’s responses should be representable in much fewer dimensions than the total number of items (see Assumption 1).
+Thus, the task of learning attentive response behavior can be seen as finding a good lower-dimensional representation for every participant’s given responses.
+Formally, the autoencoder minimizes the participant-item response prediction error over all participants and items.
+However, in the process of making predictions, the autoencoder enforces to first find a lower-dimensional representation of each participant’s given responses, and then to make a response prediction based on that lower-dimensional representation.
+In other words, the autoencoder first needs to compress $p$ dimensions into $s\ll p$ dimensions, and then decompress again that lower-dimensional representation to obtain $p$-dimensional predictions.
+The potential presence of inconsistent careless responses should not much affect the quality of the lower-dimensional representation—unless carelessness onsets very early—because inconsistent responding is characterized by its random-like nature, meaning that such responses are not informative for learning the lower-dimensional representation.
+We stress that response predictions based on a lower-dimensional representation could also be obtained through principal component analysis (PCA).
+In fact, autoencoders turn out to be a generalization of PCA that, unlike PCA, is not based on linearity assumptions (Kramer,, 1991).
+Since we wish to refrain from possibly restrictive assumptions, we choose autoencoders over PCA to construct lower-dimensional representations of the given responses.
+We are now ready to dive into the formal definition of the autoencoder.
+B.2.2 Network Architecture and Estimation
 Denote by $x_{ij}$ the response of the $i$-th participant to the $j$-th survey item.
@@
 We refer to Chapter 8.5 in Goodfellow et al., (2016) for details.
-B.1.2 Modeling Choices
+B.2.3 Modeling Choices
 Following Kramer, (1992), we specify $M=5$ layers in the autoencoder.
@@
 For fitting the autoencoder, we use the stochastic gradient descent algorithm with a batch size of 10, a learning rate of 0.0001, and 100 epochs.
-In the presudo-Huber loss (B.2), we set constant $\delta=1$.
-B.2 LongstringPattern Algorithm
+In the pseudo-Huber loss (B.2), we set constant $\delta=1$.
+B.3 LongStringPattern Algorithm
 Formal algorithms for computing the $l$-pattern and LSP sequences are provided in Algorithms 1 and 2, respectively.
@@
 Algorithm 2 LongStringPattern($\{x_{j}\}_{j=1}^{p},l_{max}$)
-B.3 Details on Changepoint Detection
+B.4 Details on Changepoint Detection
 Let $\{\bm{Y}_{j}\}_{j=1}^{p}$ be a series of length $p$ consisting of $d$-dimensional random variables.
@@
 Doing so results in negligible but nonzero variation, which renders $\bm{V}_{p}(k)$ invertible.
-Appendix C Details on the Empirical Application
-C.1 Data Preprocessing
-The data of Johnson, (2005), provided at https://osf.io/sxeq5/, contain the following information for each of the the $n=20,993$ respondents.
-First, per-item responses to $p=300$ IPIP-NEO-300 items, five domain scores (directly calculated from the responses), scores for psychometric antonyms and personal reliability, as well as respondent sex and age.
-The responses in the provided data have already underwent some preprocessing that is described in detail the accompanying README file.
-Specifically, responses to negatively-worded items were reverse-coded, that is, a given response of 5 was recoded to 1, 4 was recoded to 3, and so on.
-Hence, in order to recover the original responses, we reverted this recoding.
-Furthermore, a total of 33,691 individual responses are missing (0.53% of all responses).
-Note that Johnson, had already removed participants with too many missing responses from the provided data set, hence there are relatively few missing responses for the remaining participants.
-Due to the low number of missing responses, we treat them as a separate answer category by coding them as 0.
-C.2 Additional Results
-Figure C.1 visualizes the detected onset items for the respondents that are flagged as careless by CODERS at significance levels 0.1%, 0.5%, and 1%.
-Figure C.1: Estimated onset of careless responding for participants flagged by CODERS in the data of Johnson, (2005).
+Appendix C Additional Simulations
+C.1 Additional Results for the Main Design
+In addition to CODERS, we apply three traditional methods for the detection of careless respondents in the simulation design from the main text, which is inspired by the data collected by Johnson, (2005): personal reliability, psychometric antonym, and the longstring index.
+We choose these three methods since they were also applied in the study of Johnson, (2005).
+Following Johnson,, we reverse psychometric antonym scores, and we flag respondents as careless whose personal reliability and psychometric antonym scores are lower than $0.3$ and $-0.03$, respectively.
+For the longstring index, we flag a respondent as careless if its value exceeds 6 (i.e., $>6$), which is a cutoff value suggested by Niessen et al., (2016).
+Figure C.1 visualizes the false positive rate (FPR) and the false negative rate (FNR) of the three methods for the different carelessness onset regimes (in separate columns).
+Figure C.2 further shows the false negative rates for the four types of partially careless respondents (in separate rows).
+The longstring index performs perfectly at its intended task of detecting straightlining, whereas it hardly detects any random responding or pattern responding (which is expected).
+Furthermore, it performs better than expected at identifying extreme responding.
+Since the survey is lengthy and those respondents choose randomly from only the two extreme answer categories, there is a relatively high probability (depending on the onset) that they select a long string of the same answer category somewhere by chance.
+It turns out that personal reliability flags hardly any respondents, resulting in false positive rates close to 0 and false negative rates close to 1.
+Psychometric antonym yields inflated false positive rates, while the false negative rates are still relatively high.
+Interestingly, its performance is fairly constant across the different types of carelessness.
+As expected, the false negative rate of the traditional methods in general worsens the later the onset of careless responding (while that of CODERS deteriorates the earlier the onset, see Figures 7 and 8 in the main text).
+Figure C.1: Performance measures for the traditional methods used by Johnson, (2005).
 Note.
-Results for different significance levels $\alpha\in\{0.01,0.05,0.1\}$ are presented in separate panels.
-The $x$-axis shows the indices of the participants flagged as partially carelessness by CODERS.
-The participants are sorted in ascending order of the estimated carelessness onset item, which is in turn depicted on the $y$-axis.
-The dashed horizontal line represents the average location of carelessness onset across the flagged participants.
-Figure C.2 provides a histogram as well as a density estimate for the per-respondent average reconstruction errors.
-We can see that the density has a long right tail, which might be due to careless respondents.
-We highlight the value 0.064 by a vertical dotted line, which is the average reconstruction error of the participant from Figure 3(d) and Table 2(d).
-This value is in the 99.7% percentile of all $n=20,993$ average reconstruction errors.
-Given the extremely high average reconstruction error and the fact that this respondent was flagged by both benchmark methods, it is likely that the participant responded carelessly throughout the survey, explaining why CODERS did not flag a changepoint.
-Figure C.2: Histogram of average reconstruction errors for the data of Johnson, (2005).
+The $x$-axis displays various prevalence levels.
+False positive rates in attentive respondents (top row) and false negative rates in partially careless respondents (bottom row) for different onset regimes (columns) are averaged across 100 repetitions.
+Figure C.2: False negative rate of the traditional methods used by Johnson, (2005) for different types of partially careless respondents.
+Note. The $x$-axis displays various prevalence levels. Results for different types of careless responding (rows) in different onset regimes (columns) are averaged across 100 repetitions.
+C.2 Decreasing Survey Length
+Recall that the changepoint detection method of Shao & Zhang, (2010) used in CODERS is based on asymptotic arguments for the number of items being large (see Appendix B.4).
+It therefore seems plausible that the ability of CODERS to accurately identify partial carelessness may decline as survey length decreases.
+C.2.1 Data Generation
+We use the same simulation design of the main text, but we focus on the baseline onset regime and instead decrease the number of items $p$ by gradually removing sets of items.
+Specifically, for a given simulated dataset, we remove all items belonging to randomly selected personality traits.
+There is a total of five traits, each of which is measured by six facets comprising ten items each (300 items in total).
+Hence, removing a single trait from the measurement will remove $6\times 10=60$ items.
+We drop up to four traits, resulting in surveys comprising a total of $p\in\{240,180,120,60\}$ items.
+C.2.2 Results
+Figure C.3 plots the false positive rates for the different survey lengths (in separate columns).
+As before, CODERS rarely incorrectly flags attentive respondents as careless at any of the considered significance levels, although there is some overrejection for the shortest survey with $p=60$.
+Figure C.4 visualizes the false negative rate (top row) and the mean absolute error of onset estimation (bottom row) for careless respondents across the different surveys lengths (in separate columns).
+At a given significance level, the false negative rate slightly increases with decreasing survey length compared to the results from the main text.
+For $p=60$, the false negative rate reaches about 30%–40%, depending on the significance level.
+In addition, the location of the onset of carelessness is still estimated highly accurately throughout the investigated survey lengths, on average being off by only about one or two items.
+Figure C.3: False positive rate of CODERS for attentive respondents across decreasing survey length.
+Note. The $x$-axis displays various prevalence levels of partially careless respondents. Results for different survey lengths are shown in separate columns, averaged across 100 repetitions.
+Figure C.4: Performance measures of CODERS for partially careless respondents across decreasing survey length.
 Note.
-The $x$-axis contains the per-respondent average reconstruction errors, $\textsf{RE}_{i}=\frac{1}{p}\sum_{j=1}^{p}\textsf{RE}_{ij}$, $i=1,\dots,n$, where averages are taken over the $p=300$ items.
-On the $y$-axis, a histogram of the average reconstruction errors for $n=20,993$ respondents is displayed together with an empirical Gaussian kernel density.
-The vertical dotted line denotes the value of $\textsf{RE}_{i}$ for the participant in Figure 3(d) that was flagged by both benchmark methods, but not CODERS.
-This value, 0.064, lies in the 99.7% percentile of all average reconstruction errors.
-Appendix D Details and Additional Simulations
-D.1 Details on the Adjusted Rand Index
-In the context of our simulations, the adjusted Rand index (ARI) for a given respondent is computed as
-$ARI=\begin{cases}1&\text{ if the respondent is attentive and no changepoint isflagged},\\ \frac{A-BC/D}{(B+C)/2-BC/D}&\textnormal{ otherwise,}\end{cases}$
-where
-$A$ $=\binom{p_{11}}{2}+\binom{p_{00}}{2}+\binom{p_{10}}{2}+\binom{p_{01}}{2},$
-$B$ $=\binom{p_{00}+p_{01}}{2}+\binom{p_{10}+p_{11}}{2},$
-$C$ $=\binom{p_{00}+p_{10}}{2}+\binom{p_{01}+p_{11}}{2},$
-$D$ $=\binom{p}{2},$
-with $p_{11}$ denoting the number of careless responses that are correctly identified as careless, $p_{00}$ the number of attentive responses that are correctly identified as attentive, $p_{10}$ the number of careless responses that are incorrectly identified as attentive, and $p_{01}$ the number of attentive responses that are incorrectly identified as careless.
-Intuitively, $p_{11}$ and $p_{00}$ measure correct classification, whereas $p_{01}$ and $p_{10}$ measure misclassification.
-Observe that the ARI assumes value 1 if there is no misclassification, which is the maximum value it can take.
-Hence, for accurate estimates of the carelessness onset item, the ARI will be close to value 1.
-For very inaccurate estimates of carelessness onset, the ARI will be close to 0, or potentially even below that.
-It is unclear how to interpret negative ARI values, but such values only occur rarely.
-We refer to Steinley, (2004) for a detailed discussion of the properties of the ARI.
-If a changepoint is flagged in an attentive respondent, the ARI takes value 0.
-Similarly, if no changepoint is flagged in an attentive respondent, the ARI takes value 1.
-Hence, when averaging the ARI over attentive respondents, the ensuing average can be interpreted as the proportion of attentive respondents who are correctly identified as attentive.
-On the other hand, averages of ARIs calculated on partially careless respondents do not have a direct interpretation other than higher values indicating better performance.
-D.2 Additional Simulation Results
-We investigate the effect of various choices in our simulations.
-As the results are similar to those in the main paper other than some shifts in size or power, we only present plots of the average ARI for attentive and careless respondents (similar to Figure 5) in order to keep the appendix concise.
-D.2.1 Choosing a Different Significance Level
-Figure D.1: ARI of CODERS at a significance level of 0.5%.
-Note. The ARI is averaged over fully attentive respondents and partially careless respondents (columns) for different carelessness onset settings (rows).
-Figure D.2: ARI of CODERS at a significance level of 1%.
-Note. The ARI is averaged over fully attentive respondents and partially careless respondents (columns) for different carelessness onset settings (rows).
-We use the baseline simulation design from the main text, but we set the significance level in CODERS to 0.5% and to 1%, respectively.
-The corresponding results are shown in Figures D.1 and D.2.
-Compared to the significance level of 0.1% in the main text (cf.
-Figure 5), there is a gradual increase in power in the changepoint detection, yielding a small increase in the ARI across prevalence levels for careless respondents.
-Further investigation suggests that the increase in ARI stems mainly from better performance in detecting random and extreme responding (results not shown to keep the appendix at a reasonable length).
-However, for high prevalence levels from 40% onwards, this comes at a cost of too many false positives in attentive respondents, translating to a substantial decrease in ARI for the highest prevalence levels.
-D.2.2 Varying Correlation Structures
-In this variation of the simulation design, we keep the same settings as in the baseline design from the main text, but we vary the correlation structure in the following ways:
-  1. 1.
-    Randomly sampled moderate correlations.
-In every repetition, correlations of item pairs from the same scale are randomly drawn from the interval $[0.4,0.6]$.
-Items from different scales remain uncorrelated.
-  2. 2.
-    Randomly sampled strong correlations.
-In every repetition, correlations of item pairs from the same scale are randomly drawn from the interval $[0.6,0.8]$.
-Items from different scales remain uncorrelated.
-  3. 3.
-    Randomly sampled grouped correlation structure.
-Multiple scales form groups so that there are 5 groups consisting of 6 scales each.
-Since there are 10 items in each scale, there are $6\times 10=60$ items in each group.
-In each repetition, correlations of item pairs from the same scale are randomly drawn from the interval $[0.6,0.8]$, correlations of item pairs in different scales of the same group are randomly drawn from the interval $[0,0.2]$, and items from different groups are uncorrelated.
-While the baseline simulation design uses a relatively simple correlation structure (items from the same scale have a fixed correlation of 0.7 before recoding, items from different scales are uncorrelated), the three variations above introduce more randomness into the design.
-Due to this randomness, it may occur that a sampled correlation matrix is not positive definite.
-In such a case, we find the closest positive definite correlation matrix via the function nearPD() in the R package Matrix (Bates et al.,, 2023).
-Figure D.3: ARI of CODERS for randomly sampled moderate correlations.
-Note. The ARI is averaged over fully attentive respondents and partially careless respondents (columns) for different carelessness onset settings (rows).
-Figure D.4: ARI of CODERS for randomly sampled strong correlations.
-Note. The ARI is averaged over fully attentive respondents and partially careless respondents (columns) for different carelessness onset settings (rows).
-Figure D.5: ARI of CODERS for the randomly sampled grouped correlation structure.
-Note. The ARI is averaged over fully attentive respondents and partially careless respondents (columns) for different carelessness onset settings (rows).
-Figures D.3D.5 display the results corresponding to the different correlation structures at significance level 0.1%.
-The results of the design with randomly sampled strong correlations (Figure D.4) and the grouped correlation structure (Figure D.5) are essentially equivalent to the results of the baseline design with fixed correlation (Figure 5 in the main text).
-The design with randomly sampled moderate correlations (Figure D.3) also yields similar results, save for a small power drop in the changepoint detection (resulting in a lower ARI for careless respondents) and fewer false positives in extreme carelessness prevalence levels (yielding a higher ARI in attentive respondents).
-D.2.3 Varying Sample Size
-This variation of the simulation design again uses the same settings as in the baseline design from the main text, but now we change the sample size to $n=250$ and $n=1000$, respectively.
-Figures D.6 and D.7 visualize the corresponding results at significance level 0.1%.
-They are similar to the results for $n=500$ in the main text (cf.
-Figure 5), other than a small decrease ARI in careless respondents for $n=250$ as well as a small increase for $n=1000$.
-Figure D.6: ARI of CODERS for sample size $n=250$.
-Note. The ARI is averaged over fully attentive respondents and partially careless respondents (columns) for different carelessness onset settings (rows).
-Figure D.7: ARI of CODERS for sample size $n=1000$.
-Note. The ARI is averaged over fully attentive respondents and partially careless respondents (columns) for different carelessness onset settings (rows).
+The $x$-axis displays various prevalence levels.
+The top row shows results for the false negative rate, while the bottom row visualizes the accuracy of onset estimation via the mean absolute error.
+Results for different survey lengths are shown in separate columns, averaged across 100 repetitions.
+Figure C.5: False negative rate of CODERS for different types of partially careless respondents across decreasing survey length.
+Note. The $x$-axis displays various prevalence levels. Results for different types of careless responding (rows) in different survey lengths (columns) are averaged across 100 repetitions.
+Figure C.6: Mean absolute error of carelessness onset estimation with CODERS for different types of partially careless respondents across decreasing survey length.
+Note. The $x$-axis displays various prevalence levels. Results for different types of careless responding (rows) in different survey lengths (columns) are averaged across 100 repetitions.
+Figures C.5 and C.6 show the false negative rate and the mean absolute error of onset estimation, respectively, grouped by carelessness type (rows) for each considered survey length (columns).
+In essence, the slight increase in false negative rates seems to be primarily caused by fewer inconsistent careless respondents being detected, whereas the performance for invariable careless responding stays near-perfect.
+For $p=60$, in particular very few random respondents are correctly identified, but the performance is still acceptable for extreme respondents.
+We conclude that, as expected, the power of CODERS for detecting partially careless respondents drops as the survey length decreases.
+Yet, the drop in power is small and primarily caused by a low detection performance for random respondents.
+It is conceivable that reconstruction errors are a noisier measurement of inconsistency (compared to LSP as a measurement of invariability), as the autoencoder has less information to learn attentive responding behavior in shorter surveys, explaining why the loss of power is primarily visible for random responding.
+Nevertheless, if a changepoint is flagged, its accuracy is nearly unchanged at excellent levels.
+That is, even though CODERS has somewhat less power in identifying partially careless respondents, the onset of carelessness in the flagged respondents is sill accurately estimated.
+Furthermore, the power loss can be partially alleviated by choosing a less conservative significance level, such as 0.5% or 1%.
+C.3 Additional Sample Sizes
+Although the number of items in the survey may be the primary driver regarding power of the changepoint detection test used in CODERS, the autoencoder may also benefit from more information in the form of a larger sample size in order to learn attentive response behavior (and conversely suffer under a smaller sample size).
+C.3.1 Data Generation
+[…]
+C.3.2 Results
+Figures C.7C.10 visualize the performance measures for $n=248$ and $n=1,000$ simulated participants (in separate columns).
+The results are similar to those in the main text, with the overall FNR in partially careless respondents increasing only slightly for the smaller sample size and decreasing only slightly for the larger sample size.
+For random respondents, however, the increase in false negatives in the smaller sample size is notable, but so is the decrease in FNR in the larger sample size (see Figure C.9, top row).
+This can likely be explained by the amount of information the autoencoder has available in order to learn to distinguish between attentive responding and random responding.
+It is therefore conceivable that the power for detecting random respondents may improve further in even larger sample sizes.
+Figure C.7: False positive rate of CODERS for attentive respondents across additional sample sizes.
+Note. The $x$-axis displays various prevalence levels of partially careless respondents. Results for different sample sizes are shown in separate columns, averaged across 100 repetitions.
+Figure C.8: Performance measures of CODERS for partially careless respondents across additional sample sizes.
+Note.
+The $x$-axis displays various prevalence levels.
+The top row shows results for the false negative rate, while the bottom row visualizes the accuracy of onset estimation via the mean absolute error.
+Results for different sample sizes are shown in separate columns, averaged across 100 repetitions.
+Figure C.9: False negative rate of CODERS for different types of partially careless respondents across additional sample sizes.
+Note. The $x$-axis displays various prevalence levels. Results for different types of careless responding (rows) in different sample sizes (columns) are averaged across 100 repetitions.
+Figure C.10: Mean absolute error of carelessness onset estimation with CODERS for different types of partially careless respondents across additional sample sizes.
+Note. The $x$-axis displays various prevalence levels. Results for different types of careless responding (rows) in different sample sizes (columns) are averaged across 100 repetitions.
+C.4 Randomly Choosing from the Middle Answer Categories
+Recall that the simulation design in the main text considers four types of careless responding: random responding, extreme responding, pattern responding, and straightlining.
+However, there may be other relevant types of carelessness beyond those four, including ones that are extremely difficult to distinguish from attentive response behavior.
+C.4.1 Data Generation
+As before, we take the simulation design from the main text with the baseline onset regime, but instead of the four aforementioned types of careless responding, we generate a distinct type that we refer to as middling.
+Middling is characterized by randomly choosing between the three middle answer categories (out of five).
+As such, middling carelessness can be seen as an inconsistent response type complementing extreme responding.
+The major difference to extreme carelessness is that central response options are more frequently chosen by attentive respondents in our simulation design.
+In fact, with the chosen marginal distributions and correlation structure (see Figure 5), there is a lot of natural variation among the central response options in attentive responses.
+Hence, it is difficult to systematically distinguish middling carelessness from attentive responding, but there is also the risk that attentive respondents might be mistaken for middling careless respondents.
+C.4.2 Results
+Figures C.11 and C.12 visualizes the performance of CODERS when partially careless respondents are of the middling type.
+One the one hand, CODERS still almost never incorrectly flags an attentive respondent as partially careless.
+On the other hand, it identifies very few middling respondents with a false negative rate between about 85% and 95%, depending on the significance level.
+Hence, these simulations further demonstrate that CODERS requires inconsistent responding to be sufficiently different from attentive responding in order to distinguish between the two.
+Figure C.11: False positive rate of CODERS for attentive respondents in the presence of partially careless respondents that randomly choose from the middle answer categories.
+Note. The $x$-axis displays various prevalence levels of partially careless respondents. Results are averaged across 100 repetitions.
+Figure C.12: Performance measures of CODERS for partially careless respondents that randomly choose from the middle answer categories.
+Note.
+The $x$-axis displays various prevalence levels.
+The top row shows results for the false negative rate, while the bottom row visualizes the accuracy of onset estimation via the mean absolute error.
+Results are averaged across 100 repetitions.
+In order to further highlight the difficulty of detecting middling carelessness, we again apply the traditional methods personal reliability, psychometric antonym, and the longstring index.
+The false positive rates and false negative rates are shown in Figure C.13.
+As in the simulation design from the main text, personal reliability flags hardly any respondents, resulting in false positive rates close to 0 and false negative rates close to 1.
+The performance of the longstring index is similar, which is not surprising because it was designed for detecting straightlining (invariable carelessness) rather than inconsistent responding like middling.
+Psychometric antonym identifies more middling respondents than CODERS with a false negative rate of about 65%–75%, but it exhibits an inflated false positive rate of up to about 15%.
+Figure C.13: Performance measures of the traditional methods for partially careless respondents that randomly choose from the middle answer categories.
+Note.
+The $x$-axis displays various prevalence levels.
+False positive rates in attentive respondents (top row) and false negative rates in partially careless respondents (bottom row) are averaged across 100 repetitions.
+C.5 Temporary Carelessness
+It seems plausible that more than one changepoint may occur in a participant’s response behavior.
+For instance, a respondent may start the survey attentively, but become careless at some point due to being distracted, only to start responding attentively again once the distraction vanishes.
+C.5.1 Data Generation
+We use the same simulation design from the main text, but instead of simulating only one changepoint, we simulate temporary carelessness as follows.
+We randomly sample the first changepoint from the item set $\{30,31,\dots,149,150\}$ (between 10% and 50% of all items) and the second changepoint from the item set $\{151,152,\dots,269,270\}$ (between 50% and 90% of all items).
+Between these two changepoints, we generate careless responses using the same four types of carelessness from the main text.
+C.5.2 Results
+Figures C.14 and C.15 visualize the performance of CODERS and Figures C.16 and C.17 that of the traditional methods.
+Clearly, CODERS rarely flags any participants, with a false positive rate close to 0 and a false negative rate close to 1 (we therefore omit separate results for the four types of partially careless responding).
+The same behavior is observed for personal reliability.
+Psychometric antonym achieves a better false negative rate (about 70%–85%) but struggles with many false positives (up to about 15%).
+The longstring index yields somewhat better performance, with a relatively low false positive rate and a false negative rate around 65%.
+As in the simulation design from the main text, it successfully captures straightliners as well as some extreme respondents (see Appendix C.1 for a more detailed discussion).
+In summary, CODERS is simply not designed for detecting more than one changepoint in respondent behavior, such that an extension to multiple changepoint detection may be a fruitful direction for further research.
+Figure C.14: False positive rate of CODERS for attentive respondents in case of temporary carelessness.
+Note. The $x$-axis displays various prevalence levels of partially careless respondents. Results are averaged across 100 repetitions.
+Figure C.15: Performance measures of CODERS for partially careless respondents in case of temporary carelessness.
+Note.
+The $x$-axis displays various prevalence levels.
+The top row shows results for the false negative rate, while the bottom row visualizes the accuracy of onset estimation via the mean absolute error.
+Results are averaged across 100 repetitions.
+Figure C.16: Performance measures for the traditional methods in case of temporary carelessness.
+Note.
+The $x$-axis displays various prevalence levels.
+False positive rates in attentive respondents (top row) and false negative rates in partially careless respondents (bottom row) are averaged across 100 repetitions.
+Figure C.17: False negative rate of the traditional methods for different types of partially careless respondents in case of temporary carelessness.
+Note. The $x$-axis displays various prevalence levels. Results for different types of careless responding are shown in separate columns, averaged across 100 repetitions.
+C.6 Organizational Survey
+Until here, all simulations are based on a measurement of the Big Five personality traits.
+To study how CODERS fares in surveys of different nature, we further apply a simulation design based on a study from organizational research.
+C.6.1 Data Generation
+We simulate a hypothetical survey inspired by a study conducted among university students working in teams on a project for a Strategic Management course.
+The measurement includes various individual- and team-level constructs, among others the HEXACO-60 personality assessment of Ashton & Lee, (2009), the goal orientation scales of Vandewalle, (1997), the transactive memory systems scales of Lewis, (2003), the intra-group conflict scales of Jehn, (1995), the intra-unit justice climate scales of Li & Cropanzano, (2009), a team commitment scale based on Mowday et al., (1979), and the team performance scale of Hackman, (1986).
+In total, the hypothetical survey includes $p=198$ items measuring $q=34$ constructs (which we set as the number of nodes in the bottleneck layer of the autoencoder in CODERS).
+Figure C.18: Correlation matrix and marginal response distributions used in the simulated organizational survey.
+(a) Correlation matrix of the $p=198$ items. A more saturated blue (red) in a cell corresponds to a stronger positive (negative) correlation between the corresponding items.
+(b) Marginal response distributions of the $p=198$ items ($x$-axis), each of which has five response options ($y$-axis).
+A darker shade of gray indicates a higher probability of the corresponding response option.
+Note.
+The correlation matrix and the marginal response distributions are calculated from empirical data.
+Of the $p=198$ items, 152 are positively scored and 46 are negatively scored.
+The items in both plots are presented in an order such that items measuring the same construct are adjacent.
+Although team-level constructs are measured, we simplify the data generating process and simulate individual-level data using the correlation matrix and marginal distributions displayed in Figure C.18, which are based on empirical data from 322 students (after removing 29 respondents that contain missing values).
+Unfortunately, we are not at liberty to share the original data.
+Of the $p=198$ items, 152 are positively scored and 46 are negatively scored.
+The different constructs are clearly visible along the diagonal of the correlation matrix, as are various relationships between many of the measured constructs in the off-diagonal parts (cf.
+also Alfons et al., 2022a, ; Alfons et al., 2022b, ).
+We randomize the order of the items in a way that ensures that items measuring the same construct remain adjacent, which is better suited for surveys in organizational contexts.
+That is, we first randomize the order of the items within each construct, and we randomize the order of (groups of related) constructs.
+Otherwise, we keep the same design as in the simulations from the main text.
+In particular, we simulate responses for $n=500$ participants, using the same settings for generating partially careless respondents.
+C.6.2 Results
+Figure C.19 shows the false positive rates for attentive respondents.
+Compared to the simulations from the main text, false positive rates are inflated for high prevalence levels, but they remain reasonably close to the nominal level while carelessness prevalence does not exceed 40% (in the early onset regime) or 30% (for the other onset regimes).
+A possible explanation for this finding may be that many measured constructs do not use negatively-scored items while others do, such that response behavior may be naturally different in some parts of the survey than in others.
+Figure C.19: False positive rate of CODERS for attentive respondents in the simulated organizational survey.
+Note. The $x$-axis displays various prevalence levels of partially careless respondents. Results for different onset regimes are shown in separate columns, averaged across 100 repetitions.
+In Figures C.20C.22, the performance measures for (different types) of partially careless respondents are displayed.
+The overall false negative rate is similar to that of the simulations from the main text, despite the shorter survey length.
+But interestingly, there are slightly fewer false negatives for random responding and slightly more false negatives for extreme responding.
+This may be explained by differences in response distributions, with higher skewness in some of the measurements compared to the data by Johnson, (2005) (cf.
+Figure C.18 with Figure 5 from the main text).
+In addition, the onset of careless responding is estimated with similar accuracy, with an error of about two to three items averaged across the different types of partially careless responding.
+All in all, these results are an indication that the performance of CODERS may not depend too much on the institutional setting of the survey.
+Figure C.20: Performance measures of CODERS for partially careless respondents in the simulated organizational survey.
+Note.
+The $x$-axis displays various prevalence levels.
+The top row shows results for the false negative rate, while the bottom row visualizes the accuracy of onset estimation via the mean absolute error.
+Results for different onset regimes are shown in separate columns, averaged across 100 repetitions.
+Figure C.21: False negative rate of CODERS for different types of partially careless respondents in the simulated organizational survey.
+Note. The $x$-axis displays various prevalence levels. Results for different types of careless responding (rows) in different onset regimes (columns) are averaged across 100 repetitions.
+Figure C.22: Mean absolute error of carelessness onset estimation with CODERS for different types of partially careless respondents in the simulated organizational survey.
+Note. The $x$-axis displays various prevalence levels. Results for different types of careless responding (rows) in different onset regimes (columns) are averaged across 100 repetitions.
+Appendix D Details on the Empirical Application
+D.1 Data Preprocessing
+[…]
+D.2 Additional Results
+[…]
+Figure D.1: Histogram of average reconstruction errors for the data of Johnson, (2005).
+[…]

### v4

@@
 Abstract
-Questionnaires in the behavioral and organizational sciences tend to be lengthy.
+Questionnaires in the behavioral sciences tend to be lengthy.
 However, literature suggests that survey length is a contributing factor to careless responding, with longer questionnaires yielding higher probability that participants start responding carelessly.
@@
 Introduction
-Research in the behavioral and organizational sciences often involves the administration of lengthy (online) self-report questionnaires.
-For instance, common personality measures tend to consist of hundreds of items, such as the Revised NEO Personality Inventory (Costa & McCrae,, 1992) with 240 items or the Minnesota Multiphasic Personality Inventory-2 Restructured Form (Ben-Porath & Tellegen,, 2008) with 338 items.
+Research in the behavioral sciences often involves the administration of lengthy (online) self-report questionnaires.
+For instance, common personality measures tend to consist of hundreds of items, such as the Revised NEO Personality Inventory (Costa and McCrae, 1992) with 240 items or the Minnesota Multiphasic Personality Inventory-2 Restructured Form (Ben-Porath and Tellegen, 2008) with 338 items.
 Even if one does not use such extensive measures, the number of items can easily reach three digits by including several shorter measures.
-However, it is well known that questionnaire length can have a concerning adverse effect on measurement accuracy: Respondents may experience fatigue or boredom as they progress through a lengthy questionnaire, which can provoke careless responding (e.g., Bowling et al.,, 2021; Gibson & Bowling,, 2020; Galesic & Bosnjak,, 2009), in particular in online administrations (e.g., Johnson,, 2005; Meade & Craig,, 2012).
+However, it is well known that questionnaire length can have a concerning adverse effect on measurement accuracy: Respondents may experience fatigue or boredom as they progress through a lengthy questionnaire, which can provoke careless responding (e.g., Bowling et al., 2021; Gibson and Bowling, 2020; Galesic and Bosnjak, 2009), in particular in online administrations (e.g., Johnson, 2005; Meade and Craig, 2012).
 We call such participants partially careless respondents as they initially provide accurate responses, but resort to careless responding after a certain item and remain careless for the remainder of the questionnaire.
-According to the definition of Ward & Meade, (2023), “careless responding occurs when participants are not basing their response on the item content, and it can occur when a respondent does not read an item, does not understand an item, or is unmotivated to think about what the item is asking”.11endnote: 1Perhaps more commonly used, Huang et al., (2012) define careless responding as “a response set in which the respondent answers a survey measure with low or little motivation to comply with survey instructions, correctly interpret item content, and provide accurate responses”.
-We prefer the definition of Ward & Meade, (2023), as it highlights that careless responding may not be intentional (cf.
-Huang et al.,, 2012; Ward & Pond,, 2015; Ward et al.,, 2017).
-Also note that Huang et al., (2012) refer to careless responding as insufficient effort responding in their definition.
-Other synonyms for careless responding are participant inattention (Maniaci & Rogge,, 2014), inconsistent responding (Greene,, 1978), protocol invalidity (Johnson,, 2005), and random responding (e.g., Beach,, 1989; Berry et al.,, 1992; Emons,, 2008).
-Notably, Schroeders et al., (2022) point out that the latter—random responding—may be a misnomer since careless responding can also arise in non-random patterns (e.g., a recurring sequence of 1-2-3-4-5).
-Careless responding has been identified as a major threat to the validity of research findings through a variety of psychometric and statistical issues such as lower scale reliability, errors in statistical hypothesis testing, and deteriorated model fit (e.g., Arias et al.,, 2020; Huang et al.,, 2012, 2015; Kam & Meyer,, 2015).
-It is widely prevalent (e.g., Bowling et al.,, 2016; Curran,, 2016; DeSimone et al.,, 2015; Meade & Craig,, 2012), suspected to be present in all questionnaire data (Ward & Meade,, 2023), and already incidence rates as low as 5% can be problematic (Arias et al.,, 2020; Credé,, 2010; Welz et al., 2024a, ; Welz et al., 2024b, ).
+According to the definition of Ward and Meade (2023), “careless responding occurs when participants are not basing their response on the item content, and it can occur when a respondent does not read an item, does not understand an item, or is unmotivated to think about what the item is asking”.
+[^1] Perhaps more commonly used, Huang et al.
+(2012) define careless responding as “a response set in which the respondent answers a survey measure with low or little motivation to comply with survey instructions, correctly interpret item content, and provide accurate responses”.
+We prefer the definition of Ward and Meade (2023), as it highlights that careless responding may not be intentional (cf.
+Huang et al., 2012; Ward and Pond, 2015; Ward et al., 2017).
+Also note that Huang et al.
+(2012) refer to careless responding as insufficient effort responding in their definition.
+Other synonyms for careless responding are participant inattention (Maniaci and Rogge, 2014), inconsistent responding (Greene, 1978), protocol invalidity (Johnson, 2005), and random responding (e.g., Beach, 1989; Berry et al., 1992; Emons, 2008).
+Notably, Schroeders et al.
+(2022) point out that the latter—random responding—may be a misnomer since careless responding can also arise in non-random patterns (e.g., a recurring sequence of 1-2-3-4-5).
+Careless responding has been identified as a major threat to the validity of research findings through a variety of psychometric and statistical issues such as errors in statistical hypothesis testing, deteriorated model fit, and lower scale reliability (e.g., Stosic et al., 2024; Voss, 2024; Arias et al., 2020; Huang et al., 2012, 2015).
+It is widely prevalent (e.g., Bowling et al., 2016; Curran, 2016; DeSimone et al., 2015; Meade and Craig, 2012), suspected to be present in all questionnaire data (Ward and Meade, 2023), and already incidence rates as low as 5% can be problematic (Arias et al., 2020; Credé, 2010; Welz et al., 2024, 2026).
 Often viewed as merely an insidious source of measurement error, careless responding can be a valuable source of information about study participants and the study itself.
-For instance, carelessness has been linked to participant personality (e.g., Bowling et al.,, 2016; DeSimone et al.,, 2020) and design factors such as study length (e.g., Bowling et al.,, 2021) and item wording.
-In particular negatively-worded items are prone to misresponses (e.g., Swain et al.,, 2008; Weijters & Baumgartner,, 2012) triggered by carelessness.
-Consequently, it is widely recommended to carefully screen survey data for the presence of careless responding (e.g., DeSimone et al.,, 2015; Huang et al.,, 2015).
-Numerous methods for identifying careless respondents after questionnaire administration (so-called screeners) have been proposed, for instance inconsistency indicators such as psychometric synonyms or antonyms (Meade & Craig,, 2012), invariability indicators such as the longstring index (Johnson,, 2005), or multivariate outlier analyses (e.g., Curran,, 2016).
-We refer to Arthur et al., (2021), DeSimone et al., (2015), and Ward & Meade, (2023) for detailed reviews.22endnote: 2A popular pre-administration method for the detection of careless responding is the inclusion of detection items (e.g., Table 1 in Arthur et al.,, 2021).
+For instance, carelessness has been linked to participant personality (e.g., Bowling et al., 2016; DeSimone et al., 2020) as well as design factors such as study length (e.g., Bowling et al., 2021) and item wording.
+In particular negatively-worded items are prone to misresponses (e.g., Swain et al., 2008; Weijters and Baumgartner, 2012) triggered by carelessness.
+Consequently, it is widely recommended to carefully screen survey data for the presence of careless responding (e.g., DeSimone et al., 2015; Huang et al., 2015).
+Numerous methods for identifying careless respondents after questionnaire administration (so-called screeners) have been proposed, for instance inconsistency indicators such as (resampled) personal reliabilty (Jackson, 1976; Curran, 2016; see also Goldammer et al., 2024) or psychometric synonyms/antonyms (Meade and Craig, 2012), invariability indicators such as the longstring index (Johnson, 2005), or multivariate outlier analyses (e.g., Curran, 2016).
+We refer to Arthur et al.
+(2021), DeSimone et al.
+(2015), and Ward and Meade (2023) for detailed reviews.
+[^2] A popular pre-administration method for the detection of careless responding is the inclusion of detection items (e.g., Table 1 in Arthur et al., 2021).
 Such detection items are based on the presumption that an attentive respondent will respond in a specific way, while careless respondent may fail to do so.
-For instance, it is expected that an attentive respondent would strongly disagree to so-called bogus items such as “I am paid biweekly by leprechauns” (Meade & Craig,, 2012).
-Alternative types of detection items are instructed items and self-report items (Arthur et al.,, 2021).
-More recently, Alfons & Welz, (2024), Ozaki, (2024), and Schroeders et al., (2022) have studied machine learning techniques for identifying careless respondents.
-Existing methods and their implementations in freely available software packages such as careless (Yentes & Wilhelm,, 2023) make screening for careless respondents a simple task for empirical researchers, thus elevating reliability and validity of research findings and playing an instrumental role in gaining insights into the phenomenon of careless responding.
+For instance, it is expected that an attentive respondent would strongly disagree to so-called bogus items such as “I am paid biweekly by leprechauns” (Meade and Craig, 2012).
+Alternative types of detection items are instructed items and self-report items (Arthur et al., 2021).
+More recently, Bloy et al.
+(2025), Alfons and Welz (2024), Ozaki (2024), and Schroeders et al.
+(2022) have studied machine learning techniques for identifying careless respondents.
+Existing methods and their implementations in freely available software packages such as careless (Yentes and Wilhelm, 2023) make screening for careless respondents a simple task for empirical researchers, thus elevating reliability and validity of research findings and playing an instrumental role in gaining insights into the phenomenon of careless responding.
 In addition, many popular screeners are intuitive and easy to understand.
 It should be noted that common indicators typically compute a single score for each participant, with respondents deemed careless and removed from subsequent analysis if their score exceeds a certain cutoff value.
-Although there exist many comparative studies on such screeners (e.g., Curran,, 2016; DeSimone & Harms,, 2018; Goldammer et al.,, 2020; Huang et al.,, 2012), their behavior under partially careless responding is unfortunately not well-understood.
+Although there exist many comparative studies on such screeners (e.g., Curran, 2016; DeSimone and Harms, 2018; Goldammer et al., 2020; Huang et al., 2012), their behavior under partially careless responding is unfortunately not well-understood.
 Furthermore, excluding partially careless respondents—particularly in lengthy surveys—may leave a lot of valuable information unused, namely the respective parts of the survey in which they responded attentively.
 We therefore aim to address a more nuanced question: Can we identify at which point in the survey participants start responding carelessly, if at all?
-Indeed, there is evidence that many—though not all—careless respondents tend to begin a survey as attentive and truthful respondents, but may resort to careless responding as the survey progresses, for instance due to fatigue or boredom (Bowling et al.,, 2021; Brower,, 2018; Galesic & Bosnjak,, 2009; Gibson & Bowling,, 2020; Ward et al.,, 2017); cf.
-also the literature on back random responding (e.g., Clark et al.,, 2003) or partial random responding (e.g.
-Berry et al.,, 1992; Gallen & Berry,, 1997).
+Indeed, there is evidence that many—though not all—careless respondents tend to begin a survey as attentive and truthful respondents, but may resort to careless responding as the survey progresses, for instance due to fatigue or boredom (Bowling et al., 2021; Brower, 2018; Galesic and Bosnjak, 2009; Gibson and Bowling, 2020; Ward et al., 2017); cf.
+also the literature on back random responding (e.g., Clark et al., 2003) or partial random responding (e.g.
+Berry et al., 1992; Gallen and Berry, 1997).
 Successful identification of the onset of careless responding can therefore yield valuable, complementary insights to those from existing screeners.
@@
 Rather than excluding partially careless respondents altogether, researchers can restrict their primary analyses to the parts of attentive responses, treating the careless responses as missing values.
-Common guidelines for handling missing data (e.g., Newman,, 2014) can thereby be followed.
+Common guidelines for handling missing data (e.g., Newman, 2014) can thereby be followed.
 In addition, accurately identifying the onset of careless responding may uncover implications for survey design (e.g., if many participants start responding carelessly at a similar point in the survey, it may be beneficial to rephrase or replace those items, or to place an intervention) and may lead to new theory and a deeper understanding of careless responding behavior (e.g., potentially linking certain personality traits to a tendency for an earlier or later onset of careless responding; cf.
-Gibbert et al.,, 2021, for theory-building based on outlying data).
+Gibbert et al., 2021, for theory-building based on outlying data).
 For the purpose of identifying at which point in the survey a given participant becomes careless, if at all, we introduce a novel method called careless onset detection in extensive rating-scale surveys (CODERS).
 Specifically, CODERS estimates the item at which carelessness onsets in each questionnaire participant, or an absence thereof.
-An important issue is that careless responding may manifest in multiple distinctive ways (e.g., Curran,, 2016; DeSimone et al.,, 2015; DeSimone & Harms,, 2018; Edwards,, 2019; Ward & Meade,, 2023).
+An important issue is that careless responding may manifest in multiple distinctive ways (e.g., Curran, 2016; DeSimone et al., 2015; DeSimone and Harms, 2018; Edwards, 2019; Ward and Meade, 2023).
 To capture various types of carelessness, CODERS can combine multiple dimensions of evidence for careless responding to construct a score that, for each item, measures if a given respondent has started responding carelessly by that item.
@@
 Moreover, CODERS builds on nonparametric estimation theory for cutoff values with statistical guarantees regarding the false positive rate (i.e., the probability of attentive respondents being incorrectly flagged as partially careless).
-To the best of our knowledge, CODERS is the first attempt to systematically detect the onset of careless responding.
-Through extensive simulation experiments, we find that CODERS accurately estimates carelessness onset (or a lack thereof) and achieves excellent sensitivity and specificity in detecting partially careless responding.
-Crucially, CODERS should not be viewed as a replacement of existing screeners for careless respondents, but as a complementary tool for a related but more nuanced task, as highlighted by an empirical application.
-Re-examining data from a a seminal study on careless responding in Big Five personality measurements (Johnson,, 2005), we demonstrate novel insights that can be obtained by estimating onset items of partially careless responding.
-Reflecting on those insights, we subsequently discuss limitations and potential extensions of CODERS.
-Our concluding discussion highlights how CODERS complements various existing approaches for the detection of careless respondents.
-Finally, we provide the freely available software package carelessonset (Welz & Alfons,, 2024), which implements CODERS in the statistical programming environment R (R Core Team,, 2024) and can be downloaded from https://github.com/mwelz/carelessonset.
-As such, it is a useful and accessible tool for any researcher who is concerned with survey fatigue or careless responding in general.
+To the best of our knowledge, only one other approach for systematically detecting the onset of careless responding has been proposed to date: The procedure of Yu and Cheng (2019) is based on item-response-theory (IRT) modeling and is more narroly focused on random responding.
+[…]
 Introducing CODERS
-The literature generally identifies three distinct ways in which careless responding may manifest: low internal consistency of the given responses (inconsistency), low variability of given responses (invariability), and impossibly fast responses (e.g., Curran,, 2016; DeSimone et al.,, 2015; DeSimone & Harms,, 2018; Edwards,, 2019; Ward & Meade,, 2023).
+The literature generally identifies three distinct ways in which careless responding may manifest: low internal consistency of the given responses (inconsistency), low variability of given responses (invariability), and impossibly fast responses (e.g., Curran, 2016; DeSimone et al., 2015; DeSimone and Harms, 2018; Edwards, 2019; Ward and Meade, 2023).
 We expect partially careless responding to result in a change in respondent behavior in at least one of these dimensions.
@@
   1. 1.
-    For each participant, a nonparametric changepoint detection test (Shao & Zhang,, 2010) is applied to a combined series of participant-item level measurements of careless responding.
+    For each participant, a nonparametric changepoint detection test (Shao and Zhang, 2010) is applied to a combined series of participant-item level measurements of careless responding.
   2. 2.
-    As a measurement of inconsistency, reconstruction errors (RE) obtained via an auto-associative neural network (autoencoder) (Kramer,, 1992) are used.
+    As a measurement of inconsistency, reconstruction errors (RE) obtained via an auto-associative neural network (autoencoder) (Kramer, 1992) are used.
 An autoencoder can be viewed as a nonlinear generalization of principal component analysis (PCA), and the (participant-item level) reconstruction errors can be interpreted as the difference between observed responses and what is expected under attentive responding.
@@
     As a measurement of invariability, we propose a novel algorithm called LongStringPattern (LSP).
-LSP can be viewed as a generalization of the famous longstring index of Johnson, (2005) to the participant-item level, also allowing for recurring patterns involving multiple answer categories.
+LSP can be viewed as a generalization of the famous longstring index of Johnson (2005) to the participant-item level, also allowing for recurring patterns involving multiple answer categories.
 Regarding the first dimension, inconsistency, the autoencoder learns response patterns that characterize attentive responding (cf.
-Alfons & Welz,, 2024).
+Alfons and Welz, 2024).
 We expect that random-like, content-independent responses cannot be learned well and are thus poorly reconstructed by the autoencoder, leading to a changepoint in reconstruction performance.
@@
 Conversely, if a respondent never becomes careless, we do not expect a changepoint in any dimension.
-Combining multiple indicators is a generally recommended practice to capture various types of careless responding (Goldammer et al.,, 2020; Huang et al.,, 2012; Meade & Craig,, 2012; Ward & Meade,, 2023).
+Combining multiple indicators is a generally recommended practice to capture various types of careless responding (Goldammer et al., 2020; Huang et al., 2012; Meade and Craig, 2012; Ward and Meade, 2023).
 We provide a detailed description of our assumptions in Appendix A and additional details on CODERS in Appendix B.
 Quantifying Inconsistency with Autoencoders
-Ward & Meade, (2023) define internal consistency of responses as patterns that are expected based on theoretical/logical grounds or trends in the data.
+Ward and Meade (2023) define internal consistency of responses as patterns that are expected based on theoretical/logical grounds or trends in the data.
 For instance, items that are part of the same construct are expected to correlate strongly in most participants, provided that participants are attentive.
-In contrast, inconsistent careless responding “generate[s] responses that fail to meet an expected level of consistency” (Ward & Meade,, 2023).
-Respondents may engage in inconsistent careless responding if they attempt to conceal their carelessness, for instance by randomly choosing from all response options (Ward & Meade,, 2023).
+In contrast, inconsistent careless responding “generate[s] responses that fail to meet an expected level of consistency” (Ward and Meade, 2023).
+Respondents may engage in inconsistent careless responding if they attempt to conceal their carelessness, for instance by randomly choosing from all response options (Ward and Meade, 2023).
 For our purposes, we consider the defining characteristic of inconsistent careless responding to be content-independent responses that are approximately randomly chosen from the response options, where the probability of being chosen may differ between response options, such as preferring extreme response options.
-In order to identify inconsistent careless responding, we propose to use the machine learning method of autoencoders (Kramer,, 1992), which were originally developed to filter random noise in signal processing applications.
-An autoencoder is a neural network (e.g., Bishop,, 2006; Goodfellow et al.,, 2016) that attempts to reconstruct its input.
+In order to identify inconsistent careless responding, we propose to use the machine learning method of autoencoders (Kramer, 1992), which were originally developed to filter random noise in signal processing applications.
+An autoencoder is a neural network (e.g., Bishop, 2006; Goodfellow et al., 2016) that attempts to reconstruct its input.
 In other words, the output variables are equal to the input variables.
@@
 To achieve the goal of learning the internal structures of a dataset, autoencoders first express the input data in terms of a low-dimensional representation before reconstructing the input data again from this low-dimensional representation (cf.
-Kramer,, 1992, and Chapter 14.6 in Goodfellow et al.,, 2016).
+Kramer, 1992, and Chapter 14.6 in Goodfellow et al., 2016).
 The use of such a low-dimensional representation makes autoencoders appealing for questionnaire data, as questionnaires typically contain multiple items to measure each construct of interest, such as attitudes, sentiments, or personality traits.
@@
 Figure 1: Schematic example of a neural network with an autoencoder structure.
-$\bm{x}_{1}$$\bm{x}_{2}$$\bm{x}_{3}$$\bm{x}_{4}$$\bm{\widehat{z}}_{1}$$\bm{\widehat{z}}_{2}$$\bm{\widehat{x}}_{1}$$\bm{\widehat{x}}_{2}$$\bm{\widehat{x}}_{3}$$\bm{\widehat{x}}_{4}$nput LayerMapping LayerBottleneck LayerDemapping LayerOutput Layer
+$\bm{x}_{1}$$\bm{x}_{2}$$\bm{x}_{3}$$\bm{x}_{4}$$\bm{\widehat{z}}_{1}$$\bm{\widehat{z}}_{2}$$\bm{\widehat{x}}_{1}$$\bm{\widehat{x}}_{2}$$\bm{\widehat{x}}_{3}$$\bm{\widehat{x}}_{4}$Input LayerMapping LayerBottleneck LayerDemapping LayerOutput Layer
 Note.
@@
 The network architecture of an autoencoder is symmetric around this bottleneck layer such that it can be viewed as having an encoder-decoder architecture: the encoder compresses the data into the low-dimensional representations, while the decoder reconstructs the input data again.
-It follows that an autoencoder can be seen as a dimension reduction technique, specifically as a nonlinear generalization of principal component analysis (PCA; see Kramer,, 1991).
+It follows that an autoencoder can be seen as a dimension reduction technique, specifically as a nonlinear generalization of principal component analysis (PCA; see Kramer, 1991).
 Hence, the number of nodes in the bottleneck layer is an important design choice in the architecture of an autoencoder.
 In the context of questionnaire data, we recommend setting this number equal to the number of constructs measured in the questionnaire.
-For example, consider data from the Revised NEO Personality Inventory (NEO-PI-R; Costa & McCrae,, 1992), which contains 240 items and measures six so-called facets of each of the Big Five personality traits.
+For example, consider data from the Revised NEO Personality Inventory (NEO-PI-R; Costa and McCrae, 1992), which contains 240 items and measures six so-called facets of each of the Big Five personality traits.
 Thus, this instrument measures a total of 30 facets, such as anxiety and modesty, so that we would set the number of nodes in the bottleneck layer to 30.
-Further intuition and a thorough mathematical definition of autoencoders can be found in Appendix B.2.
-Although there are other design choices besides the number of nodes in the bottleneck layer (see Appendix B.2.3), our software implementation in package carelessonset (Welz & Alfons,, 2024) provides default values that we recommend researchers to use.
+Further intuition and a thorough mathematical definition of autoencoders can be found in Appendix B.
+Although there are other design choices besides the number of nodes in the bottleneck layer, our software implementation in package carelessonset (Welz and Alfons, 2026) provides default values that we recommend researchers to use.
 The only input required by the user is the number of constructs measured in the survey (for the number of nodes in the bottleneck layer).
@@
 Hence, if participant $i$ starts providing inconsistent careless responses at item $k\in\{1,\dots,p\}$, we expect a changepoint at position $k$ in the participant’s series of $p$ reconstruction errors, as reconstruction errors are expected to be higher from item $k$ onward.
-An example can be found in Figure 2, where we show simulated responses of a hypothetical participant in a 300-item questionnaire (5 answer categories per item) who starts choosing answer categories completely at random from the 172nd item, resulting in systemically larger autoencoder reconstruction errors from that item onward.
+An example can be found in Figure 2, where we show simulated responses of a hypothetical participant in a 300-item questionnaire (5 answer categories per item) who starts choosing answer categories completely at random from the 183rd item, resulting in systemically larger autoencoder reconstruction errors from that item onward.
 Figure 2: Simulated illustrative example of autoencoder reconstruction errors of an inconsistent partially careless respondent to a 300-item questionnaire comprising five-point Likert scales.
-Note. The simulated participant started choosing answer categories completely at random from the 172nd item onward.
+Note. The simulated participant started choosing answer categories completely at random from the 183rd item onward.
 Quantifying Invariability
-We propose to measure response invariability through a novel algorithm inspired by the longstring index of Johnson, (2005) that we call LongStringPattern (LSP).
+We propose to measure response invariability through a novel algorithm inspired by the longstring index of Johnson (2005) that we call LongStringPattern (LSP).
 This algorithm exploits that invariable careless responses are characterized by content-independent response patterns.
 A pattern sequence has the defining property that it consists of repeated occurrences of the same response pattern.
-Consider the sequence of responses $-2-1-2-\ldots-1-$ 2 - … - 1 - 2, which consists of repeated occurrences of the pattern $1-2$.
+Consider the sequence of responses $1-2-1-2-\ldots-1-2$, which consists of repeated occurrences of the pattern 121-2.
 Denoting the length of a pattern by $l$, the above example exhibits a pattern of length $l=2$.
-Analogously, the sequence $-3-1-2-3-\ldots-1-2-$ 3 - … - 1 - 2 - 3 consists of repeated occurrences of the pattern $1-2-3$ of length $l=3$, whereas the straightlining sequence $1-1-\ldots-$ 1 - … - 1 consists of repeated occurrences of 1, which is a pattern of length $l=1$.
+Analogously, the sequence $1-2-3-1-2-3-\ldots-1-2-3$ consists of repeated occurrences of the pattern 1231-2-3 of length $l=3$, whereas the straightlining sequence $1-1-\ldots-1$ consists of repeated occurrences of 1, which is a pattern of length $l=1$.
 Define $l\textsf{-pattern}_{ij}$ as the number of consecutive items in the longest sequence of a recurring pattern of length $l$ in which the response of participant $i$ to item $j$ takes part, or assign the value 1 if the response is not part of a recurring pattern of length $l$.
 Consider the following two illustrative examples.
-First, for $l=2$, the response sequence $1-2-1-2-1-4-3-5-4-5-4$ will be assigned the 2-pattern sequence $4-4-4-4-1-1-1-4-4-4-4$ because the first four responses and last four responses comprise two occurrences of the distinct patterns $1-2$ and $5-4$, respectively, which are both of length $l=2$.
-The central three responses ($1-4-3$) are not part of any recurring pattern of length $l=2$, hence they are each assigned the value 1.
-Second, for a recurring pattern of length $l=1$ (i.e., consecutive identical responses), the response sequence $3-2-3-3-1-4-1-1-1$ is assigned the 1-pattern sequence $1-1-2-2-1-1-3-3-3$, since the subsequences of consecutive identical responses $3-3$ and $1-1-$ 1 - 1 comprise two and three responses, respectively.33endnote: 3The 1-pattern sequence can be seen as a generalization of the the longstring index of Johnson, (2005), which we recover by calculating $\max_{1\leq j\leq p}1\textsf{-pattern}_{ij}$ for each respondent $i=1,\dots,n$.
+First, for $l=2$, the response sequence 121214354541-2-1-2-1-4-3-5-4-5-4 will be assigned the 2-pattern sequence 444411144444-4-4-4-1-1-1-4-4-4-4 because the first four responses and last four responses comprise two occurrences of the distinct patterns 121-2 and 545-4, respectively, which are both of length $l=2$.
+The central three responses (1431-4-3) are not part of any recurring pattern of length $l=2$, hence they are each assigned the value 1.
+Second, for a recurring pattern of length $l=1$ (i.e., consecutive identical responses), the response sequence 3233141113-2-3-3-1-4-1-1-1 is assigned the 1-pattern sequence 1122113331-1-2-2-1-1-3-3-3, since the subsequences of consecutive identical responses 333-3 and 1111-1-1 comprise two and three responses, respectively.
+[^3] The 1-pattern sequence can be seen as a generalization of the the longstring index of Johnson (2005), which we recover by calculating $\max_{1\leq j\leq p}1\textsf{-pattern}_{ij}$ for each respondent $i=1,\dots,n$.
 In short, higher values in the $l$-pattern sequence capture a tendency to respond in fixed patterns of length $l$.
@@
 In this paper, we vary $l$ from 1 to the number of answer categories in the survey.
-The rationale behind this choice is that a careless respondent may repeat, e.g., the pattern $1-2-3-4-5$ in case of five answer categories.
+The rationale behind this choice is that a careless respondent may repeat, e.g., the pattern 123451-2-3-4-5 in case of five answer categories.
 At the same time, we consider it unlikely that careless respondents choose complicated individual patterns whose length exceeds the number of answer categories.
-We provide a detailed description of the LSP algorithm in Appendix B.3.
-An example of an LSP sequence is shown in Figure 3, which shows simulated responses of a hypothetical participant in a 300-item questionnaire (5 answer categories per item) who starts responding carelessly using the recurring pattern $3-4$ from the 260th item, resulting in systematically larger LSP values from that item onward.
+We provide a detailed description of the LSP algorithm in Appendix B.
+An example of an LSP sequence is shown in Figure 3, which shows simulated responses of a hypothetical participant in a 300-item questionnaire (5 answer categories per item) who starts responding carelessly using the recurring pattern 131-3 from the 245th item, resulting in systematically larger LSP values from that item onward.
 Figure 3: Simulated illustrative example of a LongStringPattern sequence of an invariable partially careless respondent to a 300-item questionnaire comprising five-point Likert scales.
-Note. The simulated participant started responding carelessly with the recurring pattern $3-4$ from the 260th item onward.
+Note. The simulated participant started responding carelessly with the recurring pattern 131-3 from the 245th item onward.
 Identifying Carelessness Onset via Changepoint Detection
@@
 Figures 2 and 3), and no changepoint in the absence of carelessness.
-We therefore propose to apply the nonparametric cumulative sum self-normalization test of Shao & Zhang, (2010), which was developed in the context of time series analysis.
-It searches for the location of a possible changepoint in a multivariate series and has attractive theoretical guarantees (see Shao & Zhang,, 2010; Zhao et al.,, 2022, and our detailed description in Appendix B.4).
+We therefore propose to apply the nonparametric cumulative sum self-normalization test of Shao and Zhang (2010), which was developed in the context of time series analysis.
+It searches for the location of a possible changepoint in a multivariate series and has attractive theoretical guarantees (see Shao and Zhang, 2010; Zhao et al., 2022; and our detailed description in Appendix B).
 For each point in a multivariate series (we use measurements of carelessness associated with each questionnaire item), this changepoint detection test calculates a statistic that compares a quantity of interest (we take the mean) in the periods before and after the current point (item).
@@
 This gives CODERS great flexibility, as it allows us to use any combination of measurements of carelessness without making possibly unrealistic assumptions on their distributions or obtaining cutoff values in an ad-hoc manner.
-In line with literature that suggests that there should be overwhelming evidence when labeling respondents as careless (e.g., Huang et al.,, 2012), we recommend using a conservative significance level no greater than 1% for identifying changepoints.
-As an illustrative example, consider Figure 4(a).
-It illustrates two dimensions computed from a given participant’s responses, namely autoencoder reconstruction errors and the LSP sequence.
-The former series has been shown before in Figure 2.
-The bottom panel of Figure 4(a) shows the test statistics computed from this two-dimensional series, where the maximum value at the 172nd item clearly exceeds the critical value.
-That is, CODERS identifies inconsistent carelessness from the 172nd item onward, as the reconstruction errors are in general much larger than before.
-Indeed, this simulated participant started choosing responses at random at the identified item.
-Figure 4(b) is similar, but computed from a different respondent, whose LSP sequence has been shown before in Figure 3.
-CODERS flags a changepoint at the 261st item.
-From this item onward, the LSP values are clearly elevated, indicating invariable carelessness.
-In fact, this simulated participant started repeating the pattern $3-4$ just one item before the estimated onset.
+In line with literature that suggests that there should be overwhelming evidence when labeling respondents as careless (e.g., Huang et al., 2012), we recommend using a conservative significance level no greater than 1% for identifying changepoints.
 Figure 4: Illustrative example output from CODERS using responses of two simulated partially careless respondents to a 300-item questionnaire.
@@
 Note.
-The left subfigure shows the output from CODERS for a simulated partially inconsistent respondent who starts responding randomly from the 172nd item onward, while the right subfigure does so for a simulated partially invariable respondent who starts repeating the pattern $3-4$ from the 260th item onward.
+The left subfigure shows the output from CODERS for a simulated partially inconsistent respondent who starts responding randomly from the 183rd item onward, while the right subfigure does so for a simulated partially invariable respondent who starts repeating the pattern 131-3 from the 245th item onward.
 Each subfigure visualizes the autoencoder reconstruction errors (RE), LongStringPattern (LSP) sequence, and test statistics associated with the corresponding participant’s responses.
-The blue solid vertical lines indicate the carelessness onset items identified by CODERS at significance level 0.1%, which correspond to the true onset item 172 for the partially inconsistent respondent, and to item 261 (one item after true onset) for the partially invariable respondent.
-The dashed horizontal line is the critical value of the test statistic at the 0.1% significance level (192.5; see Table B.2 in Appendix B.4).
+The blue solid vertical lines indicate the carelessness onset items identified by CODERS at significance level 0.1%, which correspond to item 181 (two items before true onset) for the partially inconsistent respondent, and to the true onset item 245 for the partially invariable respondent.
+The dashed horizontal line is the critical value of the test statistic at the 0.1% significance level (192.5; see Table B2 in Appendix B).
 The reconstruction errors in the left subfigure are those shown in Figure 2, and the LSP sequence in the right subfigure is that shown in Figure 3.
+As an illustrative example, consider Figure 4(a).
+It illustrates two dimensions computed from a given participant’s responses, namely autoencoder reconstruction errors and the LSP sequence.
+The former series has been shown before in Figure 2.
+The bottom panel of Figure 4(a) shows the test statistics computed from this two-dimensional series, where the maximum value at the 181st item clearly exceeds the critical value.
+That is, CODERS identifies inconsistent carelessness from the 181st item onward, as the reconstruction errors are in general much larger than before.
+Indeed, this simulated participant started choosing responses at random just two items after the estimated onset.
+Figure 4(b) is similar, but computed from a different respondent, whose LSP sequence has been shown before in Figure 3.
+CODERS flags a changepoint at the 245th item.
+From this item onward, the LSP values are clearly elevated, indicating invariable carelessness.
+In fact, this simulated participant started repeating the pattern 131-3 at the identified item.
 Simulation Experiments
-After introducing CODERS, we now investigate its behavior via simulation experiments modeled after data collected by Johnson, (2005).
+After introducing CODERS, we now investigate its behavior via simulation experiments modeled after data collected by Johnson (2005).
 We use the data from this seminal study on careless responding in our empirical application in a later section, where the data are described in more detail.
+Unfortunately, it is not possible to compare CODERS with the approach of Yu and Cheng (2019), which is the only other method for detecting the onset of careless responding that we are aware of, as no R code seems to be available.
 Data Generation
-The survey instrument used by Johnson, (2005) is a 300-item representation of the NEO-PI-R (Goldberg,, 1999) measurement of the Big Five personality traits (openness, conscientiousness, extraversion, agreeableness, and neuroticism; Goldberg,, 1992) by means of five-point Likert scales.
+The survey instrument used by Johnson (2005) is a 300-item representation of the NEO-PI-R (Goldberg, 1999) measurement of the Big Five personality traits (openness, conscientiousness, extraversion, agreeableness, and neuroticism; Goldberg, 1992) by means of five-point Likert scales.
 Each of the five traits comprises six subdomains (facets) so that a total of $5\times 6=30$ facets are measured.
-We base our simulation design on the between-item correlation structure and per-item response distributions in the data collected by Johnson, (2005), calculated after removing any presumably careless respondents detected in our empirical application.
+We base our simulation design on the between-item correlation structure and per-item response distributions in the data collected by Johnson (2005), calculated after removing any presumably careless respondents detected in our empirical application.
 Figure 5 visualizes the resulting correlation matrix and marginal response distributions.
@@
 In the correlation matrix, one can identify the five traits and, within each block of within-trait correlations, also the six corresponding facets.
-Using these correlations and marginal response distributions, we generate rating-scale responses of $n=500$ participants to the $p=300$ items,44endnote: 4We use package simstudy (Goldfeld & Wujciak-Jens,, 2020) to generate the data, for which the target values of correlations only hold approximately on the population level.
-with the items in the same randomized order as in Johnson, (2005).
+Using these correlations and marginal response distributions, we generate rating-scale responses of $n=500$ participants to the $p=300$ items, [^4] We use package simstudy (Goldfeld and Wujciak-Jens, 2020) to generate the data, for which the target values of correlations only hold approximately on the population level.
+with the items in the same randomized order as in Johnson (2005).
 Figure 5: Correlation matrix and marginal response distributions used in simulations.
@@
 Note.
-The correlation matrix and the marginal response distributions are calculated from the data of Johnson, (2005) after removing presumably careless respondents.
+The correlation matrix and the marginal response distributions are calculated from the data of Johnson (2005) after removing presumably careless respondents.
 Of the $p=300$ items, 152 are positively scored and 148 are negatively scored.
@@
 We consider three regimes for the carelessness onset: (i) the baseline setting where the onset item is randomly sampled from $\{30,31,\dots,269,270\}$ so that carelessness starts between 10% and 90% of all items, (ii) early onset randomly sampled from items $\{30,31,\dots,149,150\}$ (between 10% and 50% of all items), and (iii) late onset randomly sampled from items $\{150,151,\dots,269,270\}$ (between 50% and 90% of all items).
-[…]
+From the carelessness onset item onward, we generate the responses from one of the following four careless response types: (i) random responding (choosing answer categories completely at random), (ii) extreme responding (randomly choosing between the two most extreme answer categories), (iii) straightlining (repeating the same, randomly determined answer category), or (iv) pattern responding (repeating a fixed pattern such as 1231-2-3 or 545-4, which is randomly determined from all possible patterns of length two to five).
+Specifically, we assign these careless response types to the partially careless respondents so that there are $\lfloor\gamma n/4\rfloor$ respondents for each response type, where $\lfloor\ .\ \rfloor$ denotes the operator for rounding down to the nearest integer.
 We repeat the above described data generating process 100 times.
@@
 Overall, CODERS rarely flags changepoints in fully attentive respondents.
-Even with our most liberal choice of significance level $\alpha$ (1%), overrejection increases only slowly to about 3% for 60% prevalence, before it reaches between 3.5% and 5.7% at 80% prevalence (depending on the onset regime).
+Even with our most liberal choice of significance level $\alpha$ (1%), overrejection increases only slowly to about 3% for 60% prevalence, before it reaches between 3.0% and 5.6% at 80% prevalence (depending on the onset regime).
 Other than for extremely high prevalence levels, the FPR stays reasonably close to the respective nominal level, indicating that CODERS controls Type I errors fairly well.
@@
 In addition, Figure 7 (bottom row) displays the MAE of the estimated onset in partially careless respondents.
-The prevalence level and onset regime have little effect on the accuracy of the estimates, with an MAE hovering around 2 to 3 depending on the significance level.
+The prevalence level and onset regime have little effect on the accuracy of the estimates, with an MAE hovering around 2 to 3, depending on the significance level.
 That is, on average the estimated onset is only two to three items away from the true onset, indicating that CODERS achieves excellent accuracy in estimating carelessness onset.
@@
 In other words, this is an issue of statistical power, which explains why it can be mitigated to some extent by choosing a less conservative significance level.
-For additional investigation, we apply traditional methods for identifying careless respondents also used by Johnson, (2005).
-They do not improve upon CODERS (see Appendix C.1), therefore further highlighting the difficulty of detecting random respondents in this simulation design.
+For additional investigation, we apply traditional methods for identifying careless respondents also used by Johnson (2005).
+They do not improve upon CODERS (see Appendix C), therefore further highlighting the difficulty of detecting random respondents in this simulation design.
 Coming back to CODERS, it shows similar but far less pronounced behavior for extreme responding (see the second row in Figure 8).
@@
 On the other hand, it struggles with random responding in this simulation design, where decent power is only achieved in case of late carelessness onset and with a not-too-conservative significance level.
-However, results for traditional methods (see Appendix C.1) suggest that distinguishing random responding from attentive responding may be an intrinsically difficult problem for the chosen correlation structure and marginal distributions.
+However, results for traditional methods (see Appendix C) suggest that distinguishing random responding from attentive responding may be an intrinsically difficult problem for the chosen correlation structure and marginal distributions.
 Overall, we find that (i) CODERS reliably discriminates between attentive and partially careless respondents, as indicated by near-zero false positive rates and low overall false negative rates, (ii) CODERS accurately estimates the location of carelessness onset, and (iii) the performance of CODERS is stable across a wide range of prevalence levels, demonstrating robustness.
@@
 In summary, other than for random respondents, overall power remains adequate when using a not-too-conservative significance level (e.g., 0.5% or 1%).
-[…]
+Second, we vary the number of participants by setting $n=248$ and $n=\text{1,000}$.
+[^5] Since we generate four types of partially careless respondents in the simulations, the sample size should be divisible by 4, explaining the unusual choice of $n=248$ (rather than $n=250$).
+The overall FNR in partially careless respondents slightly increases for the smaller sample size and slightly decreases for the larger sample size, with bigger changes being observed for random respondents.
+That is, the effect of the sample size on power is relatively small except for random responding, where further improvements in performance may be possible for even larger sample sizes.
 Third, we consider an additional type of careless responding to investigate the performance of CODERS when the lines between attentive and careless responding are blurry.
@@
 This highlights that while CODERS is effective in detecting a single changepoint, it is simply not designed for detecting multiple changepoints.
-Finally, having so far considered simulations based on personality assessments, we include a simulation inspired by a study from organizational research.
-This simulated survey is shorter, with $p=198$ items measuring a variety of organizationally relevant constructs, and we apply a different strategy for randomization better suited in organizational contexts.
-Furthermore, there exist complex relationships between (some of) the measured constructs.
-The FPR of CODERS is somewhat more elevated for high prevalence levels.
-Interestingly, the FNR is slightly lower for random responding but slightly higher for extreme responding, resulting in a comparable FNR overall.
-All in all, the results for CODERS remain relatively similar to those in the simulations based on personality assessments, indicating that its performance may not depend much on the institutional setting of the survey.
 Empirical Application
-After studying the performance of CODERS in simulations inspired by the data from Johnson, (2005), we now re-examine the responses from this seminal study for partial carelessness.
+After studying the performance of CODERS in simulations inspired by the data from Johnson (2005), we now re-examine the responses from this seminal study for partial carelessness.
 Data
-The data collected by Johnson, (2005) concern the five factor model of personality (Goldberg,, 1992), which assumes that variation in a measurement of personality can be explained by the so-called Big Five personality traits (factors): openness, conscientiousness, extraversion, agreeableness, and neuroticism.
+The data collected by Johnson (2005) concern the five factor model of personality (Goldberg, 1992), which assumes that variation in a measurement of personality can be explained by the so-called Big Five personality traits (factors): openness, conscientiousness, extraversion, agreeableness, and neuroticism.
 Each of the Big Five factors can be further decomposed into six subdomains (facets): for instance, depression and anxiety are facets of the neuroticism trait.
-Consequently, common Big Five instruments, such as the 240-item NEO-PI-R (Costa & McCrae,, 1992), typically measure a total of $5\times 6=30$ facets.
-The data set of Johnson, (2005) consists of internet-collected responses to a 300-item representation of the NEO-PI-R inventory (Goldberg,, 1999), in which each of the 30 facets is measured by 10 items on a five-point Likert scale.
-[…]
+Consequently, common Big Five instruments, such as the 240-item NEO-PI-R (Costa and McCrae, 1992), typically measure a total of $5\times 6=30$ facets.
+The data set of Johnson (2005) consists of internet-collected responses to a 300-item representation of the NEO-PI-R inventory (Goldberg, 1999), in which each of the 30 facets is measured by 10 items on a five-point Likert scale.
+Johnson (2005) identifies straightliners via the longstring index (which the author proposed in that paper) and inconsistent careless respondents via personal reliability (Jackson, 1976) and psychometric antonym (e.g., Meade and Craig, 2012) scores.
+The data from this study are made publicly available by the author at https://osf.io/sxeq5/.
+Importantly, from the original sample of 23,994 participants, the author removed 918 participants that were likely duplicates, 1,455 participants with too many straightlining responses, and 688 participants with too many missing responses, leaving $n=\text{20,933}$ participants in the provided data set.
+Unfortunately, the raw data is not available at the above link.
+The minimal preprocessing that we applied to the provided data set is described in Appendix D.
 In general, it seems likely that survey fatigue is present in a sample of responses to an online questionnaire with a lengthy battery of 300 items.
-For instance, Bowling et al., (2021) estimate that if one wishes that at least 90% of participants do not respond carelessly to more than 20% of all items, one should not include more than 176 items in an online questionnaire.
-By applying CODERS—which is specifically designed to identify partially careless responding through a change in response behavior—in addition to the traditional screeners applied by Johnson, (2005), new and complimentary insights into the incidence and nature of (partially) careless responding may emerge.
+For instance, Bowling et al.
+(2021) estimate that if one wishes that at least 90% of participants do not respond carelessly to more than 20% of all items, one should not include more than 176 items in an online questionnaire.
+By applying CODERS—which is specifically designed to identify partially careless responding through a change in response behavior—in addition to the traditional screeners applied by Johnson (2005), new and complimentary insights into the incidence and nature of (partially) careless responding may emerge.
 Results
-[…]
+Throughout the discussion of the results, it is important to keep in mind that Johnson (2005) already removed straightliners from the publicly available data set.
+To maintain consistency in reporting, any percentages of participants refer to a baseline sample size of $N=\text{22,448}$ (i.e., including straightliners).
+For instance, the 1,455 straightliners found by Johnson (2005) amount to 6.5% of participants.
 General Overview
@@
 On average, the location of the carelessness onset item for flagged participants is estimated around the 145th item, meaning that carelessness onsets on average after nearly half of the 300 items.
-These findings are broadly in line with the results of Bowling et al., (2021, Figure 1), whose model suggests a low probability that a representative participant starts responding carelessly at the 145th item.
 The earliest onset occurs at item 15, whereas the latest onset is at item 294.
 Figure 10 provides a more detailed visual overview of the detected onset items for the flagged participants at different significance levels.
-Table 1: Results of CODERS for the data of Johnson, (2005).
+Table 1: Results of CODERS for the data of Johnson (2005).
 Detected onset
@@
 $\alpha=0.01\phantom{0}$ 856 (3.8%) 15 147.9 294
-[…]
-Figure 10: Estimated onset of careless responding for participants flagged by CODERS in the data of Johnson, (2005).
+Note.
+The methods are applied to the responses of $n=\text{20,993}$ participants to $p=300$ items of a Big Five personality measurement.
+Among others, participants with too many straightlining responses have already been removed from the data provided by Johnson.
+For consistent reporting, percentages refer to a baseline sample size of $N=\text{22,448}$, which includes straightliners.
+Figure 10: Estimated onset of careless responding for participants flagged by CODERS in the data of Johnson (2005).
 Note.
@@
 The dashed horizontal line represents the average location of carelessness onset across the flagged participants.
-We now turn to analyzing the agreement between CODERS and the two traditional methods applied by Johnson, (2005), personal reliability and psychometric antonym.
-Following Johnson,, we reverse psychometric antonym scores so that low scores are associated with carelessness in both methods, and we flag respondents as careless whose personal reliability and psychometric antonym scores are lower than $0.3$ and $-0.03$, respectively.
-The results are shown in Table 2.
-Personal reliability flags 140 participants (0.6%) as careless, while psychometric antonym flags 284 respondents (1.3%).66endnote: 6We use the R package careless (Yentes & Wilhelm,, 2023) to compute personal reliability and psychometric antonym scores.
-The use of different software implementations (with potentially different default values) may explain differences with the numbers reported by Johnson, (2005).
-CODERS, personal reliability, and psychometric antonyms generally show little agreement regarding the detected careless respondents.
-Personal reliability and psychometric antonym scores correlate by a low value of 0.314, yielding only 11 participants that are flagged by both methods.
-Depending on the significance level $\alpha$, CODERS and personal reliability have between 2 and 10 flagged participants in common, while the number of jointly flagged respondents for CODERS and psychometric antonym ranges between 9 and 24.
-All three methods share a single flagged participant.
-Table 2: Complementary findings of CODERS and traditional methods for the data of Johnson, (2005).
+Table 2: Complementary findings of CODERS and traditional methods for the data of Johnson (2005).
 Also flagged by CODERS
@@
 Both 11 1 11 21
-[…]
+Note.
+The methods are applied to the responses of $n=\text{20,993}$ participants to $p=300$ items of a Big Five personality measurement.
+Among others, participants with too many straightlining responses have already been removed from the data provided by Johnson.
+For consistent reporting, percentages refer to a baseline sample size of $N=\text{22,448}$, which includes straightliners.
+Too small percentages are omitted for better readability.
+We now turn to analyzing the agreement between CODERS and the two traditional methods applied by Johnson (2005), personal reliability and psychometric antonym.
+Following Johnson, we reverse psychometric antonym scores so that low scores are associated with carelessness in both methods, and we flag respondents as careless whose personal reliability and psychometric antonym scores are lower than $0.3$ and 0.03-0.03, respectively.
+The results are shown in Table 2.
+Personal reliability flags 140 participants (0.6%) as careless, while psychometric antonym flags 284 respondents (1.3%).
+[^6] We use the R package careless (Yentes and Wilhelm, 2023) to compute personal reliability and psychometric antonym scores.
+The use of different software implementations (with potentially different default values) may explain differences with the numbers reported by Johnson (2005).
+CODERS, personal reliability, and psychometric antonyms generally show little agreement regarding the detected careless respondents.
+Personal reliability and psychometric antonym scores correlate by a low value of 0.314, yielding only 11 participants that are flagged by both methods.
+Depending on the significance level $\alpha$, CODERS and personal reliability have between 2 and 10 flagged participants in common, while the number of jointly flagged respondents for CODERS and psychometric antonym ranges between 9 and 24.
+All three methods share a single flagged participant.
 It is not surprising that there is little consensus of CODERS with the two traditional methods, as they are designed for different tasks.
@@
 While these examples are naturally selective, similar plots for all participants flagged by CODERS or the two traditional methods can be found in our online repository at https://github.com/mwelz/carelessonset_plots.
-Figure 11: Scores and changepoint item detected by CODERS in selected participants from the data of Johnson, (2005).
+Figure 11: Scores and changepoint item detected by CODERS in selected participants from the data of Johnson (2005).
 (a) Presumably inconsistent. Onset estimated by CODERS at item 148.
@@
 Summary statistics for the corresponding participants are provided in Table 3.
-Table 3: Average scores before and after the changepoint item detected by CODERS in selected participants from the data of Johnson, (2005).
+Table 3: Average scores before and after the changepoint item detected by CODERS in selected participants from the data of Johnson (2005).
 Before After
@@
 These two cases are representative for the majority of respondents flagged by CODERS, likely starting attentively but starting to respond carelessly at some point.
-In addition, more respondents flagged by CODERS seem to exhibit inconsistency than invariability, as Johnson, already removed straightliners from the publicly available data set.
+In addition, more respondents flagged by CODERS seem to exhibit inconsistency than invariability, as Johnson already removed straightliners from the publicly available data set.
 However, CODERS also identifies participants with remarkably different behavior.
@@
 Although there is a clear change in behavior, as found by CODERS, this participant seems to respond invariably throughout the survey.
-[…]
+Finally, the example in Figure 11(f) and Table 3(f) corresponds to a participant flagged by psychometric antonym but not CODERS.
+A visual inspection confirms that there is indeed no clear changepoint in either the reconstruction errors or the LSP sequence.
+However, the reconstruction errors are similar in magnitude and variability to those found in inconsistent response periods from other examples.
+To investigate this further, we computed average reconstruction errors across all $p=300$ items for the $n=\text{20,993}$ participants.
+Only 6 participants have a higher average reconstruction error (cf.
+Figure D1 in Appendix D), providing overwhelming evidence that this participant may be an inconsistent careless respondent throughout the whole questionnaire.
 In summary, we reiterate that the first two examples are representative for the majority of participants flagged by CODERS, in line with the assumption that respondents start attentive and possibly become careless.
@@
 In this sense, CODERS—which is intended for detecting partial carelessness through a change in behavior—is complementary to those existing methods.
-If one suspects that respondents who have been careless throughout all items are present, we recommend to apply established detection methods following the guidelines of Arthur et al., (2021) and Ward & Meade, (2023).
+If one suspects that respondents who have been careless throughout all items are present, we recommend to apply established detection methods following the guidelines of Arthur et al.
+(2021) and Ward and Meade (2023).
 Respondents that are flagged as careless by such methods, but for whom no changepoint was flagged by CODERS, have likely been careless throughout the survey.
@@
 For such cases, our empirical application demonstrates that CODERS—combined with a post-hoc analysis of the identified segments of responses—remains a valuable tool for gaining detailed insights into careless response behaviors.
-Third, there may be participants who alternate between periods of attentive and careless responding (Berry et al.,, 1992; Clark et al.,, 2003; Meade & Craig,, 2012).
-Our simulations indicate that CODERS remains conservative in this case and rarely flags one of the changepoints (see Appendix C.5).
+Third, there may be participants who alternate between periods of attentive and careless responding (Berry et al., 1992; Clark et al., 2003; Meade and Craig, 2012).
+Our simulations indicate that CODERS remains conservative in this case and rarely flags one of the changepoints (see Appendix C).
 CODERS should therefore be viewed as a promising first step towards detecting periods of careless responding, with an extension for the detection of multiple changepoints being an area of further research.
-Finally, the changepoint detection procedure of Shao & Zhang, (2010) employed by CODERS is based on asymptotic arguments when the number of items is large.
+Finally, the changepoint detection procedure of Shao and Zhang (2010) employed by CODERS is based on asymptotic arguments when the number of items is large.
 Hence, the number of items in the questionnaire is the main determinant for the statistical power of CODERS.
 For the detection of inconsistent careless responding, the number of participants plays a secondary role regarding power: the more attentive responses the autoencoder can learn from, the better the reconstruction errors can distinguish between attentive and careless responding (cf.
-Appendix C.3).
+Appendix C).
 The LSP sequence of a given participant, on the other hand, depends only on their own responses, meaning that the number of observations should not play a role concerning power in detecting invariable careless responding.
-CODERS is thus primarily designed for relatively lengthy questionnaires, and our simulations suggest that there is a gradual drop in statistical power for detecting inconsistency as survey length decreases (see Appendix C.2).
+CODERS is thus primarily designed for relatively lengthy questionnaires, and our simulations suggest that there is a gradual drop in statistical power for detecting inconsistency as survey length decreases (see Appendix C).
 However, partial carelessness due to survey fatigue may be less prevalent in short questionnaires (cf.
-Bowling et al.,, 2021).
+Bowling et al., 2021).
 A detailed study of partially careless responding in shorter surveys is therefore beyond the scope of this paper.
 Response Times Revisited
-In addition to inconsistency and invariability, fast responding is considered the third major manifestation of carelessness (e.g., Ward & Meade,, 2023).
+In addition to inconsistency and invariability, fast responding is considered the third major manifestation of carelessness (e.g., Ward and Meade, 2023).
 In (online) surveys, response time is typically measured by the total time a participant spent on the questionnaire, time spent on each questionnaire page, or (less common) time spent on each questionnaire item.
 If per-item response times are available, they can easily be incorporated into CODERS as an additional series to be used in the changepoint detection.
-If only per-page response times are available and pages are relatively short, an additional series can be obtained by assigning to each response the average page time, i.e., the time the participant spent on the page on which the item is located divided by the number of items on that page (which is inspired by the page time index of Bowling et al.,, 2023).
+If only per-page response times are available and pages are relatively short, an additional series can be obtained by assigning to each response the average page time, i.e., the time the participant spent on the page on which the item is located divided by the number of items on that page (which is inspired by the page time index of Bowling et al., 2023).
 If a participant at some point in the survey, for instance, starts responding without reading the items, we expect a changepoint in response time towards considerably faster responses (cf.
-Bowling et al.,, 2023, 2016; Meade & Craig,, 2012; Huang et al.,, 2012).
+Bowling et al., 2023, 2016; Meade and Craig, 2012; Huang et al., 2012).
 Nevertheless, it is unclear if a changepoint in response times is indeed indicative of careless responding in an empirical setting.
@@
 On the other hand, if all items are of similar length, a changepoint to (much) shorter response times may indicate carelessness.
-For instance, there exist instruments whose individual items comprise of single words, such as the Big Five unipolar markers (Goldberg,, 1992) with items like “timid” or “envious”.
+For instance, there exist instruments whose individual items comprise of single words, such as the Big Five unipolar markers (Goldberg, 1992) with items like “timid” or “envious”.
 In such measurements, one would not expect large differences in response times across items.
@@
 We introduced a novel method, CODERS, designed to identify the onset of careless responding (or absence thereof) for each participant in a survey.
-It is motivated by literature stressing that survey length contributes to careless responding, in the sense that lengthy surveys—which are common in the behavioral and organizational sciences—may result in a large proportion of survey participants being partially careless, for instance due to fatigue or boredom.
+It is motivated by literature stressing that survey length contributes to careless responding, in the sense that lengthy surveys—which are common in the behavioral sciences—may result in a large proportion of survey participants being partially careless, for instance due to fatigue or boredom.
 CODERS is highly flexible and can combine multiple indicators for the potential onset of carelessness, such as autoencoder reconstruction errors (RE) as an indicator of inconsistency, and the proposed LongStringPattern (LSP) sequence as an indicator of invariability.
@@
 In particular when combined with a post-hoc analysis of the identified response segments, CODERS is able to uncover various types of changes in (careless) response behavior.
-While CODERS is a promising first step towards detecting periods of careless responding rather than identifying which respondents are careless, various pertinent questions remain for future research.
-For instance, the LSP sequence as a measurement of response invariability may require additional validation, as does the incorporation of response times into CODERS, which we leave for empirical follow-up work.
-In addition, extending or adapting CODERS with changepoint detection methods that allow for multiple changepoints, or that retain higher power in shorter surveys, may be fruitful directions for further research.
-Furthermore, it could be interesting to study whether autoencoder reconstruction errors, which can be thought of comparing observed responses with an expectation, are able to distinguish between inconsistent careless responding and other sources of deviations from expected response behavior, such as misunderstanding or misinterpretation of survey items.
+While CODERS is a promising step towards detecting periods of careless responding rather than identifying which respondents are careless, pertinent questions remain for future research.
+Most notably, extending or adapting CODERS with changepoint detection methods that allow for multiple changepoints, or that retain higher power in shorter surveys, may be fruitful directions for further research.
 CODERS naturally integrates with existing techniques that address careless responding.
@@
 It should be noted that some implementations allow traditional screeners to be computed on subsets of responses.
-For instance, the implementation of the intra-individual variance (IRV) (Dunn et al.,, 2018; Marjanovic et al.,, 2015) in the R package careless (Yentes & Wilhelm,, 2023) can be calculated for a given number of subsets of equal length.
+For instance, the implementation of the intra-individual variance (IRV) (Dunn et al., 2018; Marjanovic et al., 2015) in the R package careless (Yentes and Wilhelm, 2023) can be calculated for a given number of subsets of equal length.
 Although this approach is related to the detection of periods of careless responding, it analyzes predefined subsets of responses, while CODERS aims to identify the location of a changepoint.
 Second, pre-administered detection items such as bogus, instructed, or self-report items (e.g.
-Arthur et al.,, 2021; Curran,, 2016; Ward & Meade,, 2023) may provide further validation for the findings of CODERS.
+Arthur et al., 2021; Curran, 2016; Ward and Meade, 2023) may provide further validation for the findings of CODERS.
 An advantage of detection items is that they can in principle detect any type of careless responding.
-By themselves, detection items do not provide accurate estimates of the location of a change in response behavior, at least when following recommendations of Meade & Craig, (2012) to include at most three detection items, spaced every 50–100 items.
+By themselves, detection items do not provide accurate estimates of the location of a change in response behavior, at least when following recommendations of Meade and Craig (2012) to include at most three detection items, spaced every 50–100 items.
 However, detection items can strengthen the findings of CODERS: if passed detection items (i.e., responses expected from an attentive participant) fall with the found attentive period while failed detection items fall within the found careless period, there is convergent evidence of such response behavior.
-Third, various methods have been proposed in the literature to prevent careless responding in the first place, such as promising rewards for responding accurately (Gibson & Bowling,, 2020), or threatening with punishment for careless responding (Bowling et al.,, 2021).
-There is evidence that such measures can prevent or at least delay the onset of careless responding (Bowling et al.,, 2021; Gibson & Bowling,, 2020; Huang et al.,, 2012).
+Third, various methods have been proposed in the literature to prevent careless responding in the first place, such as promising rewards for responding accurately (Gibson and Bowling, 2020), or threatening with punishment for careless responding (Bowling et al., 2021).
+There is evidence that such measures can prevent or at least delay the onset of careless responding (Bowling et al., 2021; Gibson and Bowling, 2020; Huang et al., 2012).
 However, preventive measures can be undesirable or even unethical in practice.
-For instance, Arthur et al., (2021) and Bowling et al., (2021) advise against using threats if one wishes to maintain a positive long-term relationship with survey participants, while Arthur et al., (2021) stress that promising rewards may cross ethical boundaries in organizational surveys.
+For instance, Arthur et al.
+(2021) and Bowling et al.
+(2021) advise against using threats if one wishes to maintain a positive long-term relationship with survey participants, while Arthur et al.
+(2021) stress that promising rewards may cross ethical boundaries in some contexts.
 Notwithstanding, if preventative methods are appropriate, combining them with screening techniques such as CODERS may further enhance quality and reliability of the collected data for subsequent analysis.
 Whereas we apply machine learning and nonparametric techniques in CODERS, other literature is focused on model-based approaches.
-For instance, Ulitzsch et al., 2022b propose a mixture model based on item response theory that is designed to detect careless respondents, which is adapted in Ulitzsch et al., 2022a to incorporate response times.
-Other approaches that explicitly model carelessness as part of a mixture model are proposed in Arias et al., (2020), Steinmann et al., (2022), and Van Laar & Braeken, (2022).
-One advantage of model-based approaches is that their estimates can be used to construct person-weights that can be used to downweight the adverse influence of potentially careless respondents in subsequent analyses, which is discussed in detail in Ulitzsch et al., 2022a ; Ulitzsch et al., 2022b and Hong & Cheng, (2019).
+For instance, Ulitzsch et al.
+(2022b) propose a mixture model based on item response theory that is designed to detect careless respondents, which is adapted in Ulitzsch et al.
+(2022a) to incorporate response times.
+Other approaches that explicitly model carelessness as part of a mixture model are proposed in Zhang et al.
+(2025), Steinmann et al.
+(2022), Van Laar and Braeken (2022), and Arias et al.
+(2020).
+One advantage of model-based approaches is that their estimates can be used to construct person-weights that can be used to downweight the adverse influence of potentially careless respondents in subsequent analyses, which is discussed in detail in Ulitzsch et al.
+(2024, 2022a, 2022b) and Hong and Cheng (2019).
 However, model-based approaches typically impose strong and potentially unverifiable assumptions on the properties of attentive and careless responding.
-On a final note, Arthur et al., (2021) envisage “artificial intelligence and machine learning being used to detect careless responding and response distortion in real time in the foreseeable future.” Using machine learning techniques in the form of autoencoders and nonparametric changepoint detection, our proposed method CODERS takes a step towards that future.
+On a final note, Arthur et al.
+(2021) envisage “artificial intelligence and machine learning being used to detect careless responding and response distortion in real time in the foreseeable future.” Using machine learning techniques in the form of autoencoders and nonparametric changepoint detection, our proposed method CODERS takes a step towards that future.
 Computational Details
-Throughout this paper, we have followed recent guidelines by Alfons & Welz, (2024) on the application of machine learning methods for the detection of careless responding.
-All computations were performed with R version 4.4.2 (R Core Team,, 2024) and Python version 3.10 (Van Rossum & Drake,, 2009).
-Our implementation of CODERS in the R package carelessonset (Welz & Alfons,, 2024) is available from https://github.com/mwelz/carelessonset.
-Replication files will be made publicly available upon acceptance of this manuscript and a link will be included here.
+[…]
 Acknowledgments
@@
 This assumption is very mild in survey data since surveys typically measure multiple constructs, with the number of constructs typically known to the survey designer.
-For instance, the $p=300$ item representation of the NEO-PI-R inventory (Goldberg,, 1999) measures $s=30$ constructs.
+For instance, the $p=300$ item representation of the NEO-PI-R inventory (Goldberg, 1999) measures $s=30$ constructs.
 Assumption 2.
@@
 There exist well-established and reliable survey measures for a large variety of constructs.
-For instance, the International Personality Item Pool (IPIP; Goldberg,, 1999) is a pool of more than 250 personality scales.
-It is in general recommended to use well-established measures to ensure a highly reliable measurement (Babbie,, 2020, p.
+For instance, the International Personality Item Pool (IPIP; Goldberg, 1999) is a pool of more than 250 personality scales.
+It is in general recommended to use well-established measures to ensure a highly reliable measurement (Babbie, 2020, p.
 150).
-The reliability of a measurement can be estimated in multiple ways, such as Cronbach’s alpha, McDonald’s omega and variations thereof, as well as the $H$ coefficient (see, e.g., McNeish,, 2018, for a description of these and other reliability measures).
+The reliability of a measurement can be estimated in multiple ways, such as Cronbach’s alpha, McDonald’s omega and variations thereof, as well as the $H$ coefficient (see, e.g., McNeish, 2018, for a description of these and other reliability measures).
 Hence, this assumption is satisfied when the survey data are collected with measures of high reliability.
@@
 The survey design precludes responses in (fixed) recurring patterns, meaning that such patterns are indicative of invariable careless responding.
-One way to satisfy this assumption is through randomization of the survey items, which is common practice in surveys in the social and behavioral sciences.
-If it is desirable to keep items of the same construct together to reduce the burden on respondents, as is often the case in organizational surveys, one may randomize the order of items within each construct, and randomize the order of constructs (or groups of constructs, if related constructs should be presented together).
+One way to satisfy this assumption is through randomization of the survey items, which is common practice in surveys in the behavioral sciences.
+If it is desirable to keep items of the same construct together to reduce the burden on respondents, one may randomize the order of items within each construct, and randomize the order of constructs (or groups of constructs, if related constructs should be presented together).
 The use of negatively keyed items may also be beneficial.
@@
 The onset of partial carelessness is characterized by a changepoint in one or more indicators. We use autoencoder reconstruction errors (RE) and LongStringPattern (LSP) sequences as indicators.
-A detailed motivation for this assumption is provided in the section introducing CODERS.
+A detailed motivation for this assumption is provided in the section of the main text that introduces CODERS.
 Appendix B Details on CODERS
-B.1 Motivating Novel Methodology
-Before we present a technical description of CODERS, we believe that it is instructive to first take a step back and motivate the application of novel techniques for the task of detecting partially careless responding rather than using existing post-hoc screeners such as psychometric antonym, personal reliability, and the longstring index (see e.g., Alfons & Welz,, 2024, for a recent overview of traditional screeners).
+Motivating Novel Methodology
+Before we present a technical description of CODERS, we believe that it is instructive to first take a step back and motivate the application of novel techniques for the task of detecting partially careless responding rather than using existing post-hoc screeners such as personal reliability, psychometric antonym, and the longstring index (see e.g., Alfons and Welz, 2024, for a recent overview of traditional screeners).
 A primary motivation for devising new methodology is that most existing screeners compute a single carelessness score for each participant.
@@
 In order to obtain a granular estimate of carelessness onset on the participant level, we require a carelessness score for every item for every respondent rather than a singular per-participant carelessness score.
-While some traditional screeners can be somewhat tweaked in this direction (for instance, one could calculate separate intra-individual response variances for separate sets of items; see e.g., Dunn et al.,, 2018; Marjanovic et al.,, 2015), accurate estimation of the onset of carelessness necessitates new methodology.
+While some traditional screeners can be somewhat tweaked in this direction (for instance, one could calculate separate intra-individual response variances for separate sets of items; see e.g., Dunn et al., 2018; Marjanovic et al., 2015), accurate estimation of the onset of carelessness necessitates new methodology.
 As such, CODERS is based on two different carelessness scores on the participant-item level to capture a wide variety of different types of careless responding.
-The first dimension, reconstruction errors (RE) obtained via an auto-associative neural network (autoencoder) (Kramer,, 1992), is intended to measure inconsistent careless responding, such as choosing response categories at random, on the participant-item level.
-Autoencoders are described in detail in Section B.2.
-The second dimension, a LongStringPattern (LSP) sequence, is a generalization of the famous longstring index (Johnson,, 2005) that measures on the participant-item level if a given response is part of a content-independent recurring response sequence, such as $-2-1-2-\dots-1-$ 2 - ⋯ - 1 - 2.
+The first dimension, reconstruction errors (RE) obtained via an auto-associative neural network (autoencoder) (Kramer, 1992), is intended to measure inconsistent careless responding, such as choosing response categories at random, on the participant-item level.
+The second dimension, a LongStringPattern (LSP) sequence, is a generalization of the famous longstring index (Johnson, 2005) that measures on the participant-item level if a given response is part of a content-independent recurring response sequence, such as $1-2-1-2-\dots-1-2$.
 If so, it allows one to trace back when the recurring sequence has started, thereby providing evidence for the onset of invariable carelessness.
-LSP is explained in detail in Section B.3.
-B.2 Details on Autoencoders
-B.2.1 Further Intuition
+Autoencoders and LSP are described in detail in the following sections.
+Details on Autoencoders
+Further Intuition
 Assigning a participant-item score that measures how inconsistently a participant has responded to the item in question arguably is a challenging task.
@@
 We stress that response predictions based on a lower-dimensional representation could also be obtained through principal component analysis (PCA).
-In fact, autoencoders turn out to be a generalization of PCA that, unlike PCA, is not based on linearity assumptions (Kramer,, 1991).
+In fact, autoencoders turn out to be a generalization of PCA that, unlike PCA, is not based on linearity assumptions (Kramer, 1991).
 Since we wish to refrain from possibly restrictive assumptions, we choose autoencoders over PCA to construct lower-dimensional representations of the given responses.
 We are now ready to dive into the formal definition of the autoencoder.
-B.2.2 Network Architecture and Estimation
+Network Architecture and Estimation
 Denote by $x_{ij}$ the response of the $i$-th participant to the $j$-th survey item.
 Collect the responses of the $i$-th participant to all $p$ items in a $p$-dimensional vector $\mathbf{x}_{i}=(x_{i1},x_{i2},\dots,x_{ip})^{\top}$ so that the $n\times p$ data matrix $\bm{X}$ is given by $\bm{X}=(\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{n})^{\top}$.
-Introduced by Kramer, (1992), an autoencoder is a network consisting of an uneven number of $M$ layers, in which the first layer (input layer) holds the input data $\mathbf{x}_{i}$ and the last layer (output layer) holds the reconstructed input data $\widehat{\mathbf{x}}_{i}$.
+Introduced by Kramer (1992), an autoencoder is a network consisting of an uneven number of $M$ layers, in which the first layer (input layer) holds the input data $\mathbf{x}_{i}$ and the last layer (output layer) holds the reconstructed input data $\widehat{\mathbf{x}}_{i}$.
 Denoting the number of nodes in the $\ell$-th layer by $N^{(\ell)}$, we have that $N^{(1)}=N^{(M)}=p$.
@@
 Due to the fact that each node is a function of nodes from a previous layer, the nodes in adjacent layers are often visualized by connecting edges.
-Figure B.1 provides such a schematic overview of a general autoencoder neural network.
-Figure B.1: Schematic overview of an autoencoder neural network with $M$ layers.
+Figure B1 provides such a schematic overview of a general autoencoder neural network.
+Figure B1: Schematic overview of an autoencoder neural network with $M$ layers.
 […]
@@
 For a prespecified loss function $\mathcal{L}(\cdot)$, we fit the neural network by finding the $\bm{\theta}$ that yields the best average reconstruction error,
-$\widehat{\bm{\theta}}=\arg\min_{\bm{\theta}}\Bigg{\{}\frac{1}{n}\sum_{i=1}^{n}\sum_{j=1}^{p}\mathcal{L}\Big{(}x_{ij}-f_{j,\bm{\theta}}(\mathbf{x}_{i})\Big{)}\Bigg{\}}.$ (B.1)
-The prediction function of the fitted network is subsequently given by $\bm{f}_{\widehat{\bm{\theta}}}(\cdot)$.
-For the loss function in (B.1), we choose the smooth and robust Pseudo-Huber loss, which is defined for a fixed $\delta>0$ as
-$\mathcal{L}(z)=\delta^{2}\left(\sqrt{1+(z/\delta)^{2}}-1\right).$ (B.2)
+$\widehat{\bm{\theta}}=\arg\min_{\bm{\theta}}\Bigg\{\frac{1}{n}\sum_{i=1}^{n}\sum_{j=1}^{p}\mathcal{L}\Big(x_{ij}-f_{j,\bm{\theta}}(\mathbf{x}_{i})\Big)\Bigg\}.$ (B1)
+[…]
+$\mathcal{L}(z)=\delta^{2}\left(\sqrt{1+(z/\delta)^{2}}-1\right).$ (B2)
 This choice results in quadratic loss for small values of $z$ and linear loss for large values of $z$.
 Consequently, the Pseudo-Huber loss function avoids that large individual reconstruction errors strongly affect the fit, which would make it hard for the network to distinguish attentive responses from inconsistent careless responses, as the latter typically lead to large prediction errors.
-To solve the optimization problem in (B.1), we use stochastic gradient descent.
-We refer to Chapter 8.5 in Goodfellow et al., (2016) for details.
-B.2.3 Modeling Choices
-Following Kramer, (1992), we specify $M=5$ layers in the autoencoder.
+To solve the optimization problem in (B1), we use stochastic gradient descent.
+We refer to Chapter 8.5 in Goodfellow et al.
+(2016) for details.
+Modeling Choices
+Following Kramer (1992), we specify $M=5$ layers in the autoencoder.
 We set the number of nodes in layers 2 and 4 (mapping an demapping layer, respectively) to $\lfloor 1.5\times p\rfloor$, as a relatively large number of nodes is expected to give the autoencoder flexibility to learn many different types of response behavior (attentive and careless).
 Based on Assumption 1, the number of nodes in the bottleneck layer equals the number of constructs measured in the questionnaire.
-Concerning the activation functions, we again follow a recommendation in Kramer, (1992) and propose to use nonlinear activation functions in the mapping and demapping layers, as well as a linear activation function in the bottleneck layer.
+Concerning the activation functions, we again follow a recommendation in Kramer (1992) and propose to use nonlinear activation functions in the mapping and demapping layers, as well as a linear activation function in the bottleneck layer.
 Specifically, we propose to use the hyperbolic tangent activation in the (de-)mapping layers and the identity mapping in the bottleneck layer.
-Definitions are given in Table B.1, which summarizes our proposed autoencoder architecture.
-Table B.1: Choices regarding the network architecture of the autoencoder.
+Definitions are given in Table B1, which summarizes our proposed autoencoder architecture.
+Table B1: Choices regarding the network architecture of the autoencoder.
 Layer Number of nodes Activation function
@@
 For fitting the autoencoder, we use the stochastic gradient descent algorithm with a batch size of 10, a learning rate of 0.0001, and 100 epochs.
-In the pseudo-Huber loss (B.2), we set constant $\delta=1$.
-B.3 LongStringPattern Algorithm
+In the pseudo-Huber loss (B2), we set constant $\delta=1$.
+LongStringPattern Algorithm
 Formal algorithms for computing the $l$-pattern and LSP sequences are provided in Algorithms 1 and 2, respectively.
 Input: Series of observed rating-scale responses $\{x_{j}\}_{j=1}^{p}$, pattern length $l\in\mathbb{N}$ to consider (number of items that an individual pattern encompasses).
-For instance, the individual pattern $1-2-3$ encompasses $l=3$ items in the response sequence $-3-1-2-$ 3 - 1 - 2 - 3.
+For instance, the individual pattern 1231-2-3 encompasses $l=3$ items in the response sequence 1231231-2-3-1-2-3.
 Note that the algorithm measures straightlining when setting $l=1$.
@@
 Initialization: $streak\leftarrow 1,startindex\leftarrow 1$
-1
-/* fill in the first $l-1$ indices */
-2 for $k\leftarrow 1$ to $p-l-1$ by $1$ do
-3       if $x[k]==x[k+l]$ then
-            increment $streak$ by 1 ;
+/* fill in the first l1l-1 indices */
+for $k\leftarrow 1$ to pl1p-l-1 by 11 do
+ if $x[k]==x[k+l]$ then
+     increment streakstreak by 1 ;
     // streak is ongoing
-4
-5      else
-             Assign value $streak$ to all elements in $\{l\textsf{-pattern}_{j}\}_{j=startindex}^{k}$ ;
+ else
+      Assign value streakstreak to all elements in $\{l\textsf{-pattern}_{j}\}_{j=startindex}^{k}$ ;
     // streak broke
-6             $startindex\leftarrow k+1$;
-7             streak $\leftarrow 1$;
-8
-9       end if
-10
-11 end for
-12
+    $startindex\leftarrow k+1$;
+      streak $\leftarrow 1$;
+   end if
+end for
 /* fill in the last $l$ indices (there are 3 edge cases) */
-13 if $\texttt{streak}\geq l$ then
+if $\texttt{streak}\geq l$ then
  /* case 1: current streak is ongoing and we have seen it completely at least once, so assume it will last until the end and complete the current streak (which has not yet been broken) */
-14       Assign value $streak+l-1$ to all elements in $\{l\textsf{-pattern}_{j}\}_{j=startindex}^{p}$;
-15
-16 else if $1<\texttt{streak}<l$ then
+   Assign value streak+l1streak+l-1 to all elements in $\{l\textsf{-pattern}_{j}\}_{j=startindex}^{p}$;
+else if $1<\texttt{streak}<l$ then
  /* case 2: there is an ongoing streak, but we have no chance of seeing it completely (since streak < l), so we do not know how the streak looks like.
 In this case, fill in the current value of "streak" until the (p-l)-th item and fill in 1 for the last l items (because we do not know what the complete streak looks like) */
-17       Assign value $streak-1$ to all elements in $\{l\textsf{-pattern}_{j}\}_{j=startindex}^{p-l}$;
-18       Assign value $1$ to all elements in $\{l\textsf{-pattern}_{j}\}_{j=p-l+1}^{p}$;
-19
-20 else
+   Assign value streak1streak-1 to all elements in $\{l\textsf{-pattern}_{j}\}_{j=startindex}^{p-l}$;
+   Assign value 11 to all elements in $\{l\textsf{-pattern}_{j}\}_{j=p-l+1}^{p}$;
+else
  /* case 3: there is no ongoing streak, so fill in 1 for last $l$ indices */
-21       Assign value 1 to all elements in $\{l\textsf{-pattern}_{j}\}_{j=p-l+1}^{p}$;
-22
-23 end if
+   Assign value 1 to all elements in $\{l\textsf{-pattern}_{j}\}_{j=p-l+1}^{p}$;
+end if
 Algorithm 1 $l$-pattern($\{x_{j}\}_{j=1}^{p},l$)
@@
 Output: Set of LongStringPattern measures $\{\textsf{LSP}_{j}\}_{j=1}^{p}$
-1
-/* calculate $l$-pattern for fixed pattern length $l$/
-2 for $l\leftarrow 1$ to $l_{max}$ by $1$ do
-3       $\{\textsf{LSP}^{l}_{j}\}_{j=1}^{p}\leftarrow l\textsf{-pattern}(\{x_{j}\}_{j=1}^{p},l)$;
-4
-5 end for
-6
+/* calculate $l$-pattern for fixed pattern length $l$ */
+for $l\leftarrow 1$ to $l_{max}$ by 11 do
+ $\{\textsf{LSP}^{l}_{j}\}_{j=1}^{p}\leftarrow l\textsf{-pattern}(\{x_{j}\}_{j=1}^{p},l)$;
+end for
 /* for each element $\{1,\dots,p\}$, report the largest $l$-pattern value */
-7 Initialize $\{\textsf{LSP}_{j}\}_{j=1}^{p}$;
-8 for $j\leftarrow 1$ to $p$ by $1$ do
-9       $\textsf{LSP}_{j}\leftarrow\max\big{\{}\textsf{LSP}_{j}^{1},\textsf{LSP}_{j}^{2},\dots,\textsf{LSP}_{j}^{l}\big{\}}$;
-10
-11 end for
+Initialize $\{\textsf{LSP}_{j}\}_{j=1}^{p}$;
+for $j\leftarrow 1$ to $p$ by 11 do
+ $\textsf{LSP}_{j}\leftarrow\max\big\{\textsf{LSP}_{j}^{1},\textsf{LSP}_{j}^{2},\dots,\textsf{LSP}_{j}^{l}\big\}$;
+end for
 Algorithm 2 LongStringPattern($\{x_{j}\}_{j=1}^{p},l_{max}$)
-B.4 Details on Changepoint Detection
+Details on Changepoint Detection
 Let $\{\bm{Y}_{j}\}_{j=1}^{p}$ be a series of length $p$ consisting of $d$-dimensional random variables.
-The goal of the cumulative sum self-normalization test of Shao & Zhang, (2010) is to estimate the location of a possible changepoint in the (multivariate) series, i.e., to divide the series into two subsets $\{\bm{Y}_{j}\}_{j=1}^{k-1}$ and $\{\bm{Y}_{j}\}_{j=k}^{p}$ in which a $d$-dimensional parameter of interest $\bm{\theta}$ differs.
+The goal of the cumulative sum self-normalization test of Shao and Zhang (2010) is to estimate the location of a possible changepoint $k$ in the (multivariate) series, i.e., to divide the series into two subsets $\{\bm{Y}_{j}\}_{j=1}^{k-1}$ and $\{\bm{Y}_{j}\}_{j=k}^{p}$ in which a $d$-dimensional parameter of interest $\bm{\theta}$ differs.
 Here, we focus on the mean as the parameter $\bm{\theta}$.
@@
 with
-$\bm{L}_{p}(k)=\left\{\begin{array}[]{ll}\displaystyle\sum_{i=1}^{k-2}\frac{i^{2}(k-1-i)^{2}}{p^{2}(k-1)^{2}}\left(\widehat{\bm{\theta}}_{1,i}-\widehat{\bm{\theta}}_{i+1,k-1}\right)\left(\widehat{\bm{\theta}}_{1,i}-\widehat{\bm{\theta}}_{i+1,k-1}\right)^{\top},&\qquad\text{if }k>2,\\ 0,&\qquad\text{if }k=2,\\ \end{array}\right.$
+[…]
 and
-$\bm{R}_{p}(k)=\left\{\begin{array}[]{ll}\displaystyle\sum_{i=k+1}^{p}\frac{(p-i+1)^{2}(i-k)^{2}}{p^{2}(k-1)^{2}}\left(\widehat{\bm{\theta}}_{i,p}-\widehat{\bm{\theta}}_{k,i-1}\right)\left(\widehat{\bm{\theta}}_{i,p}-\widehat{\bm{\theta}}_{k,i-1}\right)^{\top},&\qquad\text{if }k<p,\\ 0,&\qquad\text{if }k=p,\\ \end{array}\right.$
+[…]
 The final test statistic is then defined as
-$SN_{p}=\max_{k=2,\dots,p}T_{p}(k).$ (B.3)
+$SN_{p}=\max_{k=2,\dots,p}T_{p}(k).$ (B3)
 Consider a given threshold $K_{d}>0$. If the test statistic exceeds the threshold, $SN_{p}>K_{d}$, an estimate of the changepoint location is obtained via
@@
 Otherwise, no changepoint is detected.
-Shao & Zhang, (2010) and Zhao et al., (2022) derive theoretical guarantees for this procedure.
-Specifically, Shao & Zhang, (2010) obtain the asymptotic distribution of the test statistic $SN_{p}$ for $p\rightarrow\infty$ and provide the corresponding critical values $K_{d}$ for a given significance level $\alpha$ and dimension $d$.
-We summarize relevant critical values in Table B.2.
-Table B.2: Critical values $K_{d}$ for relevant significance levels $\alpha$ and dimensions $d$.
+Shao and Zhang (2010) and Zhao et al.
+(2022) derive theoretical guarantees for this procedure.
+Specifically, Shao and Zhang (2010) obtain the asymptotic distribution of the test statistic $SN_{p}$ for $p\rightarrow\infty$ and provide the corresponding critical values $K_{d}$ for a given significance level $\alpha$ and dimension $d$.
+We summarize relevant critical values in Table B2.
+Table B2: Critical values $K_{d}$ for relevant significance levels $\alpha$ and dimensions $d$.
 Dimension $d$
@@
 0.001 121.9 192.5
-Note. The critical values are taken from Table 1 in Shao & Zhang, (2010).
-In the context of this paper, we use the test statistic from (B.3) to test for changepoints in the two-dimensional series
+Note. The critical values are taken from Table 1 in Shao and Zhang (2010).
+In the context of this paper, we use the test statistic from (B3) to test for changepoints in the two-dimensional series
 $\bm{Y}_{j}=\begin{pmatrix}\textsf{RE}_{j}\\ \textsf{LSP}_{j}\end{pmatrix},\qquad j=1,\dots,p,$
@@
 That is, the detected changepoint $\widehat{k}$ may be viewed as the onset of careless responding.
-We obtain such a series for each of the $n$ survey participants and subsequently test each of the $n$ series for changepoints.
+We obtain a series $\{\bm{Y}_{j}\}_{j=1}^{p}$ for each of the $n$ survey participants and subsequently test each of the $n$ series for changepoints.
 On a final technical note, LSP sequences may exhibit low variation.
@@
 Appendix C Additional Simulations
-C.1 Additional Results for the Main Design
-In addition to CODERS, we apply three traditional methods for the detection of careless respondents in the simulation design from the main text, which is inspired by the data collected by Johnson, (2005): personal reliability, psychometric antonym, and the longstring index.
-We choose these three methods since they were also applied in the study of Johnson, (2005).
-Following Johnson,, we reverse psychometric antonym scores, and we flag respondents as careless whose personal reliability and psychometric antonym scores are lower than $0.3$ and $-0.03$, respectively.
-For the longstring index, we flag a respondent as careless if its value exceeds 6 (i.e., $>6$), which is a cutoff value suggested by Niessen et al., (2016).
-Figure C.1 visualizes the false positive rate (FPR) and the false negative rate (FNR) of the three methods for the different carelessness onset regimes (in separate columns).
-Figure C.2 further shows the false negative rates for the four types of partially careless respondents (in separate rows).
+Additional Results for the Main Design
+In addition to CODERS, we apply three traditional methods for the detection of careless respondents in the simulation design from the main text, which is inspired by the data collected by Johnson (2005): personal reliability, psychometric antonym, and the longstring index.
+We choose these three methods since they were also applied in the study of Johnson (2005).
+Following Johnson, we reverse psychometric antonym scores, and we flag respondents as careless whose personal reliability and psychometric antonym scores are lower than $0.3$ and 0.03-0.03, respectively.
+For the longstring index, we flag a respondent as careless if its value exceeds 6 (i.e., >6>6), which is a cutoff value suggested by Niessen et al.
+(2016).
+Figure C1 visualizes the false positive rate (FPR) and the false negative rate (FNR) of the three methods for the different carelessness onset regimes (in separate columns).
+Figure C2 further shows the false negative rates for the four types of partially careless respondents (in separate rows).
 The longstring index performs perfectly at its intended task of detecting straightlining, whereas it hardly detects any random responding or pattern responding (which is expected).
@@
 As expected, the false negative rate of the traditional methods in general worsens the later the onset of careless responding (while that of CODERS deteriorates the earlier the onset, see Figures 7 and 8 in the main text).
-Figure C.1: Performance measures for the traditional methods used by Johnson, (2005).
+Figure C1: Performance measures for the traditional methods used by Johnson (2005).
 Note.
@@
 False positive rates in attentive respondents (top row) and false negative rates in partially careless respondents (bottom row) for different onset regimes (columns) are averaged across 100 repetitions.
-Figure C.2: False negative rate of the traditional methods used by Johnson, (2005) for different types of partially careless respondents.
+Figure C2: False negative rate of the traditional methods used by Johnson (2005) for different types of partially careless respondents.
 Note. The $x$-axis displays various prevalence levels. Results for different types of careless responding (rows) in different onset regimes (columns) are averaged across 100 repetitions.
-C.2 Decreasing Survey Length
-Recall that the changepoint detection method of Shao & Zhang, (2010) used in CODERS is based on asymptotic arguments for the number of items being large (see Appendix B.4).
+Decreasing Survey Length
+Recall that the changepoint detection method of Shao and Zhang (2010) used in CODERS is based on asymptotic arguments for the number of items being large (see Appendix B).
 It therefore seems plausible that the ability of CODERS to accurately identify partial carelessness may decline as survey length decreases.
-C.2.1 Data Generation
+Data Generation
 We use the same simulation design of the main text, but we focus on the baseline onset regime and instead decrease the number of items $p$ by gradually removing sets of items.
@@
 We drop up to four traits, resulting in surveys comprising a total of $p\in\{240,180,120,60\}$ items.
-C.2.2 Results
-Figure C.3 plots the false positive rates for the different survey lengths (in separate columns).
+Results
+Figure C3 plots the false positive rates for the different survey lengths (in separate columns).
 As before, CODERS rarely incorrectly flags attentive respondents as careless at any of the considered significance levels, although there is some overrejection for the shortest survey with $p=60$.
-Figure C.4 visualizes the false negative rate (top row) and the mean absolute error of onset estimation (bottom row) for careless respondents across the different surveys lengths (in separate columns).
+Figure C4 visualizes the false negative rate (top row) and the mean absolute error of onset estimation (bottom row) for careless respondents across the different surveys lengths (in separate columns).
 At a given significance level, the false negative rate slightly increases with decreasing survey length compared to the results from the main text.
@@
 In addition, the location of the onset of carelessness is still estimated highly accurately throughout the investigated survey lengths, on average being off by only about one or two items.
-Figure C.3: False positive rate of CODERS for attentive respondents across decreasing survey length.
+Figure C3: False positive rate of CODERS for attentive respondents across decreasing survey length.
 Note. The $x$-axis displays various prevalence levels of partially careless respondents. Results for different survey lengths are shown in separate columns, averaged across 100 repetitions.
-Figure C.4: Performance measures of CODERS for partially careless respondents across decreasing survey length.
+Figure C4: Performance measures of CODERS for partially careless respondents across decreasing survey length.
 Note.
@@
 Results for different survey lengths are shown in separate columns, averaged across 100 repetitions.
-Figure C.5: False negative rate of CODERS for different types of partially careless respondents across decreasing survey length.
+Figure C5: False negative rate of CODERS for different types of partially careless respondents across decreasing survey length.
 Note. The $x$-axis displays various prevalence levels. Results for different types of careless responding (rows) in different survey lengths (columns) are averaged across 100 repetitions.
-Figure C.6: Mean absolute error of carelessness onset estimation with CODERS for different types of partially careless respondents across decreasing survey length.
+Figure C6: Mean absolute error of carelessness onset estimation with CODERS for different types of partially careless respondents across decreasing survey length.
 Note. The $x$-axis displays various prevalence levels. Results for different types of careless responding (rows) in different survey lengths (columns) are averaged across 100 repetitions.
-Figures C.5 and C.6 show the false negative rate and the mean absolute error of onset estimation, respectively, grouped by carelessness type (rows) for each considered survey length (columns).
+Figures C5 and C6 show the false negative rate and the mean absolute error of onset estimation, respectively, grouped by carelessness type (rows) for each considered survey length (columns).
 In essence, the slight increase in false negative rates seems to be primarily caused by fewer inconsistent careless respondents being detected, whereas the performance for invariable careless responding stays near-perfect.
@@
 Furthermore, the power loss can be partially alleviated by choosing a less conservative significance level, such as 0.5% or 1%.
-C.3 Additional Sample Sizes
+Additional Sample Sizes
 Although the number of items in the survey may be the primary driver regarding power of the changepoint detection test used in CODERS, the autoencoder may also benefit from more information in the form of a larger sample size in order to learn attentive response behavior (and conversely suffer under a smaller sample size).
-C.3.1 Data Generation
-[…]
-C.3.2 Results
-Figures C.7C.10 visualize the performance measures for $n=248$ and $n=1,000$ simulated participants (in separate columns).
+Data Generation
+We again use the same simulation design of the main text and focus on the baseline onset regime, but we now generate a smaller and a larger number of participants, namely $n=248$ and $n=\text{1,000}$, respectively.
+Note that the sample size in our simulations should be divisible by 4 since we generate four types of partially careless respondents.
+This explains the perhaps unusual choice of $n=248$ (rather than $n=250$).
+Results
+Figures C7C10 visualize the performance measures for $n=248$ and $n=1,000$ simulated participants (in separate columns).
 The results are similar to those in the main text, with the overall FNR in partially careless respondents increasing only slightly for the smaller sample size and decreasing only slightly for the larger sample size.
-For random respondents, however, the increase in false negatives in the smaller sample size is notable, but so is the decrease in FNR in the larger sample size (see Figure C.9, top row).
+For random respondents, however, the increase in false negatives in the smaller sample size is notable, but so is the decrease in FNR in the larger sample size (see Figure C9, top row).
 This can likely be explained by the amount of information the autoencoder has available in order to learn to distinguish between attentive responding and random responding.
 It is therefore conceivable that the power for detecting random respondents may improve further in even larger sample sizes.
-Figure C.7: False positive rate of CODERS for attentive respondents across additional sample sizes.
+Figure C7: False positive rate of CODERS for attentive respondents across additional sample sizes.
 Note. The $x$-axis displays various prevalence levels of partially careless respondents. Results for different sample sizes are shown in separate columns, averaged across 100 repetitions.
-Figure C.8: Performance measures of CODERS for partially careless respondents across additional sample sizes.
+Figure C8: Performance measures of CODERS for partially careless respondents across additional sample sizes.
 Note.
@@
 Results for different sample sizes are shown in separate columns, averaged across 100 repetitions.
-Figure C.9: False negative rate of CODERS for different types of partially careless respondents across additional sample sizes.
+Figure C9: False negative rate of CODERS for different types of partially careless respondents across additional sample sizes.
 Note. The $x$-axis displays various prevalence levels. Results for different types of careless responding (rows) in different sample sizes (columns) are averaged across 100 repetitions.
-Figure C.10: Mean absolute error of carelessness onset estimation with CODERS for different types of partially careless respondents across additional sample sizes.
+Figure C10: Mean absolute error of carelessness onset estimation with CODERS for different types of partially careless respondents across additional sample sizes.
 Note. The $x$-axis displays various prevalence levels. Results for different types of careless responding (rows) in different sample sizes (columns) are averaged across 100 repetitions.
-C.4 Randomly Choosing from the Middle Answer Categories
+Randomly Choosing from the Middle Answer Categories
 Recall that the simulation design in the main text considers four types of careless responding: random responding, extreme responding, pattern responding, and straightlining.
 However, there may be other relevant types of carelessness beyond those four, including ones that are extremely difficult to distinguish from attentive response behavior.
-C.4.1 Data Generation
+Data Generation
 As before, we take the simulation design from the main text with the baseline onset regime, but instead of the four aforementioned types of careless responding, we generate a distinct type that we refer to as middling.
@@
 Hence, it is difficult to systematically distinguish middling carelessness from attentive responding, but there is also the risk that attentive respondents might be mistaken for middling careless respondents.
-C.4.2 Results
-Figures C.11 and C.12 visualizes the performance of CODERS when partially careless respondents are of the middling type.
+Results
+Figures C11 and C12 visualizes the performance of CODERS when partially careless respondents are of the middling type.
 One the one hand, CODERS still almost never incorrectly flags an attentive respondent as partially careless.
@@
 Hence, these simulations further demonstrate that CODERS requires inconsistent responding to be sufficiently different from attentive responding in order to distinguish between the two.
-Figure C.11: False positive rate of CODERS for attentive respondents in the presence of partially careless respondents that randomly choose from the middle answer categories.
+Figure C11: False positive rate of CODERS for attentive respondents in the presence of partially careless respondents that randomly choose from the middle answer categories.
 Note. The $x$-axis displays various prevalence levels of partially careless respondents. Results are averaged across 100 repetitions.
-Figure C.12: Performance measures of CODERS for partially careless respondents that randomly choose from the middle answer categories.
+Figure C12: Performance measures of CODERS for partially careless respondents that randomly choose from the middle answer categories.
 Note.
@@
 In order to further highlight the difficulty of detecting middling carelessness, we again apply the traditional methods personal reliability, psychometric antonym, and the longstring index.
-The false positive rates and false negative rates are shown in Figure C.13.
+The false positive rates and false negative rates are shown in Figure C13.
 As in the simulation design from the main text, personal reliability flags hardly any respondents, resulting in false positive rates close to 0 and false negative rates close to 1.
@@
 Psychometric antonym identifies more middling respondents than CODERS with a false negative rate of about 65%–75%, but it exhibits an inflated false positive rate of up to about 15%.
-Figure C.13: Performance measures of the traditional methods for partially careless respondents that randomly choose from the middle answer categories.
+Figure C13: Performance measures of the traditional methods for partially careless respondents that randomly choose from the middle answer categories.
 Note.
@@
 False positive rates in attentive respondents (top row) and false negative rates in partially careless respondents (bottom row) are averaged across 100 repetitions.
-C.5 Temporary Carelessness
+Temporary Carelessness
 It seems plausible that more than one changepoint may occur in a participant’s response behavior.
 For instance, a respondent may start the survey attentively, but become careless at some point due to being distracted, only to start responding attentively again once the distraction vanishes.
-C.5.1 Data Generation
+Data Generation
 We use the same simulation design from the main text, but instead of simulating only one changepoint, we simulate temporary carelessness as follows.
@@
 Between these two changepoints, we generate careless responses using the same four types of carelessness from the main text.
-C.5.2 Results
-Figures C.14 and C.15 visualize the performance of CODERS and Figures C.16 and C.17 that of the traditional methods.
+Results
+Figures C14 and C15 visualize the performance of CODERS and Figures C16 and C17 that of the traditional methods.
 Clearly, CODERS rarely flags any participants, with a false positive rate close to 0 and a false negative rate close to 1 (we therefore omit separate results for the four types of partially careless responding).
@@
 The longstring index yields somewhat better performance, with a relatively low false positive rate and a false negative rate around 65%.
-As in the simulation design from the main text, it successfully captures straightliners as well as some extreme respondents (see Appendix C.1 for a more detailed discussion).
+As in the simulation design from the main text, it successfully captures straightliners as well as some extreme respondents (see the corresponding section above for a more detailed discussion).
 In summary, CODERS is simply not designed for detecting more than one changepoint in respondent behavior, such that an extension to multiple changepoint detection may be a fruitful direction for further research.
-Figure C.14: False positive rate of CODERS for attentive respondents in case of temporary carelessness.
+Figure C14: False positive rate of CODERS for attentive respondents in case of temporary carelessness.
 Note. The $x$-axis displays various prevalence levels of partially careless respondents. Results are averaged across 100 repetitions.
-Figure C.15: Performance measures of CODERS for partially careless respondents in case of temporary carelessness.
+Figure C15: Performance measures of CODERS for partially careless respondents in case of temporary carelessness.
 Note.
@@
 Results are averaged across 100 repetitions.
-Figure C.16: Performance measures for the traditional methods in case of temporary carelessness.
+Figure C16: Performance measures for the traditional methods in case of temporary carelessness.
 Note.
@@
 False positive rates in attentive respondents (top row) and false negative rates in partially careless respondents (bottom row) are averaged across 100 repetitions.
-Figure C.17: False negative rate of the traditional methods for different types of partially careless respondents in case of temporary carelessness.
+Figure C17: False negative rate of the traditional methods for different types of partially careless respondents in case of temporary carelessness.
 Note. The $x$-axis displays various prevalence levels. Results for different types of careless responding are shown in separate columns, averaged across 100 repetitions.
-C.6 Organizational Survey
-Until here, all simulations are based on a measurement of the Big Five personality traits.
-To study how CODERS fares in surveys of different nature, we further apply a simulation design based on a study from organizational research.
-C.6.1 Data Generation
-We simulate a hypothetical survey inspired by a study conducted among university students working in teams on a project for a Strategic Management course.
-The measurement includes various individual- and team-level constructs, among others the HEXACO-60 personality assessment of Ashton & Lee, (2009), the goal orientation scales of Vandewalle, (1997), the transactive memory systems scales of Lewis, (2003), the intra-group conflict scales of Jehn, (1995), the intra-unit justice climate scales of Li & Cropanzano, (2009), a team commitment scale based on Mowday et al., (1979), and the team performance scale of Hackman, (1986).
-In total, the hypothetical survey includes $p=198$ items measuring $q=34$ constructs (which we set as the number of nodes in the bottleneck layer of the autoencoder in CODERS).
-Figure C.18: Correlation matrix and marginal response distributions used in the simulated organizational survey.
-(a) Correlation matrix of the $p=198$ items. A more saturated blue (red) in a cell corresponds to a stronger positive (negative) correlation between the corresponding items.
-(b) Marginal response distributions of the $p=198$ items ($x$-axis), each of which has five response options ($y$-axis).
-A darker shade of gray indicates a higher probability of the corresponding response option.
+Appendix D Details on the Empirical Application
+Data Preprocessing
+Among other information, the data of Johnson (2005), provided at https://osf.io/sxeq5/, contain per-item responses to $p=300$ IPIP-NEO-300 items for each of the the $n=\text{20,993}$ respondents.
+The responses in the provided data have already underwent some preprocessing that is described in detail the accompanying README file.
+Specifically, responses to negatively-worded items were reverse-coded, that is, a given response of 5 was recoded to 1, 4 was recoded to 3, and so on.
+Thus, in order to recover the original responses, we reverted this recoding.
+Furthermore, a total of 33,691 individual responses are missing (0.53% of all responses).
+Note that Johnson had already removed participants with too many missing responses from the provided data set, hence there are relatively few missing responses for the remaining participants.
+Due to the low number of missing responses, we treat them as a separate answer category by coding them as 0 when applying CODERS and the traditional screeners for careless responding.
+Additional Results
+Figure D1 provides a histogram as well as a density estimate for the per-respondent average reconstruction errors $\textsf{RE}_{i}=\frac{1}{p}\sum_{j=1}^{p}\textsf{RE}_{ij}$, $i=1,\dots,n$, where averages are taken over the $p=300$ items.
+We can see that the density has a long right tail, which might be due to careless respondents.
+We highlight the value 0.079 by a vertical dotted line, which is the average reconstruction error of the participant from Figure 11(f) and Table 3(f) in the main text.
+This value is in the 99.98-th percentile of all $n=\text{20,993}$ average reconstruction errors, with only 6 participants having a higher value.
+Given the extremely high average reconstruction error and the fact that this respondent is also flagged by psychometric antonym, it is likely that the participant responded carelessly throughout the survey, explaining why CODERS did not flag a changepoint.
+Figure D1: Histogram of average reconstruction errors for the data of Johnson (2005).
 Note.
-The correlation matrix and the marginal response distributions are calculated from empirical data.
-Of the $p=198$ items, 152 are positively scored and 46 are negatively scored.
-The items in both plots are presented in an order such that items measuring the same construct are adjacent.
-Although team-level constructs are measured, we simplify the data generating process and simulate individual-level data using the correlation matrix and marginal distributions displayed in Figure C.18, which are based on empirical data from 322 students (after removing 29 respondents that contain missing values).
-Unfortunately, we are not at liberty to share the original data.
-Of the $p=198$ items, 152 are positively scored and 46 are negatively scored.
-The different constructs are clearly visible along the diagonal of the correlation matrix, as are various relationships between many of the measured constructs in the off-diagonal parts (cf.
-also Alfons et al., 2022a, ; Alfons et al., 2022b, ).
-We randomize the order of the items in a way that ensures that items measuring the same construct remain adjacent, which is better suited for surveys in organizational contexts.
-That is, we first randomize the order of the items within each construct, and we randomize the order of (groups of related) constructs.
-Otherwise, we keep the same design as in the simulations from the main text.
-In particular, we simulate responses for $n=500$ participants, using the same settings for generating partially careless respondents.
-C.6.2 Results
-Figure C.19 shows the false positive rates for attentive respondents.
-Compared to the simulations from the main text, false positive rates are inflated for high prevalence levels, but they remain reasonably close to the nominal level while carelessness prevalence does not exceed 40% (in the early onset regime) or 30% (for the other onset regimes).
-A possible explanation for this finding may be that many measured constructs do not use negatively-scored items while others do, such that response behavior may be naturally different in some parts of the survey than in others.
-Figure C.19: False positive rate of CODERS for attentive respondents in the simulated organizational survey.
-Note. The $x$-axis displays various prevalence levels of partially careless respondents. Results for different onset regimes are shown in separate columns, averaged across 100 repetitions.
-In Figures C.20C.22, the performance measures for (different types) of partially careless respondents are displayed.
-The overall false negative rate is similar to that of the simulations from the main text, despite the shorter survey length.
-But interestingly, there are slightly fewer false negatives for random responding and slightly more false negatives for extreme responding.
-This may be explained by differences in response distributions, with higher skewness in some of the measurements compared to the data by Johnson, (2005) (cf.
-Figure C.18 with Figure 5 from the main text).
-In addition, the onset of careless responding is estimated with similar accuracy, with an error of about two to three items averaged across the different types of partially careless responding.
-All in all, these results are an indication that the performance of CODERS may not depend too much on the institutional setting of the survey.
-Figure C.20: Performance measures of CODERS for partially careless respondents in the simulated organizational survey.
-Note.
-The $x$-axis displays various prevalence levels.
-The top row shows results for the false negative rate, while the bottom row visualizes the accuracy of onset estimation via the mean absolute error.
-Results for different onset regimes are shown in separate columns, averaged across 100 repetitions.
-Figure C.21: False negative rate of CODERS for different types of partially careless respondents in the simulated organizational survey.
-Note. The $x$-axis displays various prevalence levels. Results for different types of careless responding (rows) in different onset regimes (columns) are averaged across 100 repetitions.
-Figure C.22: Mean absolute error of carelessness onset estimation with CODERS for different types of partially careless respondents in the simulated organizational survey.
-Note. The $x$-axis displays various prevalence levels. Results for different types of careless responding (rows) in different onset regimes (columns) are averaged across 100 repetitions.
-Appendix D Details on the Empirical Application
-D.1 Data Preprocessing
-[…]
-D.2 Additional Results
-[…]
-Figure D.1: Histogram of average reconstruction errors for the data of Johnson, (2005).
-[…]
+The $x$-axis contains the per-respondent average reconstruction errors over the $p=300$ items.
+On the $y$-axis, a histogram of the average reconstruction errors for $n=\text{20,993}$ respondents is displayed together with a density estimate.
+The vertical dotted line denotes the average reconstruction error for the participant in Figure 11(f) and Table 3(f) in the main text.
+This value, 0.079, lies in the 99.98-th percentile of all average reconstruction errors.